### v1

# The Diophantine problem in isotropic reductive groups

Abstract
We begin to study model-theoretic properties of non-split isotropic reductive group schemes. In this paper we show that the base ring $K$ is e-interpretable in the point group $G(K)$ of every sufficiently isotropic reductive group scheme $G$. In particular, the Diophantine problems in $K$ and $G(K)$ are equivalent. We also compute the centralizer of the elementary subgroup of $G(K)$ and the common normalizer of all its root subgroups.
1 Introduction
Let $M$ be a model of some first-order language $L$ with equality, i.e. $M$ is a set with interpretation of constant symbols, functional symbols, and predicate symbols from $L$. For example, $M$ may be a group with constant $1$, operations of multiplication and inversion, and the equality predicate. Recall that an elementary formula with the variables from $\vec{x}=(x_{1},\ldots,x_{n})$ is a formula in the language $L$ consisting of a single predicate symbol with substituted terms involving only variables from $\vec{x}$ and constants from $L$. A formula
$\exists\vec{y}\enskip\bigwedge_{i=1}^{n}P_{i}(\vec{x},\vec{y})$
for elementary formulae $P_{i}(\vec{x},\vec{y})$ is called Diophantine (also positive-primitive or regular) in the variables $\vec{x}$. Finally, a subset $X\subseteq M^{n}$ is called Diophantine with respect to parameters $\vec{a}\in M^{m}$ if
$X=\{\vec{x}\in M^{n}\mid\varphi(\vec{x},\vec{a})\},$
where $\varphi$ is a Diophantine formula in the variables $\vec{x}$ and $\vec{a}$.
A closely related is the notion of e-interpretation. An e-interpretation of a model $M$ in a model $N$ (of possibly another language) is a surjection $X\to M$ for some Diophantine $X\subseteq N^{n}$ such that the lifts of all relations on $M$ to $X$ from its language are Diophantine (including the kernel pair lifting the equality), as well as lifts of all graphs of operations on $M$. For example, if $K$ is a commutative ring and $G$ is a finitely presented affine group scheme over $K$, then the group $G(K)$ is e-interpretable in the ring $K$.
Suppose that $M$ and $L$ are enumerated by $\mathbb{N}$ or its finite subsets. The Diophantine problem $\mathcal{D}(M)$ is decidable if the non-emptiness of generic Diophantine set $\{\vec{x}\mid\varphi(\vec{x},\vec{a})\}$ may be checked by an algorithm knowing only the codes of $\varphi$ and $\vec{a}$ (note that the set of Diophantine formulae is decidable). For example, the Diophantine problem $\mathcal{D}(\mathbb{Z})$ (for the language of rings and a natural enumeration) is undecidable by famous Matiyasevich’s theorem. On the other hand, $\mathcal{D}(\mathbb{Q}^{\mathrm{alg}})$ is decidable (again for the language of rings and a natural enumeration) since every algebraically closed field admits quantifier elimination.
Now let $K$ be a commutative ring and $G$ be a reductive group scheme over $K$ in the sense of [3]. If $G$ is splits, i.e. it is a Chevalley–Demazure group scheme, simple, and of rank at least $2$, then by the main result of Elena Bunina, Alexey Myasnikov, and Eugene Plotkin [2] the Diophantine problems $\mathcal{D}(K)$ and $\mathcal{D}(G(K))$ are equivalent. More precisely, if $K$ is countable with fixed enumeration, then these two problems are reducible to each other by explicit algorithms.
We are going to generalize this result to isotropic $G$, not necessarily split. There are several possible definitions of such group schemes, for example, in Victor Petrov and Anastasia Stavrova’s paper [5] the term isotropic means that there exists a strictly proper parabolic subgroup. There is also a more general notion of locally isotropic reductive group schemes [9], namely, that the isotropicity condition holds locally in the Zariski topology. In this paper we impose an even stronger condition than Petrov and Stavrova, but for local rings all these definitions coincide. We plan to cover the locally isotropic case in a sequel paper using localization technique from [9]. Of course, we actually need not only that $G$ is isotropic, but also that its suitable “isotropic rank” is at least $2$.
We follow the general strategy of Bunina – Myasnikov – Plotkin. Namely, root subgroups of $G(K)$ turn out to be Diophantine, see theorem 4 below. This allows us to construct e-interpretation of the base ring $K$ in the group $G(K)$, this is theorem 5. Finally, since $K$ and $G(K)$ e-interpret each other, their Diophantine problems are equivalent (theorem 6). As an application of preparatory technical results we also find the centralizer and the normalizer of all root subgroups together, this is a generalization of [1] to the isotropic case.
For example, it follows that the Diophantine problem for $\mathrm{SO}(E_{8}\perp H\perp H)$ is unsolvable, where $E_{8}$ is the lattice spanned by the root system of type $\mathsf{E}_{8}$ and $H$ is a hyperbolic plane over $\mathbb{Z}$.
The author wants to thank Eugene Plotkin for the problem statement and great motivation.
2 Isotropic reductive groups
Let $K$ be a commutative ring (all rings in this paper are unital) and $G$ a reductive group scheme over $K$ [3, Exp. XIX]. By root systems we mean crystallographic root systems possibly of type $\mathsf{BC}_{\ell}$. Recall that there is an étale extension $K\subseteq\widetilde{K}$ such that $G_{\widetilde{K}}$ splits with possibly non-constant root datum, i.e. there is a decomposition $\widetilde{K}=\widetilde{K}_{1}\times\ldots\times\widetilde{K}_{n}$ such that each $G_{\widetilde{K}_{i}}$ is isomorphic to a split reductive group scheme $\mathrm{G}(X^{\vee},\widetilde{\Phi}^{\vee},X,\widetilde{\Phi};\widetilde{K}_{i})$ over $\widetilde{K}_{i}$ [3, corollary XXII.2.3]. We always assume that $n=1$, i.e. the root datum $(X^{\vee},\widetilde{\Phi}^{\vee},X,\widetilde{\Phi})$ is constant over $\widetilde{K}$, and that the root system $\widetilde{\Phi}$ is irreducible. But $G$ is not necessarily semisimple, i.e. the scheme center $\operatorname{C}(G)$ may contain a non-trivial torus (such generality is required e.g. in applications of lemma 5 below).
In this paper we call $G$ isotropic with root system $\Phi$ if there are subgroups $L\leq G$ and $G_{\alpha}\leq G$ for $\alpha\in\Phi$ such that after some (each sufficiently large) étale extension $K\subseteq\widetilde{K}$ there are an isomorphism
$F\colon G_{\widetilde{K}}\to\mathrm{G}(X^{\vee},\widetilde{\Phi}^{\vee},X,\widetilde{\Phi};\widetilde{K})$
with a split reductive group scheme and a map $u\colon\widetilde{\Phi}\to\Phi\sqcup\{0\}$ induced by a linear map of the ambient vector spaces such that
  • $\Phi$ is contained in the image of $u$;
  • $F(G_{\alpha,\widetilde{K}})=\langle U_{\beta}\mid u(\beta)\in\{\alpha,2\alpha\}\rangle$ is a standard unipotent subgroup, where $U_{\beta}$ are the root subgroups of the split reductive group scheme;
  • $F(L_{\widetilde{K}})=\mathrm{G}(X^{\vee},\widetilde{\Phi}^{\vee},X,\widetilde{\Phi};\widetilde{K})^{0}_{u^{-1}(0)}$ is a reductive subsystem subgroup, namely, the subgroup of type (RC) corresponding to the subset $u^{-1}(0)$ [3, §XXII.5.11];
  • the group $G$ together with the family $(G_{\alpha})_{\alpha\in\Phi}$ is $\Phi$-graded, i.e. $G_{2\alpha}\leq G_{\alpha}$ for ultrashort $\alpha$,
    $[G_{\alpha},G_{\beta}]\leq\langle G_{\gamma}\mid\gamma\in(\mathbb{N}_{+}\alpha+\mathbb{N}_{+}\beta)\cap\Phi\rangle$
    (this automatically follows from the previous condition), and there are Weyl elements
    $n_{\alpha}\in G_{\alpha}(K)\,G_{-\alpha}(K)\,G_{\alpha}(K)$
    for all $\alpha\in\Phi$, i.e. elements with the property ${{}^{n_{\alpha}}\!{G_{\beta}}}=G_{s_{\alpha}(\beta)}$;
  • The map $u\colon\widetilde{\Phi}\to\Phi\sqcup\{0\}$ comes from one of irreducible Tits indices.
For example, all reductive groups over semi-local rings (and even LG-rings [4]) with connected spectra are naturally isotropic using their minimal parabolic subgroups by [3, Exp. XXVI] and [6]. All these isotropic structures are conjugate by elements of $G(K)$.
Recall [7] that an irreducible Tits index $(\widetilde{\Phi},\Gamma,J)$ consists of a reduced irreducible root system $\widetilde{\Phi}$, a subgroup $\Gamma$ of its group of outer automorphisms (i.e. the automorphism group of its Dynkin diagram), and a $\Gamma$-invariant subset $J$ of vertices of the Dynkin diagram satisfying an additional condition (namely, that it may be constructed by a reductive group scheme over a field using its minimal parabolic subgroup). Then $\Phi$ is the image of $\widetilde{\Phi}$ in the factor-space of the ambient vector space by $\Gamma$ and the span of basic roots not in $J$, the rank of $\Phi$ is the number of $\Gamma$-orbits in $J$. An index is usually denoted as ${{}^{g}\!{\mathsf{X}_{n,r}^{t}}}$, where $g=|\Gamma|$, $\mathsf{X}_{n}$ is the type of $\widetilde{\Phi}$, $r$ is the rank of $\Phi$, and $t$ is an additional parameter (see the list below).
Here is the list of irreducible Tits indices using the standard (Bourbaki) numeration of simple roots.
  • ${{}^{1}\!{\mathsf{A}_{n,r}^{(d)}}}$, where $d(r+1)=n+1$ and $n\geq 1$. Here $J=\{d,2d,\ldots,rd\}$, $\Phi=\mathsf{A}_{r}$, modulo the center $G(K)=\mathrm{PGL}(r+1,A)$ for an Azumaya algebra $A$ over $K$ of rank $d$. Also, $u^{-1}(0)=(r+1)A_{d-1}$ and $|u^{-1}(\alpha)|=d^{2}$ for every root $\alpha$.
  • ${{}^{2}\!{\mathsf{A}_{n,r}^{(d)}}}$, where $d\mid n+1$, $n\geq 1$, and $12rd\leq n+$ 1. Here
    $J=\{d,2d,\ldots,rd;n+1-rd,\ldots,n+1-2d,n+1-d\}.$
    The relative root system is $\Phi=\mathsf{C}_{r}$ for $12rd=n+$ 1 and $\Phi=\mathsf{BC}_{r}$ otherwise. Modulo the center $G(K)=\mathrm{SU}(A,h)$ for an Azumaya algebra $A$ of rank $d$ over a quadratic étale extension $K\subseteq K^{\prime}$ and for a non-degenerate hermitian form $h$ over $A$ of rank $\frac{n+1}{d}$ and Witt index $r$ (i.e. with a chosen hyperbolic subspace of rank $2r$). Also, $u^{-1}(0)=2r\mathsf{A}_{d-1}+\mathsf{A}_{n-2rd}$ (the last summand is vacuous for $2rd\in\{n,n+1\}$), $|u^{-1}(\alpha)|=d^{2}$ for long $\alpha$, $|u^{-1}(\alpha)|=2d^{2}$ for short $\alpha$, $|u^{-1}(\alpha)|=2d(n+1-2rd)$ for ultrashort $\alpha$ (and $2rd\leq n$).
  • $\mathsf{B}_{n,r}$, where $n\geq 2$ and $r\leq n$. Here $J=\{1,2,\ldots,r\}$, $\Phi=\mathsf{B}_{r}$, and $G(K)=\mathrm{SO}(K,q)$ modulo the center, where $q$ is a semi-regular quadratic form of rank $12n+$ 1 and Witt index $r$. Also, $u^{-1}(0)=\mathsf{B}_{n-r}$, $|u^{-1}(\alpha)|=1$ for long $\alpha$, and $|u^{-1}(\alpha)|=2n+1-2r$ for short $\alpha$.
  • $\mathsf{C}_{n,r}^{(d)}$, where $n\geq 3$, $d=2^{k}\mid 2n$, $rd\leq n$, and $n=r$ in the case $d=1$. Here $J=\{d,2d,\ldots,rd\}$, $\Phi=\mathsf{C}_{r}$ for $rd=n$ and $\Phi=\mathsf{BC}_{r}$ otherwise. Modulo the center $G(K)=\mathrm{U}(A,h)$, where $A$ is an Azumaya algebra over $K$ of rank $d$ with symplectic involution, $h$ is a non-degenerate anti-hermitian form of rank $\frac{2n}{d}$ and Witt index $r$ (with the additional condition that $h$ comes from a symplectic form, not an arbitrary alternating form). Also, $u^{-1}(0)=r\mathsf{A}_{d-1}+\mathsf{C}_{n-rd}$, $|u^{-1}(\alpha)|=\frac{d(d+1)}{2}$ for long $\alpha$, $|u^{-1}(\alpha)|=d^{2}$ for short $\alpha$, $|u^{-1}(\alpha)|=2d(n-rd)$ for ultrashort $\alpha$ (and $rd<n$).
  • ${{}^{1}\!{\mathsf{D}_{n,r}^{(d)}}}$, where $n\geq 4$, $d=2^{k}\mid 2n$, $rd\leq n$, and $n\neq rd+1$. Here $J=\{d,2d,\ldots,rd\}$, $\Phi=\mathsf{D}_{r}$ for $rd=n$ and $d=1$, $\Phi=\mathsf{C}_{r}$ for $rd=n$ and $d>1$, $\Phi=\mathsf{B}_{r}$ for $rd<n$ and $d=1$, and $\Phi=\mathsf{BC}_{r}$ for $rd<n$ and $d>1$. Modulo the center $G(K)=\mathrm{SU}(A,h,q)$, where $A$ is an Azumaya algebra over $K$ of rank $d$ with orthogonal involution, $h$ is a non-degenerate hermitian form of rank $\frac{2n}{d}$ and Witt index $r$, and $q$ is a suitable associated quadratic form. Also, $u^{-1}(0)=r\mathsf{A}_{d-1}+\mathsf{D}_{n-rd}$, $|u^{-1}(\alpha)|=\frac{d(d-1)}{2}$ for long $\alpha$ (and $d>1$), $|u^{-1}(\alpha)|=d^{2}$ for short $\alpha$, $|u^{-1}(\alpha)|=2d(n-rd)$ for ultrashort $\alpha$ (and $rd<n$). Here roots of the relative root systems of types $\mathsf{D}_{r}$ and $\mathsf{B}_{r}$ are called short and ultrashort depending on their length, not long and short.
  • ${{}^{2}\!{\mathsf{D}_{n,r}^{(d)}}}$, where $n\geq 4$, $d=2^{k}\mid 2n$, and $rd\leq n-1$. Here $J=\{d,2d,\ldots,rd\}$ for $rd<n-1$ and $J=\{d,2d,\ldots,n-1,n\}$ otherwise, $\Phi=\mathsf{BC}_{r}$ for $d>1$ and $\Phi=\mathsf{B}_{r}$ otherwise. Modulo the center $G(K)=\mathrm{SU}(A,h,q)$, where $A$ is an Azumaya algebra over $K$ of rank $d$ with orthogonal involution, $h$ is a non-degenerate hermitian form of rank $\frac{2n}{d}$ and Witt index $r$, and $q$ is a suitable associated quadratic form. Also, $u^{-1}(0)=r\mathsf{A}_{d-1}+\mathsf{D}_{n-rd}$, $|u^{-1}(\alpha)|=\frac{d(d-1)}{2}$ for long $\alpha$ (and $d>1$), $|u^{-1}(\alpha)|=d^{2}$ for short $\alpha$, $|u^{-1}(\alpha)|=2d(n-rd)$ for ultrashort $\alpha$. As in the previous case, roots of the relative root system of type $\mathsf{B}_{r}$ are called short and ultrashort.
Exceptional Tits indices are given in the following table. Sizes and types of preimages are given by increasing the root length. The indices of labels in $J$ denote the lengths of the corresponding roots in $\Phi$, namely, long, short, and ultrashort ones.
Tits index $\Phi$ $J$ $u^{-1}(0)$ $|u^{-1}(\alpha)|$ $u^{-1}(\mathbb{Z}\alpha\cap(\Phi\cup\{0\}))$
${{}^{3}\!{\mathsf{D}_{4,0}^{28}}}$, ${{}^{6}\!{\mathsf{D}_{4,0}^{28}}}$ 00 $\varnothing$ $\mathsf{D}_{4}$
${{}^{1}\!{\mathsf{E}_{6,0}^{78}}}$, ${{}^{2}\!{\mathsf{E}_{6,0}^{78}}}$ 00 $\varnothing$ $\mathsf{E}_{6}$
$\mathsf{E}_{7,0}^{133}$ 00 $\varnothing$ $\mathsf{E}_{7}$
$\mathsf{E}_{8,0}^{248}$ 00 $\varnothing$ $\mathsf{E}_{8}$
$\mathsf{F}_{4,0}^{52}$ 00 $\varnothing$ $\mathsf{F}_{4}$
$\mathsf{G}_{2,0}^{14}$ 00 $\varnothing$ $\mathsf{G}_{2}$
$\mathsf{E}_{7,1}^{78}$ $\mathsf{A}_{1}$$\{7\}$ $\mathsf{E}_{6}$ $27$ $\mathsf{E}_{7}$
${{}^{3}\!{\mathsf{D}_{4,1}^{9}}}$, ${{}^{6}\!{\mathsf{D}_{4,1}^{9}}}$ $\mathsf{BC}_{1}$$\{2\}$ $3\mathsf{A}_{1}$ $8$, $1$ $\mathsf{D}_{4}$, $4\mathsf{A}_{1}$
${{}^{2}\!{\mathsf{E}_{6,1}^{35}}}$ $\mathsf{BC}_{1}$$\{2\}$ $\mathsf{A}_{5}$ $20$, $1$ $\mathsf{E}_{6}$, $\mathsf{A}_{1}+\mathsf{A}_{5}$
$\mathsf{E}_{7,1}^{66}$ $\mathsf{BC}_{1}$$\{1\}$ $\mathsf{D}_{6}$ $32$, $1$ $\mathsf{E}_{7}$, $\mathsf{A}_{1}+\mathsf{D}_{6}$
$\mathsf{E}_{8,1}^{133}$ $\mathsf{BC}_{1}$$\{8\}$ $\mathsf{E}_{7}$ $56$, $1$ $\mathsf{E}_{8}$, $\mathsf{A}_{1}+\mathsf{E}_{7}$
$\mathsf{F}_{4,1}^{21}$ $\mathsf{BC}_{1}$$\{4\}$ $\mathsf{B}_{3}$ $8$, $7$ $\mathsf{F}_{4}$, $\mathsf{B}_{4}$
${{}^{2}\!{\mathsf{E}_{6,1}^{29}}}$ $\mathsf{BC}_{1}$$\{1,6\}$ $\mathsf{D}_{4}$ $16$, $8$ $\mathsf{E}_{6}$, $\mathsf{D}_{5}$
$\mathsf{E}_{7,1}^{48}$ $\mathsf{BC}_{1}$$\{6\}$ $\mathsf{A}_{1}+\mathsf{D}_{5}$ $32$, $10$ $\mathsf{E}_{7}$, $\mathsf{D}_{6}$
$\mathsf{E}_{8,1}^{91}$ $\mathsf{BC}_{1}$$\{1\}$ $\mathsf{D}_{7}$ $64$, $14$ $\mathsf{E}_{8}$, $\mathsf{D}_{8}$
${{}^{1}\!{\mathsf{E}_{6,2}^{28}}}$ $\mathsf{A}_{2}$$\{1,6\}$ $\mathsf{D}_{4}$ $8$ $\mathsf{D}_{5}$
$\mathsf{G}_{2,2}^{0}$ $\mathsf{G}_{2}$$\{1_{\mathrm{s}},2_{\mathrm{l}}\}$varnothing $1$, $1$ $\mathsf{A}_{1}$, $\mathsf{A}_{1}$
${{}^{3}\!{\mathsf{D}_{4,2}^{2}}}$, ${{}^{6}\!{\mathsf{D}_{4,2}^{2}}}$ $\mathsf{G}_{2}$$\{1_{\mathrm{s}},2_{\mathrm{l}},3_{\mathrm{s}},4_{\mathrm{s}}\}$varnothing $3$, $1$ $3\mathsf{A}_{1}$, $\mathsf{A}_{1}$
${{}^{1}\!{\mathsf{E}_{6,2}^{16}}}$, ${{}^{2}\!{\mathsf{E}_{6,2}^{16^{\prime\prime}}}}$ $\mathsf{G}_{2}$$\{2_{\mathrm{l}},4_{\mathrm{s}}\}$ $2\mathsf{A}_{2}$ $9$, $1$ $\mathsf{A}_{5}$, $\mathsf{A}_{1}+2\mathsf{A}_{2}$
$\mathsf{E}_{8,2}^{72}$ $\mathsf{G}_{2}$$\{7_{\mathrm{s}},8_{\mathrm{l}}\}$ $\mathsf{E}_{6}$ $27$, $1$ $\mathsf{E}_{7}$, $\mathsf{A}_{1}+\mathsf{E}_{6}$
${{}^{2}\!{\mathsf{E}_{6,2}^{16^{\prime}}}}$ $\mathsf{BC}_{2}$$\{1_{\mathrm{us}},2_{\mathrm{s}},6_{\mathrm{us}}\}$ $\mathsf{A}_{3}$ $8$, $6$, $1$ $\mathsf{A}_{5}$, $\mathsf{D}_{4}$, $2\mathsf{A}_{1}+\mathsf{A}_{3}$
$\mathsf{E}_{7,2}^{31}$ $\mathsf{BC}_{2}$$\{1_{\mathrm{s}},6_{\mathrm{us}}\}$ $\mathsf{A}_{1}+\mathsf{D}_{4}$ $16$, $8$, $1$ $\mathsf{D}_{6}$, $\mathsf{A}_{1}+\mathsf{D}_{5}$, $2\mathsf{A}_{1}+\mathsf{D}_{4}$
$\mathsf{E}_{8,2}^{66}$ $\mathsf{BC}_{2}$$\{1_{\mathrm{us}},8_{\mathrm{s}}\}$ $\mathsf{D}_{6}$ $32$, $12$, $1$ $\mathsf{E}_{7}$, $\mathsf{D}_{7}$, $\mathsf{A}_{1}+\mathsf{D}_{6}$
$\mathsf{E}_{7,3}^{28}$ $\mathsf{C}_{3}$$\{1_{\mathrm{s}},6_{\mathrm{s}},7_{\mathrm{l}}\}$ $\mathsf{D}_{4}$ $8$, $1$ $\mathsf{D}_{5}$, $\mathsf{A}_{1}+\mathsf{D}_{4}$
$\mathsf{F}_{4,4}^{0}$ $\mathsf{F}_{4}$$\{1_{\mathrm{l}},2_{\mathrm{l}},3_{\mathrm{s}},4_{\mathrm{s}}\}$varnothing $1$, $1$ $\mathsf{A}_{1}$, $\mathsf{A}_{1}$
${{}^{2}\!{\mathsf{E}_{6,4}^{2}}}$ $\mathsf{F}_{4}$$\{1_{\mathrm{s}},2_{\mathrm{l}},3_{\mathrm{s}},4_{\mathrm{l}},5_{\mathrm{s}},6_{\mathrm{s}}\}$varnothing $2$, $1$ $2\mathsf{A}_{1}$, $\mathsf{A}_{1}$
$\mathsf{E}_{7,4}^{9}$ $\mathsf{F}_{4}$$\{1_{\mathrm{l}},3_{\mathrm{l}},4_{\mathrm{s}},6_{\mathrm{s}}\}$ $3\mathsf{A}_{1}$ $4$, $1$ $\mathsf{A}_{1}+\mathsf{A}_{3}$, $4\mathsf{A}_{1}$
$\mathsf{E}_{8,4}^{28}$ $\mathsf{F}_{4}$$\{1_{\mathrm{s}},6_{\mathrm{s}},7_{\mathrm{l}},8_{\mathrm{l}}\}$ $\mathsf{D}_{4}$ $8$, $1$ $\mathsf{D}_{5}$, $\mathsf{A}_{1}+\mathsf{D}_{4}$
${{}^{1}\!{\mathsf{E}_{6,6}^{0}}}$ $\mathsf{E}_{6}$$\{1,2,3,4,5,6\}$varnothing $1$ $\mathsf{A}_{1}$
$\mathsf{E}_{7,7}^{0}$ $\mathsf{E}_{7}$$\{1,2,3,4,5,6,7\}$varnothing $1$ $\mathsf{A}_{1}$
$\mathsf{E}_{8,8}^{0}$ $\mathsf{E}_{8}$$\{1,2,3,4,5,6,7,8\}$varnothing $1$ $\mathsf{A}_{1}$
3 Technical lemmas
Recall that a subset $\Sigma\subseteq\Phi$ is called closed if $(\Sigma+\Sigma)\cap\Phi\subseteq\Sigma$. A closed subset $\Sigma$ is
  • unipotent if it is contained in an open half-space (equivalently, if it does not contain opposite roots);
  • closed root subsystem if $\Sigma=-\Sigma$, so it is a root system itself;
  • parabolic if $\Phi=\Sigma\cup(-\Sigma)$;
  • saturated if $\Sigma=\Phi\cap\mathbb{R}_{\geq 0}\Sigma$.
Note that any parabolic set is saturated (and every saturated set is an intersection of parabolic ones), but $\mathsf{A}_{2}\subseteq\mathsf{G}_{2}$ is not saturated. Any closed subset $\Sigma$ admits a unique decomposition $\Sigma=\Sigma_{\mathrm{r}}\sqcup\Sigma_{\mathrm{u}}$, where $\Sigma_{\mathrm{r}}=\Sigma\cap(-\Sigma)$ is a closed root subsystem and $\Sigma_{\mathrm{u}}=\Sigma\setminus(-\Sigma)$ is a unipotent set. Conversely, if $\Sigma_{\mathrm{r}}$ is a closed root subsystem, $\Sigma_{\mathrm{u}}$ is a unipotent set disjoint with $\Sigma_{\mathrm{r}}$, and $(\Sigma_{\mathrm{r}}+\Sigma_{\mathrm{u}})\cap\Phi\subseteq\Sigma_{\mathrm{u}}$, then $\Sigma_{\mathrm{r}}\cup\Sigma_{\mathrm{u}}$ is a closed set. A closed set $\Sigma$ is parabolic if and only if there are $w\in\mathrm{W}(\Phi)$ and $J\subseteq\Delta$ such that $w\Sigma=(\Pi+\mathbb{Z}J)\cap\Phi$. Here $\Pi\subseteq\Phi$ is the set of positive roots and $\Delta\subseteq\Pi$ is the set of basic roots. The classes of closed sets, unipotent sets, closed root subsystems, and saturated sets are closed under intersection.
If $\Sigma\subseteq\Phi$ is closed, then $u^{-1}(\Sigma\cup\{0\})\subseteq\widetilde{\Phi}$ is also closed. More precisely,
  • if $\Sigma$ is unipotent, then $u^{-1}(\Sigma)$ is unipotent;
  • if $\Sigma$ is a closed root subsystem, then $u^{-1}(\Sigma\cup\{0\})$ is a closed root subsystem;
  • if $\Sigma$ is parabolic, then $u^{-1}(\Sigma\cup\{0\})$ is parabolic;
  • if $\Sigma$ is saturated, then $u^{-1}(\Sigma\cup\{0\})$ is saturated.
If $\Sigma\subseteq\Phi$ is a closed root subsystem, then the smallest closed set containing $u^{-1}(\Sigma)$ consists of $u^{-1}(\Sigma)$ and sums of pairs of roots of $u^{-1}(\Sigma)$ (with opposite images in $\Sigma$), it is also a closed root subsystem.
For any closed subset $\Sigma\subseteq\Phi$ let $G_{\Sigma}\leq G$ and $G^{0}_{\Sigma}\leq G$ be the group subsheaves generated by $\bigcup_{\alpha\in\Sigma}U_{\alpha}$ and $L\cup\bigcup_{\alpha\in\Sigma}U_{\alpha}$, they are actually smooth closed group subschemes with connected geometric fibers over $K$ and $G^{0}_{\Sigma}$ is of type (RC) [3, §XXII.5]. More precisely, if $\Sigma$ is a root subsystem, then $G_{\Sigma}$ and $G^{0}_{\Sigma}$ are reductive group subschemes (with induced isotropic structure). If $\Sigma$ is unipotent, then $G_{\Sigma}$ is a unipotent subgroup, i.e. it is an étale twisted form of some $\mathbb{A}^{N}$ as a scheme. Finally, in general $G_{\Sigma}=G_{\Sigma_{\mathrm{r}}}\rtimes G_{\Sigma_{\mathrm{u}}}$, and $G^{0}_{\Sigma}=G^{0}_{\Sigma_{\mathrm{r}}}\rtimes G_{\Sigma_{\mathrm{u}}}$. If $\Sigma$ is parabolic, then $G^{0}_{\Sigma}$ is a parabolic group subscheme. For any parabolic $\Sigma\subseteq\Phi$ the multiplication morphism
$G_{-\Sigma_{\mathrm{u}}}\times G^{0}_{\Sigma_{\mathrm{r}}}\times G_{\Sigma_{\mathrm{u}}}\to G$
is an open scheme embedding (so for any closed $\Sigma$ such a morphism is just a scheme embedding). Also, for any unipotent $\Sigma$ the multiplication morphism $\prod_{\alpha\in\Sigma\setminus 2\Sigma}G_{\alpha}\to G_{\Sigma}$ is an isomorphism of schemes for any order of the factors.
Lemma 1.
Suppose that $K$ is semi-local. Then every isotropic reductive group $G$ has a Gauss decomposition
$G(K)=G_{\Pi}(K)\,G_{-\Pi}(K)\,G_{\Pi}(K)\,L(K).$
Moreover, for any closed subset $\Sigma\subseteq\Phi$ there is a decomposition
$G^{0}_{\Sigma}(K)=G_{\Sigma_{\mathrm{r}}\cap\Pi}(K)\,G_{\Sigma\cap(-\Pi)}(K)\,G_{\Sigma\cap\Pi}(K)\,L(K).$
Proof.
We may assume that $\operatorname{Spec}(K)$ is connected. In this case there is a maximal isotropic structure $(L^{\prime},G^{\prime}_{\alpha})_{\alpha\in\Phi^{\prime}}$ and a map $v\colon\Phi^{\prime}\to\Phi\cup\{0\}$ such that $\Phi\subseteq v(\Phi^{\prime})$, $G^{\prime}_{\alpha}=G_{v^{-1}(\alpha)}$, $L^{\prime}=G_{\Phi\cap v^{-1}(0)}$ [3, XXVI.7.4.2]. Gauss decomposition holds for $G$ with respect to a maximal isotropic structure by [3, corollary XXVI.5.2], i.e.
$G(K)=G^{\prime}_{\Pi^{\prime}}(K)\,G^{\prime}_{-\Pi^{\prime}}(K)\,G^{\prime}_{\Pi^{\prime}}(K)\,L^{\prime}(K).$
It follows that
$G(K)=G_{\Pi}(K)\,G_{-\Pi}(K)\,G_{\Pi}(K)\,L(K)$
and
$G^{0}_{\Sigma}(K)=G_{\Sigma\cap\Pi}(K)\,G_{\Sigma\cap(-\Pi)}(K)\,G_{\Sigma\cap\Pi}(K)\,L(K)$
for any root subsystem $\Sigma$. If $\Sigma\subseteq\Phi$ is a closed subset, then
$G^{0}_{\Sigma}(K)$ $=G_{\Sigma_{\mathrm{r}}\cap\Pi}(K)\,G_{\Sigma_{\mathrm{r}}\cap(-\Pi)}(K)\,G_{\Sigma_{\mathrm{u}}}(K)\,G_{\Sigma_{\mathrm{r}}\cap\Pi}(K)\,L(K)$
$=G_{\Sigma_{\mathrm{r}}\cap\Pi}(K)\,G_{\Sigma\cap(-\Pi)}(K)\,G_{\Sigma\cap\Pi}(K)\,L(K).\qed$
For any closed subsets $\Sigma,\Sigma^{\prime}\subseteq\Phi$ the intersection $G^{0}_{\Sigma}\cap G^{0}_{\Sigma^{\prime}}$ is a smooth closed subscheme and its fiberwise connected component is $G^{0}_{\Sigma\cap\Sigma^{\prime}}$ by [3, proposition XXII.5.4.5].
Lemma 2.
Let $\Sigma,\Sigma^{\prime}\subseteq\Phi$ be closed subsets. If $\Sigma^{\prime}$ is saturated, then $G^{0}_{\Sigma}\cap G^{0}_{\Sigma^{\prime}}=G^{0}_{\Sigma\cap\Sigma^{\prime}}$. If $\Sigma^{\prime}$ is unipotent, then $G^{0}_{\Sigma}\cap G_{\Sigma^{\prime}}=G_{\Sigma}\cap G_{\Sigma^{\prime}}=G_{\Sigma\cap\Sigma^{\prime}}$.
Proof.
For the first claim it suffices to prove that $G^{0}_{\Sigma}\cap G^{0}_{\Sigma^{\prime}}=G^{0}_{\Sigma\cap\Sigma^{\prime}}$ if $\Sigma^{\prime}\supseteq\Pi$ is parabolic and $K$ is local. Take $g\in G^{0}_{\Sigma}(K)\cap G^{0}_{\Sigma^{\prime}}(K)$. By lemma 1 we may multiply $g$ from both sides by elements of $G^{0}_{\Sigma\cap\Sigma^{\prime}}$ to get a new element $g^{\prime}\in G_{\Sigma\cap(-\Sigma^{\prime}_{\mathrm{u}})}(K)\cap G^{0}_{\Sigma^{\prime}}(K)$. But such $g^{\prime}$ is necessarily trivial.
Now let us prove the second claim. Without loss of generality $\Sigma^{\prime}\subseteq\Pi$, so
$G^{0}_{\Sigma}\cap G_{\Sigma^{\prime}}=G^{0}_{\Sigma}\cap G^{0}_{\Pi}\cap G_{\Sigma^{\prime}}=G^{0}_{\Sigma\cap\Pi}\cap G_{\Sigma^{\prime}}=G_{\Sigma\cap\Pi\cap\Sigma^{\prime}}=G_{\Sigma\cap\Sigma^{\prime}}.\qed$
Recall [8, §4] that a $2$-step $K$-module $(M,M_{0})$ consists of
[…]
  • a central subgroup $M_{0}\leq M$ such that $[M,M]^{\cdot}\leq M_{0}$;
  • a left $K$-module structure on $M_{0}$;
  • a right action $({-})\cdot({=})\colon M\times K\to M$ of the multiplicative monoid by group automorphisms such that $[m\cdot k,m^{\prime}\cdot k^{\prime}]^{\cdot}=kk^{\prime}[m,m^{\prime}]^{\cdot}$, $m\cdot(k+k^{\prime})=m\cdot k\dotplus kk^{\prime}\tau(m)\dotplus m\cdot k^{\prime}$ for some (uniquely determined) $\tau(m)\in M_{0}$, $m_{0}\cdot k=k^{2}m_{0}$ for $m_{0}\in M_{0}$.
Then
$m\cdot 0$ $=\dot{0},$ $m\cdot(-k)$ $=k^{2}\tau(m)\mathbin{\dot{-}}m\cdot k,$
$\tau(\dot{0})$ $=\dot{0},$ $\tau(m\dotplus m^{\prime})$ $=\tau(m)+[m,m^{\prime}]^{\cdot}+\tau(m^{\prime}),$
$\tau(\mathbin{\dot{-}}m)$ =˙$=\mathbin{\dot{-}}\tau(m),$ $\tau(m\cdot k)$ $=k^{2}\tau(m),$
$\tau(m_{0})$ $=2m_{0}\text{ for }m_{0}\in M_{0}.$
A pair of subsets $(X,X_{0})\subseteq(M,M_{0})$ generates $(M,M_{0})$ if $X_{0}$ generates the $K$-module $M_{0}$ and $X$ generates the $K$-module $M/M_{0}$.
We say that a $2$-step nilpotent $K$-module $(M,M_{0})$ is locally free if $M_{0}$ and $M/M_{0}$ are finitely generated projective $K$-modules. In this case $(M,M_{0})$ splits, i.e. $M=M_{0}\mathbin{\dot{\oplus}}M_{1}$ for some $K$-module $M_{1}\cong M/M_{0}$ with bilinear map $c\colon M_{1}\times M_{1}\to M_{0}$ and the operations are given by
$(m_{0}\mathbin{\dot{\oplus}}m_{1})\dotplus(m^{\prime}_{0}\mathbin{\dot{\oplus}}m^{\prime}_{1})$ $=(m_{0}+c(m_{1},m^{\prime}_{1})+m^{\prime}_{0})\mathbin{\dot{\oplus}}(m_{1}+m^{\prime}_{1}),$
$(m_{0}\mathbin{\dot{\oplus}}m_{1})\cdot k$ $=k^{2}m_{0}\mathbin{\dot{\oplus}}km_{1},$
$\tau(m_{0}\mathbin{\dot{\oplus}}m_{1})$ $=(2m_{0}-c(m_{1},m_{1}))\mathbin{\dot{\oplus}}0.$
Every locally free $2$-step nilpotent $K$-module $(M,M_{0})$ determines a representable fpqc sheaf $\mathbb{M}_{0}(E)=E\otimes_{K}M_{0}$, $\mathbb{M}(E)=M\boxtimes E=E\otimes_{K}M_{0}\mathbin{\dot{\oplus}}E\otimes_{K}M_{1}$ of locally free $2$-step nilpotent modules (where $K\to E$ is a ring homomorphism), the latter one is independent on the choice of the splitting [8, §4]. Also, locally free $2$-step nilpotent $K$-modules satisfy the fpqc descent. If $G$ is a group scheme acting on such a sheaf $(\mathbb{M},\mathbb{M}_{0})$ by automorphisms of $2$-step nilpotent modules and $G$ stabilizes some generating set $(X,X_{0})\subseteq(M,M_{0})$, then the action is trivial.
By [9, §2] there are canonical homomorphisms $t_{\alpha}\colon\mathfrak{g}_{\alpha}\to G_{\alpha}(K)$ for non-ultrashort $\alpha\in\Phi$ and $t_{\alpha}\colon\mathfrak{g}_{\alpha}\mathbin{\dot{\oplus}}\mathfrak{g}_{2\alpha}\to G_{\alpha}(K)$ for ultrashort $\alpha$. In the second case $\mathfrak{g}_{\alpha}\dotplus\mathfrak{g}_{2\alpha}$ is a locally free $2$-step nilpotent $K$-module. Since $G$ is root graded, the Lie bracket $\mathfrak{g}_{\alpha}\times\mathfrak{g}_{\beta}\to\mathfrak{g}_{\alpha+\beta}$ is non-degenerate on the second argument if $\alpha$ is long and $\frac{\pi}{2}<\angle(\alpha,\beta)<\pi$, i.e. $\mathfrak{g}_{\beta}\to\mathrm{Hom}(\mathfrak{g}_{\alpha},\mathfrak{g}_{\alpha+\beta})$ is injective. There are also several other instances of non-degeneracy, we check them case by case.
Lemma 3.
Suppose that $\Phi$ is of type $\mathsf{C}_{2}$. Then for any long $\alpha$ and short $\beta$ such that $\angle(\alpha,\beta)=\frac{3\pi}{4}$ the Lie bracket $\mathfrak{g}_{\alpha}\times\mathfrak{g}_{\beta}\to\mathfrak{g}_{\alpha+\beta}$ is non-degenerate.
Proof.
In the orthogonal-like case $\mathfrak{g}_{\alpha}\cong R$ for a ring $R$ (an Azumaya algebra over $K$ or its quadratic étale extension) and $\mathfrak{g}_{\beta}\cong M_{R}$ for some right faithfully projective $A$-module, the pairing is just $M\times R\to M,\,(m,r)\mapsto mr$ up to an isomorphism. In the symplectic-like case $\mathfrak{g}_{\beta}\cong R$ for a ring $R$ with a $\lambda$-involution and $\mathfrak{g}_{\alpha}\cong\Lambda$ for a form parameter $\Lambda\leq R$, the pairing is $R\times\Lambda\to R,\,(r,u)\mapsto ru$ up to an isomorphism. This pairing is also non-degenerate, the intersection $\Lambda\cap R^{*}$ is non-empty because $G$ is root graded. ∎
Lemma 4.
Suppose that $\Phi$ is of type $\mathsf{BC}_{2}$ and $\alpha,\beta\in\Phi$ are orthogonal ultrashort roots. Then the Lie bracket $\mathfrak{g}_{\alpha}\times\mathfrak{g}_{\beta}\to\mathfrak{g}_{\alpha+\beta}$ is non-degenerate.
Proof.
If $G$ is classical, then this follows from the non-degeneracy of the hermitian form on the defining representation. Otherwise there are the cases ${{}^{2}\!{\mathrm{E}_{6,2}^{16^{\prime}}}}$, $\mathrm{E}_{7,2}^{31}$, and $\mathsf{E}_{8,2}^{66}$. It is possible to check using the root diagrams of $\mathsf{E}_{\ell}$ that the required Lie bracket is non-degenerate. ∎
4 Some normalizers and centralizers
Theorem 1.
Let $G$ be an isotropic reductive group over $K$ with root system $\Phi$ of rank $\geq 2$. Let also $\alpha\in\Phi$ be a long root and $X\subseteq\mathfrak{g}_{\alpha}$ be a generating set of a $K$-module. Then the group scheme
$N=\{g\in G\mid\forall x\in X\enskip{{}^{g}\!{t_{\alpha}(x)}}\in G_{\alpha}\}$
coincides with the parabolic subgroup $P=G^{0}_{\{\beta\mid\angle(\alpha,\beta)\leq\pi/2\}}$.
Proof.
Clearly, $P\leq N$ and we may assume that $K$ is local. By lemma 1 it suffices to check that $G_{\Sigma}\cap N=1$, where $\Sigma=\{\beta\mid\angle(\alpha,\beta)>\frac{\pi}{2}\}$. Choose a linear map $f\colon\mathbb{R}\Phi\to\mathbb{R}$ such that $\operatorname{Ker}(f)\cap\Phi=\mathbb{R}\alpha\cap\Phi$ and let $\Phi_{s}=\{\beta\in\Phi\mid\mathrm{sign}(f(\beta))=s\}$, $\Sigma_{s}=\Phi_{s}\cap\Sigma$ for $s\in\{{-},0,{+}\}$, so $\Phi_{\pm}$ are unipotent subsets.
Take any element $g=g_{-}g_{0}g_{+}\in G_{\Sigma}\cap N$ with $g_{s}\in G_{\Sigma_{s}}$, then $g_{\pm}=\prod_{\beta\in\Sigma_{\pm}\setminus 2\Sigma_{\pm}}t_{\beta}(y_{\beta})$ for any linear orders on $\Sigma_{\pm}\setminus 2\Sigma_{\pm}$ and $g_{0}\in G_{-\alpha}$ (or $g_{0}\in G_{-\alpha/2}$ if $\Phi$ is of type $\mathsf{BC}_{\ell}$). For any $x\in X$ there is $x^{\prime}\in\mathfrak{g}_{\alpha}$ such that $g\,t_{\alpha}(x)=t_{\alpha}(x^{\prime})\,g$. Comparing both sides of $g_{-}\,g_{0}\,t_{\alpha}(x)\,g_{+}^{t_{\alpha}(x)}={{}^{t_{\alpha}(x^{\prime})}\!{g_{-}}}\,t_{\alpha}(x^{\prime})\,g_{0}\,g_{+}$ we get $[g_{+},t_{\alpha}(x)]=1$ and $[g_{-},t_{\alpha}(x^{\prime})]=1$. Since $x$ and $x^{\prime}$ runs generating sets of $\mathfrak{g}_{\alpha}$ and $G$ is root graded (so the commutator map $\mathfrak{g}_{\alpha}\times\mathfrak{g}_{\beta}\to\mathfrak{g}_{\alpha+\beta}$ is non-degenerate on the second argument for $\beta\in\Sigma_{\pm}$) both $g_{+}$ and $g_{-}$ are trivial.
It follows that $g=g_{0}\in G_{-\alpha^{\prime}}$ for $\alpha^{\prime}\in\{\alpha,\frac{\alpha}{2}\}$. Now suppose that $\Phi$ is not of the type $\mathsf{BC}_{\ell}$ and take a root $\beta\in\Phi$ such that $\frac{\pi}{2}<\angle(\alpha,\beta)<\pi$ and $\beta$ is long if $\Phi$ is of the type $\mathsf{G}_{2}$. Conjugate $t_{\beta}(z)$ by both sides of $g\,t_{\alpha}(x)=t_{\alpha}(x^{\prime})\,g$. We get $\bigl{[}\log g,[x,z]\bigr{]}=1$, where $[{-},{=}]\colon\mathfrak{g}_{\alpha}\times\mathfrak{g}_{\beta}\to\mathfrak{g}_{\alpha+\beta}$ is the commutator map and $\log g\in\mathfrak{g}_{-\alpha}$ is the element such that $g=t_{-\alpha}(\log g)$. Since $G$ is root graded, $\mathfrak{g}_{\alpha+\beta}=[\mathfrak{g}_{\alpha},\mathfrak{g}_{\beta}]$, so $[\log g,\mathfrak{g}_{\alpha+\beta}]=0$ and $\log g=0$ (by lemma 3 if $\beta$ is short), i.e. $g=1$.
Finally, suppose that $\Phi$ is of type $\mathsf{BC}_{\ell}$. Take an ultrashort root $\beta$ orthogonal to $\alpha$ and conjugate $t_{\beta}(u)$ by both sides of $g\,t_{\alpha}(x)=t_{\alpha}(x^{\prime})\,g$. We get $[t_{\alpha}(x^{\prime}),[g,t_{\beta}(u)]]=1$, so $[g,t_{\beta}(u)]=1$ and $g\in G_{-\alpha}$ by lemma 4. In other words, we reduce to the previous case of $\mathsf{C}_{\ell}\subseteq\mathsf{BC}_{\ell}$. ∎
We need another example of locally free $2$-step nilpotent $K$-modules. Let $G$ be an isotropic reductive group scheme over $K$ with root system $\Phi$ and $f\colon\mathbb{R}\Phi\to\mathbb{R}$ be a linear map such that $f(\Phi)\subseteq\{-2,-1,0,1,2\}$, i.e. a “$5$-grading”. Then
$\bigl{(}G_{\Phi\cap f^{-1}(\{1,2\})},G_{\Phi\cap f^{-1}(2)}\bigr{)}$
is a sheaf of locally free $2$-step nilpotent modules. The action of the group scheme $G^{0}_{\Phi\cap\operatorname{Ker}(f)}$ on it commutes with the operations of $2$-step nilpotent modules, this may be easily checked by passing to a split form of $G$.
Lemma 5.
Let $G$ be an isotropic reductive group scheme and $f\colon\mathbb{R}\Phi\to\mathbb{R}$ a linear map such that $f(\Phi)\subseteq\{-2,-1,0,1,2\}$ and $\Phi\cap f^{-1}(2)\neq\varnothing$. Then
$C=\{g\in G^{0}_{\Phi_{0}}\mid[g,t_{\alpha}(x_{\alpha,i})]=1\text{ for }f(\alpha)>0\}$
coincides with the scheme center $\mathrm{C}(G)\leq\mathrm{C}(L)$, where $\Phi_{0}=\Phi\cap\operatorname{Ker}(f)$ and $x_{\alpha,i}\in\mathfrak{g}_{\alpha}$ are $K$-module generators if $\alpha$ is not ultrashort and $x_{\alpha,i}\in\mathfrak{g}_{\alpha}\mathbin{\dot{\oplus}}\mathfrak{g}_{2\alpha}$ generate the $K$-module $\mathfrak{g}_{\alpha}$ otherwise.
Proof.
Clearly, $C$ is the scheme centralizer of $G_{\Phi_{>0}}$ in $G^{0}_{\Phi_{0}}$, where $\Phi_{>0}=\{\alpha\in\Phi\mid f(\alpha)>0\}$. We may assume that $G$ splits and $\Phi=\widetilde{\Phi}$. Firstly, let us show that $C$ is contained in the maximal torus $T=L$. By theorem 1 and lemma 2 the group scheme $C$ is contained in $G^{0}_{\Sigma}$, where $\Sigma$ consists of the roots $\alpha\in\Phi_{0}$ such that $\angle(\alpha,\beta)\leq\frac{\pi}{2}$ for all long roots $\beta\in\Phi_{>0}$. By considering rank $2$ root subsystems it may be shown that roots of $\Sigma$ are actually orthogonal to long roots from $\Phi_{>0}$.
Now if the set of long roots of $\Phi$ is indecomposable root system (of full rank), then the set of long roots from $\Phi\setminus\Phi_{0}$ generates this root subsystem e.g. by [9, lemma 1]. In this case necessarily $\Sigma=\varnothing$. Otherwise $\Phi=\mathsf{C}_{\ell}$, possibly $\mathsf{B}_{2}=\mathsf{C}_{2}$. Up to the choice of the base $\Delta\subseteq\Phi$ any $5$-grading of $\mathsf{C}_{\ell}$ maps basic roots to 00 except one, and this distinguished root is mapped to $1$ (if it is short) or $2$ (if it is long). It easily follows that $\Sigma=\mathsf{C}_{m}\subseteq\mathsf{C}_{\ell}$ is a proper root subsystem generated by a connected Dynkin subdiagram containing the long basic root or it is empty. But the representation of $\mathrm{Sp}(2m,K)$ on $G_{\{\alpha\in\mathsf{C}_{m+1}\mid\alpha\cdot\mu>0\}}/G_{\mu}$ is the defining representation of symplectic group scheme, where $\mu\in\mathsf{C}_{m+1}$ is the highest root. It follows that $C\leq T$ in all cases.
It remains to show that $C$ trivially acts on all $G_{\alpha}$. But all roots from $\Phi_{>0}$ are trivial on $C$ (recall that roots are some homomorphisms $T\to\mathbb{G}_{\mathrm{m}}$) and such roots generate the root lattice $\mathbb{Z}\Phi$. ∎
If $\Phi$ is of type $\mathsf{BC}_{\ell}$ let $\mathrm{C}^{\mathrm{us}}(G)=\bigcap_{\alpha\in\mathsf{C}_{\ell}}\operatorname{C}_{L}(G_{\alpha})$ be the scheme centralizer of non-ultrashort root subgroups. This group scheme clearly contains the center. It turns out that it is reductive or finite of multiplicative type, the second case holds only for $\mathsf{E}_{8,2}^{66}$. The derived subgroup of $\mathrm{C}^{\mathrm{us}}(G)$ corresponds to the odd component of $u^{-1}(0)$ in the classical cases and to the component $\mathsf{A}_{1}$ in the case $\mathsf{E}_{7,2}^{31}$. More precisely, assuming that $G$ splits and $\operatorname{C}(G)=1$ we have
Tits index $\operatorname{C}^{\mathrm{us}}(G)$ conditions
${{}^{2}\!{\mathsf{A}_{n,r}^{(d)}}}$ $\mathbb{GL}_{n+1-2rd}$ $d\mid n+1$, $n\geq 1$, $2rd\leq n$, $r\geq 2$
$\mathsf{C}_{n,r}^{(d)}$ $\mathbb{S}\mathrm{p}_{2n-2rd}$ $2\leq d=2^{k}\mid 2n$, $rd<n$, $r\geq 2$
${{}^{1}\!{\mathsf{D}_{n,r}^{(d)}}}$, ${{}^{2}\!{\mathsf{D}_{n,r}^{(d)}}}$ $\mathbb{SO}_{2n-2rd}$ $r\geq 2$, $2\leq d=2^{k}\mid 2n$, $rd<n$
${{}^{2}\!{\mathsf{E}_{6,2}^{16^{\prime}}}}$ $\mathbb{G}_{\mathrm{m}}$ none
$\mathsf{E}_{7,2}^{31}$ $\mathbb{SL}_{2}$ none
$\mathsf{E}_{8,2}^{66}$ $\mu_{2}$ none
In the classical cases the group scheme $\operatorname{C}^{\mathrm{us}}(G)$ can be computed using the block structure on $G$ as a matrix group. In the exceptional cases we apply lemma 1 to the splitting isotropic structure (the pinning) on $G$ and get $\operatorname{C}^{\mathrm{us}}(G)\leq T$ (for ${{}^{2}\!{\mathsf{E}_{6,2}^{16^{\prime}}}}$ and $\mathsf{E}_{8,2}^{66}$) or $\mathbb{SL}_{2}\leq\operatorname{C}^{\mathrm{us}}(G)\leq T\,\mathbb{SL}_{2}$. The intersection $T\cap\operatorname{C}^{\mathrm{us}}(G)$ is easy to compute using the root diagrams.
For any root $\alpha\in\Phi$ let $\Gamma_{\alpha}=\{\beta\in\Phi\mid\alpha+\beta\notin\Phi\cup\{0\}\}$ and $Z_{\alpha}=\bigcup_{\beta\in\Gamma_{\alpha}}t_{\beta}(X_{\beta})\subseteq G(K)$ for some generating sets $X_{\beta}\subseteq\mathfrak{g}_{\beta}$ if $\beta$ is not ultrashort and generating sets $(X_{\beta},X_{2\beta})\subseteq(\mathfrak{g}_{\beta}\mathbin{\dot{\oplus}}\mathfrak{g}_{2\beta},\mathfrak{g}_{2\beta})$ for ultrashort $\beta$. It is easy to see by considering all rank $2$ and rank $3$ irreducible root systems that $\Gamma_{\alpha}$ is closed.
Theorem 2.
Let $K$ be a commutative ring, $G$ an isotropic reductive group scheme over $K$ with root system $\Phi$ of rank at least $2$, and $\alpha\in\Phi$ a root.
  • If the type of $\Phi$ is $\mathsf{BC}_{\ell}$ and $\alpha$ is ultrashort, then $\operatorname{C}_{G}(Z_{\alpha})=\operatorname{C}^{\mathrm{us}}(G)\,G_{\alpha}$.
  • If $\alpha=\mathrm{e}_{i}+\mathrm{e}_{j}$ (using the convention $\mathrm{e}_{-i}=-\mathrm{e}_{i}$) is short and the type of $\Phi$ is $\mathsf{C}_{\ell}$ or $\mathsf{BC}_{\ell}$, including $\mathsf{B}_{2}=\mathsf{C}_{2}$, then $\operatorname{C}_{G}(Z_{\alpha})=\operatorname{C}(G)\,G_{2\mathrm{e}_{i}}\,G_{\mathrm{e}_{i}+\mathrm{e}_{j}}\,G_{2\mathrm{e}_{j}}$.
  • Otherwise $\operatorname{C}_{G}(Z_{\alpha})=\operatorname{C}(G)\,G_{\alpha}$.
Proof.
By theorem 1 and lemma 2 $\mathrm{C}_{G}(Z_{\alpha})\leq G^{0}_{\Sigma}$, where
$\Sigma=\{\gamma\in\Phi\mid\angle(\beta,\gamma)\leq\frac{\pi}{2}\text{ for all long }\beta\in\Gamma_{\alpha}\}.$
Considering the saturated root subsystem of rank $2$ (or rank $3$ in the doubly laced cases) containing $\alpha$ and arbitrary $\beta\in\Sigma\setminus\mathbb{R}\alpha$ we see that
  • If $\Phi$ is of type $\mathsf{C}_{\ell}$ with $\ell\geq 2$ and $\alpha=\mathrm{e}_{i}+\mathrm{e}_{j}$ is short (recall that $\mathrm{e}_{-i}=-\mathrm{e}_{i}$), then $\Sigma=\{2\mathrm{e}_{i},\mathrm{e}_{i}+\mathrm{e}_{j},2\mathrm{e}_{j}\}$.
  • If $\Phi$ is of type $\mathsf{BC}_{\ell}$ with $\ell\geq 2$ and $\alpha$ is long, then $\Sigma=\{\alpha,\frac{1}{2}\alpha\}$.
  • If $\Phi$ is of type $\mathsf{BC}_{\ell}$ with $\ell\geq 2$ and $\alpha=\mathrm{e}_{i}+\mathrm{e}_{j}$ is short, then $\Sigma=\{\mathrm{e}_{i},2\mathrm{e}_{i},\mathrm{e}_{i}+\mathrm{e}_{j},\mathrm{e}_{j},2\mathrm{e}_{j}\}$.
  • If $\Phi$ is of type $\mathsf{BC}_{\ell}$ with $\ell\geq 2$ and $\alpha$ is ultrashort, then $\Sigma=\{\alpha,2\alpha\}$.
  • Otherwise $\Sigma=\{\alpha\}$.
Now apply lemma 4 to get
$\operatorname{C}_{G}(Z_{\alpha})\leq\begin{cases}G^{0}_{\{\alpha,2\alpha\}}&\text{ if $\Phi$ is of type $\mathsf{BC}_{\ell}$, $\ell\geq 2$, $\alpha$ is ultrashort};\\ G^{0}_{\{2\mathrm{e}_{i},\mathrm{e}_{i}+\mathrm{e}_{j},2\mathrm{e}_{j}\}}&\text{ if $\Phi$ is of type $\mathsf{C}_{\ell}$ or $\mathsf{BC}_{\ell}$, $\ell\geq 2$, $\alpha=\mathrm{e}_{i}+\mathrm{e}_{j}$ is short};\\ G^{0}_{\alpha}&\text{ otherwise}.\end{cases}$
It remains to show that $\operatorname{C}_{G}(Z_{\alpha})\cap L=\mathrm{C}(G)$ or $\mathrm{C}^{\mathrm{us}}(G)$. If $\alpha$ is not ultrashort we are done by lemma 5 applied to $f(\beta)=2\frac{\alpha\cdot\beta}{\alpha\cdot\alpha}$. Otherwise note that $\Gamma_{\alpha}$ is a parabolic subset of $\mathsf{C}_{\ell}\subseteq\mathsf{BC}_{\ell}$. Applying lemma 5 to $G^{0}_{\{-\beta,\beta\}}$ for various $\beta\in\Gamma_{\alpha}$ with $f(\gamma)=2\frac{\gamma\cdot\beta}{\beta\cdot\beta}$ (more precisely, to its simple subgroup containing $G_{\beta}$ after a splitting étale extension) we see that $\operatorname{C}_{L}(t_{\beta}(X_{\beta}))$ centralizes both $G_{\beta}$ and $G_{-\beta}$, so $\operatorname{C}_{G}(Z_{\alpha})\cap L=\operatorname{C}^{\mathrm{us}}(G)$. ∎
Theorem 3.
Let $G$ an isotropic reductive group scheme over commutative ring $K$ with root system $\Phi$ of rank at least $2$. Let also $X_{\alpha}\subseteq\mathfrak{g}_{\alpha}$ be generating subsets for non-ultrashort $\beta$ and generating sets $(X_{\alpha},X_{2\alpha})\subseteq(\mathfrak{g}_{\alpha}\mathbin{\dot{\oplus}}\mathfrak{g}_{2\alpha},\mathfrak{g}_{2\alpha})$ otherwise. Then
$\operatorname{C}_{G}\bigl{(}\bigcup_{\alpha\in\Phi}t_{\alpha}(X_{\alpha})\bigr{)}=\operatorname{C}(G),\quad\{g\in G\mid\forall\alpha\in\Phi\enskip{{}^{g}\!{t_{\alpha}(X_{\alpha})}}\subseteq G_{\alpha}\}=L.$
Proof.
The first claim follows from lemma 2 applied to opposite long roots and lemma 2. The second one is a corollary of theorem 1 applied to all long roots and lemma 2. ∎
5 Main results
Theorem 4.
Let $G$ be an isotropic reductive group scheme over $K$ such that the rank of $\Phi$ is at least $2$. Then the sets $L(K)$, $\operatorname{C}(G)(K)$, and all $\operatorname{C}(G)(K)\,G_{\alpha}(K)$ are Diophantine in the group $G(K)$. Moreover, the root subgroups $G_{\alpha}(K)$ are Diophantine in the group $G(K)$ unless the type is $\mathsf{B}_{2}=\mathsf{C}_{2}$, $\mathsf{BC}_{2}$, or $\alpha$ is short and the type is $\mathsf{G}_{2}$.
Proof.
By theorem 2 the following groups are Diophantine.
  • $\operatorname{C}^{\mathrm{us}}(G)(K)\,G_{\alpha}(K)$ for ultrashort $\alpha$ and $\Phi$ of type $\mathsf{BC}_{\ell}$;
  • $\operatorname{C}(G)(K)\,G_{2\mathrm{e}_{i}}(K)\,G_{\mathrm{e}_{i}+\mathrm{e}_{j}}(K)\,G_{2\mathrm{e}_{j}}(K)$ for $\Phi$ of type $\mathsf{C}_{\ell}$ or $\mathsf{BC}_{\ell}$, $\ell\geq 2$;
  • $\mathrm{C}(G)(K)\,G_{\alpha}(K)$ for $\alpha$ long or $\alpha$ short and $\Phi$ of type $\mathsf{B}_{\ell}$, $\ell\geq 3$, $\mathsf{F}_{4}$, or $\mathsf{G}_{2}$.
Intersections of Diophantine groups, their products and their commutators with single elements are also Diophantine.
Clearly,
$\operatorname{C}(G)(K)=\operatorname{C}(G)(K)\,G_{\alpha}(K)\cap\operatorname{C}(G)(K)\,G_{-\alpha}(K)$
is Diophantine, where $\alpha$ is any long root. By theorem 1 applied to $G/\operatorname{C}(G)$ we have
$\{g\in G(K)\mid{{}^{g}\!{t_{\alpha}(X)}}\subseteq\operatorname{C}(G)(K)\,G_{\alpha}(K)\}=G^{0}_{\{\beta\mid\angle(\alpha,\beta)\leq\pi/2\}}(K)$
for every long root $\alpha$, where $X\subseteq\mathfrak{g}_{\alpha}$ is a finite generating set. The intersection of all these parabolic subgroups is $L(K)$ by lemma 2.
If $\alpha$ is long and the type of $\Phi$ is neither $\mathsf{C}_{\ell}$ nor $\mathsf{BC}_{\ell}$, then
$G_{\alpha}(K)=\bigl{[}\operatorname{C}(G)(K)\,G_{\beta}(K),t_{\alpha-\beta}(e)\bigr{]}$
is Diophantine, where $\beta$ is long, $\angle(\alpha,\beta)=\frac{\pi}{3}$, and $t_{\alpha-\beta}(e)\in G_{\alpha-\beta}$ is a factor of a Weyl element.
If $\alpha=\mathrm{e}_{i}$ is short and the type of $\Phi$ is $\mathsf{B}_{\ell}$, $\ell\geq 3$, then
$G_{\alpha}(K)=\operatorname{C}(G)(K)\,G_{\alpha}(K)\cap\bigl{[}\operatorname{C}(G)(K)\,G_{\mathrm{e}_{j}}(K),t_{\mathrm{e}_{i}-\mathrm{e}_{j}}(e)\bigr{]}\,G_{\mathrm{e}_{i}+\mathrm{e}_{j}}(K)$
is Diophantine, where $t_{\mathrm{e}_{i}-\mathrm{e}_{j}}(e)$ is a factor of a Weyl element. It follows that the same holds for the type $\mathsf{F}_{4}$.
Now suppose that $\Phi$ is of type $\mathsf{C}_{\ell}$ or $\mathsf{BC}_{\ell}$ and $\ell\geq 3$. We have
$G_{\mathrm{e}_{i}+\mathrm{e}_{j}}(K)=\bigl{[}\bigl{[}\operatorname{C}(G)(K)\,G_{2\mathrm{e}_{j}}(K)\,G_{\mathrm{e}_{j}+\mathrm{e}_{k}}(K)\,G_{2\mathrm{e}_{k}}(K),t_{\mathrm{e}_{i}-\mathrm{e}_{j}}(e)\bigr{]},t_{\mathrm{e}_{j}-\mathrm{e}_{k}}(e^{\prime})\bigr{]},$
where constants are factors of Weyl elements. Next, the group
$G_{2\mathrm{e}_{i}}(K)=\operatorname{C}(G)\,G_{2\mathrm{e}_{i}}(K)\cap\bigl{[}\operatorname{C}(G)\,G_{2\mathrm{e}_{j}}(K),t_{\mathrm{e}_{i}-\mathrm{e}_{j}}(e)]\,G_{\mathrm{e}_{i}+\mathrm{e}_{j}}(K)$
is Diophantine. If $\Phi$ is of type $\mathsf{BC}_{\ell}$, then
$G_{\mathrm{e}_{i}}(K)=\operatorname{C}^{\mathrm{us}}(G)(K)\,G_{\mathrm{e}_{i}}(K)\cap\bigl{[}\operatorname{C}^{\mathrm{us}}(G)(K)\,G_{\mathrm{e}_{j}}(K),t_{\mathrm{e}_{i}-\mathrm{e}_{j}}(e)\bigr{]}\,G_{\mathrm{e}_{i}+\mathrm{e}_{j}}(K)\,G_{\mathrm{e}_{i}-\mathrm{e}_{j}}(K)$
is also Diophantine.
Finally, consider the exceptional cases $\mathsf{C}_{2}$ and $\mathsf{BC}_{2}$. The group $\operatorname{C}(G)(K)\,G_{\mathrm{e}_{i}}(K)$ is Diophantine (in the case of $\mathsf{BC}_{2}$) since it contains $g\in\operatorname{C}^{\mathrm{us}}(K)\,G_{\mathrm{e}_{i}}(K)$ if and only if
$\bigl{[}g,t_{\mathrm{e}_{i}}(X_{0})\bigr{]}\subseteq\operatorname{C}(G)(K)\,G_{2\mathrm{e}_{i}}(K)$
for some generating set $X_{0}\subseteq\mathfrak{g}_{\mathrm{e}_{i}}\mathbin{\dot{\oplus}}\mathfrak{g}_{2\mathrm{e}_{i}}$. We have
$\operatorname{C}(G)\,G_{2\mathrm{e}_{i}}(K)\,G_{\mathrm{e}_{i}+\mathrm{e}_{j}}(K)=\operatorname{C}(G)\,G_{2\mathrm{e}_{i}}(K)\,\prod_{x\in X}\bigl{[}\operatorname{C}(G)\,G_{2\mathrm{e}_{i}}(K)\,G_{\mathrm{e}_{i}+\mathrm{e}_{j}}(K)\,G_{2\mathrm{e}_{j}}(K),t_{\mathrm{e}_{i}-\mathrm{e}_{j}}(x)\bigr{]}$
for a finite generating set $X\subseteq\mathfrak{g}_{\mathrm{e}_{i}-\mathrm{e}_{j}}$ by lemma 3, so $\operatorname{C}(G)\,G_{\mathrm{e}_{i}+\mathrm{e}_{j}}(K)$ is also Diophantine. ∎
Since $G$ is a finitely presented affine scheme, the point group $G(K)$ is e-interpretable in $K$. The following theorem shows the converse.
Theorem 5.
Let $G$ be an isotropic reductive group scheme over $K$ with the root system $\Phi$ of rank at least $2$. Then the ring $K$ is e-interpretable in $G(K)$.
Proof.
We assume that the type of $\Phi$ is neither $\mathsf{G}_{2}$ nor $\mathsf{BC}_{\ell}$ since these cases easily reduce to $\mathsf{A}_{2}$ and $\mathsf{C}_{\ell}$ respectively. Fix some finite $K$-module generating sets $X_{\alpha}\subseteq\mathfrak{g}_{\alpha}$ for all roots $\alpha$ and let $Y_{\alpha}=X_{\alpha}\cup\bigcup_{\beta,\alpha-\beta\in\Phi}[X_{\beta},X_{\alpha-\beta}]$.
Let $\widetilde{K}$ be the set of families
$(k_{\alpha}\colon Y_{\alpha}\to\operatorname{C}(G)(K)\,G_{\alpha}(K)/\operatorname{C}(G)(K))_{\alpha\in\Phi}$
such that
$[k_{\alpha}(y_{\alpha}),t_{\beta}(y_{\beta})]\equiv[t_{\alpha}(y_{\alpha}),k_{\beta}(y_{\beta})]\pmod{\operatorname{C}(G)(K)\,\prod_{\begin{subarray}{c}i\alpha+j\beta\in\Phi\\ i,j\geq 1\\ (i,j)\neq(1,1)\end{subarray}}G_{i\alpha+j\beta}}$
for $y_{\alpha}\in Y_{\alpha}$, $y_{\beta}\in Y_{\beta}$ and
$[k_{\alpha}(x_{\alpha}),t_{\beta}(x_{\beta})]\equiv k_{\alpha+\beta}([x_{\alpha},x_{\beta}])\pmod{\operatorname{C}(G)(K)\,\prod_{\begin{subarray}{c}i\alpha+j\beta\in\Phi\\ i,j\geq 1\\ (i,j)\neq(1,1)\end{subarray}}G_{i\alpha+j\beta}}$
for $x_{\alpha}\in X_{\alpha}$, $x_{\beta}\in X_{\beta}$. Here $\alpha$ and $\beta$ are all roots such that $\alpha+\beta$ is also a root. Clearly, $\widetilde{K}$ is the factor-set of a Diophantine set by a Diophantine equivalence relation.
Every $k\in K$ determines a corresponding element $\widetilde{k}\in\widetilde{K}$, $\widetilde{k}(y_{\alpha})=t_{\alpha}(ky_{\alpha})$. Conversely, consider elements $u_{y_{\alpha}}\in\mathfrak{g}_{\alpha}$ for $y_{\alpha}\in Y_{\alpha}$ such that $k_{\alpha}(y_{\alpha})\in t_{\alpha}(y_{x_{\alpha}})\,\operatorname{C}(G)$. For any root $\alpha$ there is long $\beta\in\Phi$ such that $\alpha+\beta\in\Phi$, so $[u_{y_{\alpha}},y_{\beta}]=[y_{\alpha},u_{y_{\beta}}]$. Recall that the Lie bracket $[{-},{=}]\colon\mathfrak{g}_{\alpha}\times\mathfrak{g}_{\beta}\to\mathfrak{g}_{\alpha+\beta}$ is non-degenerate (by lemma 3 if $\alpha$ is short). For any linear relation $\sum_{y_{\alpha}\in Y_{\alpha}}y_{\alpha}a_{\alpha}=0$ the equation implies that $\sum_{y_{\alpha}\in Y_{\alpha}}u_{y_{\alpha}}a_{\alpha}=0$, so there is a unique linear map $U_{\alpha}\colon\mathfrak{g}_{\alpha}\to\mathfrak{g}_{\alpha}$ such that $u_{y_{\alpha}}=U_{\alpha}(y_{\alpha})$.
The defining congruences imply that $[U_{\alpha}(x),y]=U_{\alpha+\beta}([x,y])=[x,U_{\beta}(y)]$. We claim that such maps $U_{\alpha}$ are necessarily scalar. It suffices to consider $\Phi$ of rank $2$ and to check the claim for only one root $\alpha$. This is clear if one of $\mathfrak{g}_{\alpha}$ is one-dimensional.
  • In the linear-like case (${{}^{1}\!{\mathsf{E}_{6,2}^{28}}}$ and ${{}^{1}\!{\mathsf{A}_{3d-1,2}^{(d)}}}$) we have
    $U_{\mathrm{e}_{1}-\mathrm{e}_{2}}(x)\,y=U_{\mathrm{e}_{1}-\mathrm{e}_{3}}(xy)=x\,U_{\mathrm{e}_{2}-\mathrm{e}_{3}}(y)$
    for $x,y\in A$, where $A$ is a composition algebra of rank $8$ or an Azumaya algebra over $K$. Hence $U_{\mathrm{e}_{1}-\mathrm{e}_{2}}(x)=U_{\mathrm{e}_{1}-\mathrm{e}_{3}}(x)$ (i.e. all maps $U_{\alpha}$ coincide) and $U_{\alpha}(x)=x\,U_{\alpha}(1)=U_{\alpha}(1)\,x$, so $U_{\alpha}(1)$ lies in the center of $A$, and this center is precisely $K$.
  • In the orthogonal-like case ($\mathsf{B}_{n,2}$, ${{}^{1}\!{\mathsf{D}_{n,2}^{(d)}}}$, ${{}^{2}\!{\mathsf{D}_{n,2}^{(d)}}}$) we have
    $U_{\mathrm{e}_{2}}(m)\,r=m\,U_{\mathrm{e}_{1}-\mathrm{e}_{2}}(r)=U_{\mathrm{e}_{1}}(mr)$
    for $r\in R$ and $m\in M$, where $R$ is an Azumaya algebra over $K$ and $M_{R}$ is a faithfully projective module. We have $U_{\mathrm{e}_{1}}(m)=U_{\mathrm{e}_{2}}(m)=m\,U_{\mathrm{e}_{1}-\mathrm{e}_{2}}(1)$ and $U_{\mathrm{e}_{1}-\mathrm{e}_{2}}(r)=r\,U_{\mathrm{e}_{1}-\mathrm{e}_{2}}(1)=U_{\mathrm{e}_{1}-\mathrm{e}_{2}}(1)\,r$, so $U_{\mathrm{e}_{1}-\mathrm{e}_{2}}(1)\in K=\operatorname{C}(R)$.
  • In the remaining symplectic-like case we have
    $U_{\mathrm{e}_{1}-\mathrm{e}_{2}}(r)\,u=r\,U_{2\mathrm{e}_{2}}(u)=U_{\mathrm{e}_{1}+\mathrm{e}_{2}}(ru)$
    for $r\in R$ and $u\in\Lambda$, where $R$ is an Azumaya algebra over $K$ or its quadratic étale extension and $\Lambda\leq R$ is a form parameter containing invertible element $\iota$. It follows that $U_{2\mathrm{e}_{2}}(u)=U_{\mathrm{e}_{1}-\mathrm{e}_{2}}(1)\,u$, $U_{\mathrm{e}_{1}-\mathrm{e}_{2}}(r)=r\,U_{\mathrm{e}_{1}-\mathrm{e}_{2}}(1)$, and $U_{\mathrm{e}_{1}+\mathrm{e}_{2}}(r)=r\iota^{-1}\,U_{\mathrm{e}_{1}-\mathrm{e}_{2}}(1)\,\iota$. Then $U_{\mathrm{e}_{1}-\mathrm{e}_{2}}(1)\,u\iota^{-1}=u\iota^{-1}\,U_{\mathrm{e}_{1}-\mathrm{e}_{2}}(1)$ and $U_{\mathrm{e}_{1}-\mathrm{e}_{2}}(1)\in\Lambda\iota^{-1}$. In the subcase ${{}^{2}\!{\mathsf{A}_{n,2}^{(d)}}}$ the group $\Lambda=R_{0}$ is an Azumaya algebra over $K$ (so $R=R_{0}\otimes\operatorname{C}(R)$) and $U_{\mathrm{e}_{1}-\mathrm{e}_{2}}(1)\in\Lambda\iota^{-1}\cap\operatorname{C}(R)=K$. In the subcase $\mathsf{C}_{n,2}^{(d)}$ the form parameter is
    $\Lambda=\{x\in\mathrm{M}(d,K)\mid x_{ij}=x_{d+1-j,d+1-i}\}$
    and $\iota=1$tale locally, so $U_{\mathrm{e}_{1}-\mathrm{e}_{2}}(1)\in\operatorname{C}_{R}(\Lambda)=K$. In the last subcase ${{}^{1}\!{\mathsf{D}_{n,2}^{(d)}}}$ we have
    $\Lambda=\{x\in\mathrm{M}(d,K)\mid x_{ij}=-x_{d+1-j,d+1-i},x_{i,d+1-i}=0\}$
    and $\iota=e_{11}+\ldots+e_{\frac{d}{2},\frac{d}{2}}-e_{\frac{d}{2}+1,\frac{d}{2}+1}-\ldots-e_{dd}$tale locally, so again $U_{\mathrm{e}_{1}-\mathrm{e}_{2}}(1)\in\operatorname{C}_{R}(\Lambda\iota^{-1})\cap\Lambda\iota^{-1}=K$ (if $d\geq 2$, then already $\operatorname{C}_{R}(\Lambda\iota^{-1})=K$).
It remains to define ring operations on $\widetilde{K}$. The group operation is directly induced from $G(K)$. Finally, the product of $(k_{\alpha})_{\alpha}$ and $(l_{\alpha})_{\alpha}$ is such a family $(m_{\alpha})_{\alpha}$ that
$[k_{\alpha}(x_{\alpha}),l_{\beta}(x_{\beta})]\equiv m_{\alpha+\beta}([x_{\alpha},x_{\beta}])\pmod{\operatorname{C}(G)(K)\,\prod_{\begin{subarray}{c}i\alpha+j\beta\in\Phi\\ i,j\geq 1\\ (i,j)\neq(1,1)\end{subarray}}G_{i\alpha+j\beta}}$
for $x_{\alpha}\in X_{\alpha}$, $x_{\beta}\in X_{\beta}$, and roots $\alpha$, $\beta$ such that $\alpha+\beta$ is also a root. It suffices to impose this condition only for one pair $(\alpha,\beta)$. ∎
Theorem 6.
Let $K$ be a commutative ring with an enumeration by $\mathbb{N}$ or its finite subset such that the ring operations are computable. Let also $G$ be an isotropic reductive group over $K$ such that the rank of $\Phi$ is at least $2$. Then the Diophantine problems for $K$ and $G(K)$ are equivalent, i.e. they reduce to each other by algorithms. Here the enumeration on $G(K)$ is obtained from some closed embedding $G\subseteq\mathbb{A}^{n}_{K}$, it is independent of the embedding up to a computable permutation.
Proof.
Clearly, the group operations on $G(K)$ are also computable. Every Diophantine subset of $G(K)^{k}$ may be considered as a Diophantine subset of $K^{kn}$ using the e-interpretation of $G(K)$ in $K$ and this transformation is clearly computable in terms of defining formulae. In the other direction apply theorem 5. ∎

### v2

@@
 A formula
-$\exists\vec{y}\enskip\bigwedge_{i=1}^{n}P_{i}(\vec{x},\vec{y})$
+$\exists\vec{y}\enskip\bigl(P_{1}(\vec{x},\vec{y})\wedge\ldots\wedge P_{n}(\vec{x},\vec{y})\bigr)$
 for elementary formulae $P_{i}(\vec{x},\vec{y})$ is called Diophantine (also positive-primitive or regular) in the variables $\vec{x}$.
@@
 A closely related is the notion of e-interpretation.
-An e-interpretation of a model $M$ in a model $N$ (of possibly another language) is a surjection $X\to M$ for some Diophantine $X\subseteq N^{n}$ such that the lifts of all relations on $M$ to $X$ from its language are Diophantine (including the kernel pair lifting the equality), as well as lifts of all graphs of operations on $M$.
+An e-interpretation of a model $M$ in a model $N$ (of possibly another language) is a surjection $X\to M$ for some Diophantine $X\subseteq N^{n}$ such that the lifts of all relations on $M$ to $X$ from its language are Diophantine (including the equivalence relation lifting the equality), as well as lifts of all graphs of operations on $M$.
 For example, if $K$ is a commutative ring and $G$ is a finitely presented affine group scheme over $K$, then the group $G(K)$ is e-interpretable in the ring $K$.
@@
 For example, the Diophantine problem $\mathcal{D}(\mathbb{Z})$ (for the language of rings and a natural enumeration) is undecidable by famous Matiyasevich’s theorem.
-On the other hand, $\mathcal{D}(\mathbb{Q}^{\mathrm{alg}})$ is decidable (again for the language of rings and a natural enumeration) since every algebraically closed field admits quantifier elimination.
+On the other hand, $\mathcal{D}(\mathbb{Q}^{\mathrm{alg}})$ is decidable (again for the language of rings and a natural enumeration) because every algebraically closed field admits quantifier elimination.
 Now let $K$ be a commutative ring and $G$ be a reductive group scheme over $K$ in the sense of [13].
 If $G$ is splits, i.e.
-it is a Chevalley–Demazure group scheme, simple, and of rank at least $2$, then by the main result of Elena Bunina, Alexey Myasnikov, and Eugene Plotkin [2] the Diophantine problems $\mathcal{D}(K)$ and $\mathcal{D}(G(K))$ are equivalent.
-More precisely, if $K$ is countable with fixed enumeration, then these two problems are reducible to each other by explicit algorithms.
+it is a Chevalley–Demazure group scheme, simple, and of rank at least $2$, then by the main result of Elena Bunina, Alexey Myasnikov, and Eugene Plotkin [12] the ring $K$ and the group $G(K)$ are e-interpretable in each other and the Diophantine problems $\mathcal{D}(K)$ and $\mathcal{D}(G(K))$ are equivalent.
+More precisely, if $K$ is countable with fixed enumeration, then the Diophantine problems are reducible to each other by explicit algorithms.
+See also [2, 3, 4, 5, 7, 8, 9, 10, 11, 14, 16, 17, 18, 19, 23] for other results on model theory of linear groups over rings, including e-interpretability and first order interpretability.
 We are going to generalize this result to isotropic $G$, not necessarily split.
@@
 Finally, since $K$ and $G(K)$ e-interpret each other, their Diophantine problems are equivalent (theorem 6).
-As an application of preparatory technical results we also find the centralizer and the normalizer of all root subgroups together, this is a generalization of [1] to the isotropic case.
+As an application of preparatory technical results we also find the centralizer and the normalizer of all root subgroups (and even their finite subsets) together, this is a generalization of [1, 15].
 For example, it follows that the Diophantine problem for $\mathrm{SO}(E_{8}\perp H\perp H)$ is unsolvable, where $E_{8}$ is the lattice spanned by the root system of type $\mathsf{E}_{8}$ and $H$ is a hyperbolic plane over $\mathbb{Z}$.
@@
 2 Isotropic reductive groups
-Let $K$ be a commutative ring (all rings in this paper are unital) and $G$ a reductive group scheme over $K$ [3, Exp.
-XIX].
-By root systems we mean crystallographic root systems possibly of type $\mathsf{BC}_{\ell}$.
-Recall that there is an étale extension $K\subseteq\widetilde{K}$ such that $G_{\widetilde{K}}$ splits with possibly non-constant root datum, i.e.
-there is a decomposition $\widetilde{K}=\widetilde{K}_{1}\times\ldots\times\widetilde{K}_{n}$ such that each $G_{\widetilde{K}_{i}}$ is isomorphic to a split reductive group scheme $\mathrm{G}(X^{\vee},\widetilde{\Phi}^{\vee},X,\widetilde{\Phi};\widetilde{K}_{i})$ over $\widetilde{K}_{i}$ [3, corollary XXII.2.3].
-We always assume that $n=1$, i.e.
-the root datum $(X^{\vee},\widetilde{\Phi}^{\vee},X,\widetilde{\Phi})$ is constant over $\widetilde{K}$, and that the root system $\widetilde{\Phi}$ is irreducible.
-But $G$ is not necessarily semisimple, i.e.
-the scheme center $\operatorname{C}(G)$ may contain a non-trivial torus (such generality is required e.g.
-in applications of lemma 5 below).
-In this paper we call $G$ isotropic with root system $\Phi$ if there are subgroups $L\leq G$ and $G_{\alpha}\leq G$ for $\alpha\in\Phi$ such that after some (each sufficiently large) étale extension $K\subseteq\widetilde{K}$ there are an isomorphism
-$F\colon G_{\widetilde{K}}\to\mathrm{G}(X^{\vee},\widetilde{\Phi}^{\vee},X,\widetilde{\Phi};\widetilde{K})$
-with a split reductive group scheme and a map $u\colon\widetilde{\Phi}\to\Phi\sqcup\{0\}$ induced by a linear map of the ambient vector spaces such that
-  • $\Phi$ is contained in the image of $u$;
-  • $F(G_{\alpha,\widetilde{K}})=\langle U_{\beta}\mid u(\beta)\in\{\alpha,2\alpha\}\rangle$ is a standard unipotent subgroup, where $U_{\beta}$ are the root subgroups of the split reductive group scheme;
-  • $F(L_{\widetilde{K}})=\mathrm{G}(X^{\vee},\widetilde{\Phi}^{\vee},X,\widetilde{\Phi};\widetilde{K})^{0}_{u^{-1}(0)}$ is a reductive subsystem subgroup, namely, the subgroup of type (RC) corresponding to the subset $u^{-1}(0)$ [3, §XXII.5.11];
-  • the group $G$ together with the family $(G_{\alpha})_{\alpha\in\Phi}$ is $\Phi$-graded, i.e. $G_{2\alpha}\leq G_{\alpha}$ for ultrashort $\alpha$,
-    $[G_{\alpha},G_{\beta}]\leq\langle G_{\gamma}\mid\gamma\in(\mathbb{N}_{+}\alpha+\mathbb{N}_{+}\beta)\cap\Phi\rangle$
-    (this automatically follows from the previous condition), and there are Weyl elements
-    $n_{\alpha}\in G_{\alpha}(K)\,G_{-\alpha}(K)\,G_{\alpha}(K)$
-    for all $\alpha\in\Phi$, i.e. elements with the property ${{}^{n_{\alpha}}\!{G_{\beta}}}=G_{s_{\alpha}(\beta)}$;
-  • The map $u\colon\widetilde{\Phi}\to\Phi\sqcup\{0\}$ comes from one of irreducible Tits indices.
-For example, all reductive groups over semi-local rings (and even LG-rings [4]) with connected spectra are naturally isotropic using their minimal parabolic subgroups by [3, Exp.
-XXVI] and [6].
-All these isotropic structures are conjugate by elements of $G(K)$.
+In this paper all root systems are crystallographic, but not necessarily reduced, i.e.
+we allow components of type $\mathsf{BC}_{\ell}$.
+Let $K$ be a non-zero unital commutative ring and $\widetilde{\Phi}$ be a reduced root system.
+Choose a weight lattice
+$\mathbb{Z}\widetilde{\Phi}\leq\Lambda\leq\bigl\{x\mid\forall\alpha\in\widetilde{\Phi}\enskip 2\frac{x\cdot\alpha}{\alpha\cdot\alpha}\in\mathbb{Z}\bigr\}$
+[…]
+$L=\bigl\langle\widetilde{T},\widetilde{U}_{\alpha}\mid u(\alpha)=0\bigr\rangle,\quad U_{\alpha}=\bigl\langle\widetilde{U}_{\alpha^{\prime}}\mid u(\alpha^{\prime})\in\{\alpha,2\alpha\}\bigr\rangle$
+[…]
+[…]
+This class of group schemes is interesting because every simple reductive group scheme $G$ over a local ring $K$ is anisotropic or isotropic in the above sense in a canonical way up to conjugation, where
+$L\rtimes\bigl\langle U_{\alpha}\mid\alpha\in\Phi^{+}\bigr\rangle$
+is a minimal parabolic subgroup, see [13, XXVI] and [21] for details.
+Moreover, if $G$ is a locally isotropic simple reductive group scheme over arbitrary ring $K$ and $\mathfrak{p}\trianglelefteq K$ is a prime ideal, then the isotropic structure of $G_{\mathfrak{p}}$ is well defined on $G_{s}$ for sufficiently large $s\in K\setminus\mathfrak{p}$.
 Recall [24] that an irreducible Tits index $(\widetilde{\Phi},\Gamma,J)$ consists of a reduced irreducible root system $\widetilde{\Phi}$, a subgroup $\Gamma$ of its group of outer automorphisms (i.e.
-the automorphism group of its Dynkin diagram), and a $\Gamma$-invariant subset $J$ of vertices of the Dynkin diagram satisfying an additional condition (namely, that it may be constructed by a reductive group scheme over a field using its minimal parabolic subgroup).
+the automorphism group of its Dynkin diagram), and a $\Gamma$-invariant subset $J$ of vertices of the Dynkin diagram from the explicit list [24, table II] or [21, appendix].
 Then $\Phi$ is the image of $\widetilde{\Phi}$ in the factor-space of the ambient vector space by $\Gamma$ and the span of basic roots not in $J$, the rank of $\Phi$ is the number of $\Gamma$-orbits in $J$.
-An index is usually denoted as ${{}^{g}\!{\mathsf{X}_{n,r}^{t}}}$, where $g=|\Gamma|$, $\mathsf{X}_{n}$ is the type of $\widetilde{\Phi}$, $r$ is the rank of $\Phi$, and $t$ is an additional parameter (see the list below).
-Here is the list of irreducible Tits indices using the standard (Bourbaki) numeration of simple roots.
-  • ${{}^{1}\!{\mathsf{A}_{n,r}^{(d)}}}$, where $d(r+1)=n+1$ and $n\geq 1$.
-Here $J=\{d,2d,\ldots,rd\}$, $\Phi=\mathsf{A}_{r}$, modulo the center $G(K)=\mathrm{PGL}(r+1,A)$ for an Azumaya algebra $A$ over $K$ of rank $d$.
-Also, $u^{-1}(0)=(r+1)A_{d-1}$ and $|u^{-1}(\alpha)|=d^{2}$ for every root $\alpha$.
-  • ${{}^{2}\!{\mathsf{A}_{n,r}^{(d)}}}$, where $d\mid n+1$, $n\geq 1$, and $12rd\leq n+$ 1. Here
-    $J=\{d,2d,\ldots,rd;n+1-rd,\ldots,n+1-2d,n+1-d\}.$
-    The relative root system is $\Phi=\mathsf{C}_{r}$ for $12rd=n+$ 1 and $\Phi=\mathsf{BC}_{r}$ otherwise.
-Modulo the center $G(K)=\mathrm{SU}(A,h)$ for an Azumaya algebra $A$ of rank $d$ over a quadratic étale extension $K\subseteq K^{\prime}$ and for a non-degenerate hermitian form $h$ over $A$ of rank $\frac{n+1}{d}$ and Witt index $r$ (i.e.
-with a chosen hyperbolic subspace of rank $2r$).
-Also, $u^{-1}(0)=2r\mathsf{A}_{d-1}+\mathsf{A}_{n-2rd}$ (the last summand is vacuous for $2rd\in\{n,n+1\}$), $|u^{-1}(\alpha)|=d^{2}$ for long $\alpha$, $|u^{-1}(\alpha)|=2d^{2}$ for short $\alpha$, $|u^{-1}(\alpha)|=2d(n+1-2rd)$ for ultrashort $\alpha$ (and $2rd\leq n$).
-  • $\mathsf{B}_{n,r}$, where $n\geq 2$ and $r\leq n$.
-Here $J=\{1,2,\ldots,r\}$, $\Phi=\mathsf{B}_{r}$, and $G(K)=\mathrm{SO}(K,q)$ modulo the center, where $q$ is a semi-regular quadratic form of rank $12n+$ 1 and Witt index $r$.
-Also, $u^{-1}(0)=\mathsf{B}_{n-r}$, $|u^{-1}(\alpha)|=1$ for long $\alpha$, and $|u^{-1}(\alpha)|=2n+1-2r$ for short $\alpha$.
-  • $\mathsf{C}_{n,r}^{(d)}$, where $n\geq 3$, $d=2^{k}\mid 2n$, $rd\leq n$, and $n=r$ in the case $d=1$.
-Here $J=\{d,2d,\ldots,rd\}$, $\Phi=\mathsf{C}_{r}$ for $rd=n$ and $\Phi=\mathsf{BC}_{r}$ otherwise.
-Modulo the center $G(K)=\mathrm{U}(A,h)$, where $A$ is an Azumaya algebra over $K$ of rank $d$ with symplectic involution, $h$ is a non-degenerate anti-hermitian form of rank $\frac{2n}{d}$ and Witt index $r$ (with the additional condition that $h$ comes from a symplectic form, not an arbitrary alternating form).
-Also, $u^{-1}(0)=r\mathsf{A}_{d-1}+\mathsf{C}_{n-rd}$, $|u^{-1}(\alpha)|=\frac{d(d+1)}{2}$ for long $\alpha$, $|u^{-1}(\alpha)|=d^{2}$ for short $\alpha$, $|u^{-1}(\alpha)|=2d(n-rd)$ for ultrashort $\alpha$ (and $rd<n$).
-  • ${{}^{1}\!{\mathsf{D}_{n,r}^{(d)}}}$, where $n\geq 4$, $d=2^{k}\mid 2n$, $rd\leq n$, and $n\neq rd+1$.
-Here $J=\{d,2d,\ldots,rd\}$, $\Phi=\mathsf{D}_{r}$ for $rd=n$ and $d=1$, $\Phi=\mathsf{C}_{r}$ for $rd=n$ and $d>1$, $\Phi=\mathsf{B}_{r}$ for $rd<n$ and $d=1$, and $\Phi=\mathsf{BC}_{r}$ for $rd<n$ and $d>1$.
-Modulo the center $G(K)=\mathrm{SU}(A,h,q)$, where $A$ is an Azumaya algebra over $K$ of rank $d$ with orthogonal involution, $h$ is a non-degenerate hermitian form of rank $\frac{2n}{d}$ and Witt index $r$, and $q$ is a suitable associated quadratic form.
-Also, $u^{-1}(0)=r\mathsf{A}_{d-1}+\mathsf{D}_{n-rd}$, $|u^{-1}(\alpha)|=\frac{d(d-1)}{2}$ for long $\alpha$ (and $d>1$), $|u^{-1}(\alpha)|=d^{2}$ for short $\alpha$, $|u^{-1}(\alpha)|=2d(n-rd)$ for ultrashort $\alpha$ (and $rd<n$).
-Here roots of the relative root systems of types $\mathsf{D}_{r}$ and $\mathsf{B}_{r}$ are called short and ultrashort depending on their length, not long and short.
-  • ${{}^{2}\!{\mathsf{D}_{n,r}^{(d)}}}$, where $n\geq 4$, $d=2^{k}\mid 2n$, and $rd\leq n-1$.
-Here $J=\{d,2d,\ldots,rd\}$ for $rd<n-1$ and $J=\{d,2d,\ldots,n-1,n\}$ otherwise, $\Phi=\mathsf{BC}_{r}$ for $d>1$ and $\Phi=\mathsf{B}_{r}$ otherwise.
-Modulo the center $G(K)=\mathrm{SU}(A,h,q)$, where $A$ is an Azumaya algebra over $K$ of rank $d$ with orthogonal involution, $h$ is a non-degenerate hermitian form of rank $\frac{2n}{d}$ and Witt index $r$, and $q$ is a suitable associated quadratic form.
-Also, $u^{-1}(0)=r\mathsf{A}_{d-1}+\mathsf{D}_{n-rd}$, $|u^{-1}(\alpha)|=\frac{d(d-1)}{2}$ for long $\alpha$ (and $d>1$), $|u^{-1}(\alpha)|=d^{2}$ for short $\alpha$, $|u^{-1}(\alpha)|=2d(n-rd)$ for ultrashort $\alpha$.
-As in the previous case, roots of the relative root system of type $\mathsf{B}_{r}$ are called short and ultrashort.
-Exceptional Tits indices are given in the following table.
+An index is usually denoted as $^{g}\!{\mathsf{X}_{n,r}^{t}}$}, where $g=|\Gamma|$, $\mathsf{X}_{n}$ is the type of $\widetilde{\Phi}$, $r$ is the rank of $\Phi$, and $t$ is an additional parameter.
+We are interested only in isotropic Tits indices, i.e.
+with $r\neq 0$.
+An outer automorphism group of an irreducible Tits index is the group of such automorphisms $f$ of the root system $\widetilde{\Phi}$ that $f$ preserves the basis and $u\circ f=u$.
+Now we give the list of all isotropic irreducible Tits indices.
+It is divided in two parts, classical indices (forming infinite families) and finitely many exceptional ones.
+The numeration of simple roots in all Dynkin diagrams follows Bourbaki [6].
+For classical root systems we use the conventions $\mathsf{A}_{-1}=\mathsf{A}_{0}=\mathsf{B}_{0}=\mathsf{C}_{0}=\mathsf{BC}_{0}=\mathsf{D}_{0}=\mathsf{D}_{1}=\varnothing$ and $\mathsf{D}_{2}=2\mathsf{A}_{1}$, as well as the exceptional isomorphisms $\mathsf{A}_{1}=\mathsf{B}_{1}=\mathsf{C}_{1}$, $\mathsf{B}_{2}=\mathsf{C}_{2}$, $\mathsf{A}_{3}=\mathsf{D}_{3}$.
+  • $^{1}\!{\mathsf{A}_{n,r}^{(d)}}$}, where $d(r+1)=n+1$, $d\geq 1$, and $n\geq 1$.
+Here $\widetilde{\Phi}=\mathsf{A}_{n}$, $J=\{d,2d,\ldots,rd\}$, and $\Phi=\mathsf{A}_{r}$.
+The kernel of $u$ has the type $(r+1)A_{d-1}$ and $|u^{-1}(\alpha)|=d^{2}$ for every root $\alpha$.
+The outer automorphism group is trivial for $r\geq 2$ or $^{1}\!{\mathsf{A}_{1,1}^{(1)}}$}, it is cyclic of order $2$ for $r=1$ and $d\geq 2$.
+    Such split isotropic structure on $\mathbb{SL}_{n+1}$ corresponds to the block decomposition by $d\times d$ blocks, $U_{\alpha}$ are the groups subschemes of matrices differing from $1$ only in one non-diagonal block, and $L$ is the group subscheme of block diagonal matrices.
+  • $^{2}\!{\mathsf{A}_{n,r}^{(d)}}$}, where $d\mid n+1$, $2rd\leq n+1$, $d\geq 1$, $r\geq 1$, and $n\geq 2$. Here $\widetilde{\Phi}=\mathsf{A}_{n}$,
+    $J=\{d,2d,\ldots,rd\}\sqcup\{n+1-rd,n+1-(r-1)d,\ldots,n+1-d\}$
+    for $2rd\leq n$ and $J=\{d,2d,\ldots,n+1-d\}$ for $2rd=n+1$, $\Phi=\mathsf{BC}_{r}$ for $2rd\leq n$ and $\Phi=\mathsf{C}_{r}$ for $2rd=n+1$.
+The kernel of $u$ has the type, $2r\mathsf{A}_{d-1}+\mathsf{A}_{n-2rd}$, $|u^{-1}(\alpha)|=2d(n+1-2rd)$ for ultrashort $\alpha$, $|u^{-1}(\alpha)|=2d^{2}$ for short $\alpha$, $|u^{-1}(\alpha)|=d^{2}$ for long $\alpha$.
+The outer automorphism group is cyclic of order $2$.
+    Such split isotropic structure on $\mathbb{SL}_{n+1}$ corresponds to the block decomposition by $r$ strips of width $d$ from each side and, possibly, the central strip of width n+12rdn+1-2rd, both vertical and horizontal.
+We number these strips by integers from r-r to $r$, 0 corresponds to the central strip.
+The group subscheme $L$ consists of block-diagonal matrices.
+If $\alpha$ is short, then $U_{\alpha}$ consists of matrices differing from $1$ only in two blocks with indices $(i,j)$ and (j,i)(-j,-i) for some $0\neq|i|\neq|j|\neq 0$.
+If $\alpha$ is long, then $U_{\alpha}$ consists of matrices differing from $1$ only in the block with the indices (i,i)(-i,i) for some $i\neq 0$.
+Finally, if $\alpha$ is ultrashort, then $U_{\alpha}$ consists of matrices differing from $1$ only in the blocks with the indices $(0,i)$, (i,i)(-i,i), and (i,0)(-i,0) for some $i\neq 0$.
+  • $\mathsf{B}_{n,r}$, where $1\leq r\leq n$.
+Here $\widetilde{\Phi}=\mathsf{B}_{n}$, $J=\{1,2,\ldots,r\}$, and $\Phi=\mathsf{B}_{r}$.
+The kernel of $u$ had the type $\mathsf{B}_{n-r}$, $|u^{-1}(\alpha)|=2n+1-2r$ for short $\alpha$, and $|u^{-1}(\alpha)|=1$ for long $\alpha$.
+The outer automorphism group is trivial.
+    Such split isotropic structure on $\mathbb{SO}_{2n+1}$ is induced from the block decomposition $^{2}\!{\mathsf{A}_{2n+1,r}^{(1)}}$} on the standard representation, i.e.
+$\mathbb{SO}_{2n+1}\leq\mathbb{SL}_{2n+1}$ is the stabilizer of the quadratic form
+    $q(\vec{x})=x_{1}x_{2n+1}+x_{2}x_{2n}+\ldots+x_{n}x_{n+1}+x_{n+1}^{2}.$
+  • $\mathsf{C}_{n,r}^{(d)}$, where $d=2^{k}\mid 2n$, $rd\leq n$, $d\geq 1$, $r\geq 1$, and $n=r$ in the case $d=1$.
+Here $\widetilde{\Phi}=\mathsf{C}_{n}$, $J=\{d,2d,\ldots,rd\}$, $\Phi=\mathsf{BC}_{r}$ for $rd\leq n-1$ and $\Phi=\mathsf{C}_{r}$ for $rd=n$.
+The kernel of $u$ has the type $u^{-1}(0)=r\mathsf{A}_{d-1}+\mathsf{C}_{n-rd}$, $|u^{-1}(\alpha)|=2d(n-rd)$ for ultrashort $\alpha$, $|u^{-1}(\alpha)|=d^{2}$ for short $\alpha$, $|u^{-1}(\alpha)|=\frac{d(d+1)}{2}$ for long $\alpha$.
+The outer automorphism group is trivial.
+    Such split isotropic structure on $\mathbb{Sp}_{2n}$ is induced from the block decomposition $^{2}\!{\mathsf{A}_{2n,r}^{(1)}}$} on the standard representation, i.e.
+$\mathbb{Sp}_{2n}\leq\mathbb{GL}_{2n}$ is the stabilizer of the symplectic form
+    $B(\vec{x},\vec{y})=x_{1}y_{2n}+x_{2}y_{2n-1}+\ldots+x_{n}y_{n+1}-x_{n+1}y_{n}-x_{n+2}y_{n-1}-\ldots-x_{2n}y_{1}.$
+  • $^{1}\!{\mathsf{D}_{n,r}^{(d)}}$}, where $d=2^{k}\mid 2n$, $rd\leq n$, $rd\neq n-1$, $d\geq 1$, $r\geq 1$, $n\geq 3$.
+Here $\widetilde{\Phi}=D_{n}$, $J=\{d,2d,\ldots,rd\}$, $\Phi=\mathsf{BC}_{r}$ for $rd\leq n-2$ and $d\geq 2$, $\Phi=\mathsf{B}_{r}$ for $rd\leq n-2$ and $d=1$, $\Phi=\mathsf{C}_{r}$ for $rd=n$ and $d\geq 2$, $\Phi=\mathsf{D}_{r}$ for $rd=n$ and $d=1$.
+The kernel of $u$ has the type $r\mathsf{A}_{d-1}+\mathsf{D}_{n-rd}$, $|u^{-1}(\alpha)|=2d(n-rd)$ for ultrashort $\alpha$, $|u^{-1}(\alpha)|=d^{2}$ for short $\alpha$, $|u^{-1}(\alpha)|=\frac{d(d-1)}{2}$ for long $\alpha$ (if $d=1$, then $|u^{-1}(\alpha)|=2d(n-rd)$ for short $\alpha$ and $|u^{-1}(\alpha)|=d^{2}$ for long $\alpha$).
+The outer automorphism group is trivial for $rd=n$ except the cases $^{1}\!{\mathsf{D}_{4,2}^{(2)}}$} and $^{1}\!{\mathsf{D}_{4,1}^{(4)}}$}; it is cyclic of order two for $^{1}\!{\mathsf{D}_{4,2}^{(2)}}$}, $^{1}\!{\mathsf{D}_{4,1}^{(4)}}$}, and for $rd\leq n-2$ except the case $^{1}\!{\mathsf{D}_{4,1}^{(2)}}$}; it is the whole triality group of order $6$ in the case $^{1}\!{\mathsf{D}_{4,1}^{(2)}}$}.
+    Such split isotropic structure on $\mathbb{SO}_{2n}$ is induced from the block decomposition $^{2}\!{\mathsf{A}_{2n,r}^{(1)}}$} on the standard representation, i.e.
+$\mathbb{SO}_{2n}\leq\mathbb{GL}_{2n}$ is the stabilizer of the quadratic form
+    $q(\vec{x})=x_{1}x_{2n+1}+x_{2}x_{2n}+\ldots+x_{n}x_{n+1}$
+    and the Dickson invariant (or the determinant if $2$ is invertible).
+  • $^{2}\!{\mathsf{D}_{n,r}^{(d)}}$}, where $d=2^{k}\mid 2n$, $rd\leq n-1$, $d\geq 1$, $r\geq 1$, $n\geq 3$.
+Here $\widetilde{\Phi}=\mathsf{D}_{n}$, $J=\{d,2d,\ldots,rd\}$ for rd<n2rd<n-2 and $J=\{d,2d,\ldots,(r-1)d,n-1,n\}$ for rd=n1rd=n-1, $\Phi=\mathsf{BC}_{r}$ for $d\geq 2$ and $\Phi=\mathsf{B}_{r}$ for $d=1$.
+The kernel of $u$ has the type $r\mathsf{A}_{d-1}+\mathsf{D}_{n-rd}$, $|u^{-1}(\alpha)|=2d(n-rd)$ for ultrashort $\alpha$, $|u^{-1}(\alpha)|=d^{2}$ for short $\alpha$, $|u^{-1}(\alpha)|=\frac{d(d-1)}{2}$ for long $\alpha$ (if $d=1$, then $|u^{-1}(\alpha)|=2d(n-rd)$ for short $\alpha$ an $|u^{-1}(\alpha)|=d^{2}$ for long $\alpha$).
+The outer automorphism group is cyclic of order $2$ except the case $^{2}\!{\mathsf{D}_{4,1}^{(2)}}$} and it is the triality group of order $6$ in this exceptional case.
+    Such split isotropic structure on $\mathbb{SO}_{2n}$ is induced from the block decomposition $^{2}\!{\mathsf{A}_{2n,r}^{(1)}}$} on the standard representation as in the previous case, independently of whether $d=1$ or not.
+Actually, all classical isotropic reductive group schemes can be described in terms of involution algebras [25].
+Exceptional isotropic irreducible Tits indices are given in the following table.
 Sizes and types of preimages are given by increasing the root length.
 The indices of labels in $J$ denote the lengths of the corresponding roots in $\Phi$, namely, long, short, and ultrashort ones.
-Tits index $\Phi$ $J$ $u^{-1}(0)$ $|u^{-1}(\alpha)|$ $u^{-1}(\mathbb{Z}\alpha\cap(\Phi\cup\{0\}))$
-${{}^{3}\!{\mathsf{D}_{4,0}^{28}}}$, ${{}^{6}\!{\mathsf{D}_{4,0}^{28}}}$ 00 $\varnothing$ $\mathsf{D}_{4}$
-${{}^{1}\!{\mathsf{E}_{6,0}^{78}}}$, ${{}^{2}\!{\mathsf{E}_{6,0}^{78}}}$ 00 $\varnothing$ $\mathsf{E}_{6}$
-$\mathsf{E}_{7,0}^{133}$ 00 $\varnothing$ $\mathsf{E}_{7}$
-$\mathsf{E}_{8,0}^{248}$ 00 $\varnothing$ $\mathsf{E}_{8}$
-$\mathsf{F}_{4,0}^{52}$ 00 $\varnothing$ $\mathsf{F}_{4}$
-$\mathsf{G}_{2,0}^{14}$ 00 $\varnothing$ $\mathsf{G}_{2}$
-$\mathsf{E}_{7,1}^{78}$ $\mathsf{A}_{1}$$\{7\}$ $\mathsf{E}_{6}$ $27$ $\mathsf{E}_{7}$
-${{}^{3}\!{\mathsf{D}_{4,1}^{9}}}$, ${{}^{6}\!{\mathsf{D}_{4,1}^{9}}}$ $\mathsf{BC}_{1}$$\{2\}$ $3\mathsf{A}_{1}$ $8$, $1$ $\mathsf{D}_{4}$, $4\mathsf{A}_{1}$
-${{}^{2}\!{\mathsf{E}_{6,1}^{35}}}$ $\mathsf{BC}_{1}$$\{2\}$ $\mathsf{A}_{5}$ $20$, $1$ $\mathsf{E}_{6}$, $\mathsf{A}_{1}+\mathsf{A}_{5}$
-$\mathsf{E}_{7,1}^{66}$ $\mathsf{BC}_{1}$$\{1\}$ $\mathsf{D}_{6}$ $32$, $1$ $\mathsf{E}_{7}$, $\mathsf{A}_{1}+\mathsf{D}_{6}$
-$\mathsf{E}_{8,1}^{133}$ $\mathsf{BC}_{1}$$\{8\}$ $\mathsf{E}_{7}$ $56$, $1$ $\mathsf{E}_{8}$, $\mathsf{A}_{1}+\mathsf{E}_{7}$
-$\mathsf{F}_{4,1}^{21}$ $\mathsf{BC}_{1}$$\{4\}$ $\mathsf{B}_{3}$ $8$, $7$ $\mathsf{F}_{4}$, $\mathsf{B}_{4}$
-${{}^{2}\!{\mathsf{E}_{6,1}^{29}}}$ $\mathsf{BC}_{1}$$\{1,6\}$ $\mathsf{D}_{4}$ $16$, $8$ $\mathsf{E}_{6}$, $\mathsf{D}_{5}$
-$\mathsf{E}_{7,1}^{48}$ $\mathsf{BC}_{1}$$\{6\}$ $\mathsf{A}_{1}+\mathsf{D}_{5}$ $32$, $10$ $\mathsf{E}_{7}$, $\mathsf{D}_{6}$
-$\mathsf{E}_{8,1}^{91}$ $\mathsf{BC}_{1}$$\{1\}$ $\mathsf{D}_{7}$ $64$, $14$ $\mathsf{E}_{8}$, $\mathsf{D}_{8}$
-${{}^{1}\!{\mathsf{E}_{6,2}^{28}}}$ $\mathsf{A}_{2}$$\{1,6\}$ $\mathsf{D}_{4}$ $8$ $\mathsf{D}_{5}$
-$\mathsf{G}_{2,2}^{0}$ $\mathsf{G}_{2}$$\{1_{\mathrm{s}},2_{\mathrm{l}}\}$varnothing $1$, $1$ $\mathsf{A}_{1}$, $\mathsf{A}_{1}$
-${{}^{3}\!{\mathsf{D}_{4,2}^{2}}}$, ${{}^{6}\!{\mathsf{D}_{4,2}^{2}}}$ $\mathsf{G}_{2}$$\{1_{\mathrm{s}},2_{\mathrm{l}},3_{\mathrm{s}},4_{\mathrm{s}}\}$varnothing $3$, $1$ $3\mathsf{A}_{1}$, $\mathsf{A}_{1}$
-${{}^{1}\!{\mathsf{E}_{6,2}^{16}}}$, ${{}^{2}\!{\mathsf{E}_{6,2}^{16^{\prime\prime}}}}$ $\mathsf{G}_{2}$$\{2_{\mathrm{l}},4_{\mathrm{s}}\}$ $2\mathsf{A}_{2}$ $9$, $1$ $\mathsf{A}_{5}$, $\mathsf{A}_{1}+2\mathsf{A}_{2}$
-$\mathsf{E}_{8,2}^{72}$ $\mathsf{G}_{2}$$\{7_{\mathrm{s}},8_{\mathrm{l}}\}$ $\mathsf{E}_{6}$ $27$, $1$ $\mathsf{E}_{7}$, $\mathsf{A}_{1}+\mathsf{E}_{6}$
-${{}^{2}\!{\mathsf{E}_{6,2}^{16^{\prime}}}}$ $\mathsf{BC}_{2}$$\{1_{\mathrm{us}},2_{\mathrm{s}},6_{\mathrm{us}}\}$ $\mathsf{A}_{3}$ $8$, $6$, $1$ $\mathsf{A}_{5}$, $\mathsf{D}_{4}$, $2\mathsf{A}_{1}+\mathsf{A}_{3}$
-$\mathsf{E}_{7,2}^{31}$ $\mathsf{BC}_{2}$$\{1_{\mathrm{s}},6_{\mathrm{us}}\}$ $\mathsf{A}_{1}+\mathsf{D}_{4}$ $16$, $8$, $1$ $\mathsf{D}_{6}$, $\mathsf{A}_{1}+\mathsf{D}_{5}$, $2\mathsf{A}_{1}+\mathsf{D}_{4}$
-$\mathsf{E}_{8,2}^{66}$ $\mathsf{BC}_{2}$$\{1_{\mathrm{us}},8_{\mathrm{s}}\}$ $\mathsf{D}_{6}$ $32$, $12$, $1$ $\mathsf{E}_{7}$, $\mathsf{D}_{7}$, $\mathsf{A}_{1}+\mathsf{D}_{6}$
-$\mathsf{E}_{7,3}^{28}$ $\mathsf{C}_{3}$$\{1_{\mathrm{s}},6_{\mathrm{s}},7_{\mathrm{l}}\}$ $\mathsf{D}_{4}$ $8$, $1$ $\mathsf{D}_{5}$, $\mathsf{A}_{1}+\mathsf{D}_{4}$
-$\mathsf{F}_{4,4}^{0}$ $\mathsf{F}_{4}$$\{1_{\mathrm{l}},2_{\mathrm{l}},3_{\mathrm{s}},4_{\mathrm{s}}\}$varnothing $1$, $1$ $\mathsf{A}_{1}$, $\mathsf{A}_{1}$
-${{}^{2}\!{\mathsf{E}_{6,4}^{2}}}$ $\mathsf{F}_{4}$$\{1_{\mathrm{s}},2_{\mathrm{l}},3_{\mathrm{s}},4_{\mathrm{l}},5_{\mathrm{s}},6_{\mathrm{s}}\}$varnothing $2$, $1$ $2\mathsf{A}_{1}$, $\mathsf{A}_{1}$
-$\mathsf{E}_{7,4}^{9}$ $\mathsf{F}_{4}$$\{1_{\mathrm{l}},3_{\mathrm{l}},4_{\mathrm{s}},6_{\mathrm{s}}\}$ $3\mathsf{A}_{1}$ $4$, $1$ $\mathsf{A}_{1}+\mathsf{A}_{3}$, $4\mathsf{A}_{1}$
-$\mathsf{E}_{8,4}^{28}$ $\mathsf{F}_{4}$$\{1_{\mathrm{s}},6_{\mathrm{s}},7_{\mathrm{l}},8_{\mathrm{l}}\}$ $\mathsf{D}_{4}$ $8$, $1$ $\mathsf{D}_{5}$, $\mathsf{A}_{1}+\mathsf{D}_{4}$
-${{}^{1}\!{\mathsf{E}_{6,6}^{0}}}$ $\mathsf{E}_{6}$$\{1,2,3,4,5,6\}$varnothing $1$ $\mathsf{A}_{1}$
-$\mathsf{E}_{7,7}^{0}$ $\mathsf{E}_{7}$$\{1,2,3,4,5,6,7\}$varnothing $1$ $\mathsf{A}_{1}$
-$\mathsf{E}_{8,8}^{0}$ $\mathsf{E}_{8}$$\{1,2,3,4,5,6,7,8\}$varnothing $1$ $\mathsf{A}_{1}$
+When working with these group schemes it is useful to use explicit root diagrams of all exceptional root systems [22], such diagrams split into the subdiagram $u^{-1}(0)$ and blocks $u^{-1}(\alpha)$ connected by edges and outer automorphisms of the Tits index.
+Tits index $\widetilde{\Phi}$ $J$ $\Phi$ $u^{-1}(0)$ $|u^{-1}(\alpha)|$ $|\mathrm{Out}|$
+$\mathsf{E}_{7,1}^{78}$ $\mathsf{E}_{7}$ $\{7\}$ $\mathsf{A}_{1}$ $\mathsf{E}_{6}$ $27$ 11
+$^{3}\!{\mathsf{D}_{4,1}^{9}}$}, $^{6}\!{\mathsf{D}_{4,1}^{9}}$} $\mathsf{D}_{4}$ $\{2\}$ $\mathsf{BC}_{1}$ $3\mathsf{A}_{1}$ 88, 11 66
+$^{2}\!{\mathsf{E}_{6,1}^{35}}$} $\mathsf{E}_{6}$ $\{2\}$ $\mathsf{BC}_{1}$ $\mathsf{A}_{5}$ 2020, 11 22
+$\mathsf{E}_{7,1}^{66}$ $\mathsf{E}_{7}$ $\{1\}$ $\mathsf{BC}_{1}$ $\mathsf{D}_{6}$ $32$, 11 11
+$\mathsf{E}_{8,1}^{133}$ $\mathsf{E}_{8}$ $\{8\}$ $\mathsf{BC}_{1}$ $\mathsf{E}_{7}$ $56$, 11 11
+$\mathsf{F}_{4,1}^{21}$ $\mathsf{F}_{4}$ $\{4\}$ $\mathsf{BC}_{1}$ $\mathsf{B}_{3}$ 88, 77 11
+$^{2}\!{\mathsf{E}_{6,1}^{29}}$} $\mathsf{E}_{6}$ $\{1,6\}$ $\mathsf{BC}_{1}$ $\mathsf{D}_{4}$ $16$, 88 22
+$\mathsf{E}_{7,1}^{48}$ $\mathsf{E}_{7}$ $\{6\}$ $\mathsf{BC}_{1}$ $\mathsf{A}_{1}+\mathsf{D}_{5}$ $32$, $10$ 11
+$\mathsf{E}_{8,1}^{91}$ $\mathsf{E}_{8}$ $\{1\}$ $\mathsf{BC}_{1}$ $\mathsf{D}_{7}$ $64$, $14$ 11
+$^{1}\!{\mathsf{E}_{6,2}^{28}}$} $\mathsf{E}_{6}$ $\{1,6\}$ $\mathsf{A}_{2}$ $\mathsf{D}_{4}$ 88 11
+$\mathsf{G}_{2,2}^{0}$ $\mathsf{G}_{2}$ $\{1_{\mathrm{s}},2_{\mathrm{l}}\}$ $\mathsf{G}_{2}$ $\varnothing$ 11, 11 11
+$^{3}\!{\mathsf{D}_{4,2}^{2}}$}, $^{6}\!{\mathsf{D}_{4,2}^{2}}$} $\mathsf{D}_{4}$ $\{1_{\mathrm{s}},2_{\mathrm{l}},3_{\mathrm{s}},4_{\mathrm{s}}\}$ $\mathsf{G}_{2}$ $\varnothing$ 33, 11 66
+$^{1}\!{\mathsf{E}_{6,2}^{16}}$}, $^{2}\!{\mathsf{E}_{6,2}^{16^{\prime\prime}}}$} $\mathsf{E}_{6}$ $\{2_{\mathrm{l}},4_{\mathrm{s}}\}$ $\mathsf{G}_{2}$ $2\mathsf{A}_{2}$ 99, 11 22
+$\mathsf{E}_{8,2}^{78}$ $\mathsf{E}_{8}$ $\{7_{\mathrm{s}},8_{\mathrm{l}}\}$ $\mathsf{G}_{2}$ $\mathsf{E}_{6}$ $27$, 11 11
+$^{2}\!{\mathsf{E}_{6,2}^{16^{\prime}}}$} $\mathsf{E}_{6}$ $\{1_{\mathrm{us}},2_{\mathrm{s}},6_{\mathrm{us}}\}$ $\mathsf{BC}_{2}$ $\mathsf{A}_{3}$ 88, 66, 11 22
+$\mathsf{E}_{7,2}^{31}$ $\mathsf{E}_{7}$ $\{1_{\mathrm{s}},6_{\mathrm{us}}\}$ $\mathsf{BC}_{2}$ $\mathsf{A}_{1}+\mathsf{D}_{4}$ $16$, 88, 11 11
+$\mathsf{E}_{8,2}^{66}$ $\mathsf{E}_{8}$ $\{1_{\mathrm{us}},8_{\mathrm{s}}\}$ $\mathsf{BC}_{2}$ $\mathsf{D}_{6}$ $32$, $12$, 11 11
+$\mathsf{E}_{7,3}^{28}$ $\mathsf{E}_{7}$ $\{1_{\mathrm{s}},6_{\mathrm{s}},7_{\mathrm{l}}\}$ $\mathsf{C}_{3}$ $\mathsf{D}_{4}$ 88, 11 11
+$\mathsf{F}_{4,4}^{0}$ $\mathsf{F}_{4}$ $\{1_{\mathrm{l}},2_{\mathrm{l}},3_{\mathrm{s}},4_{\mathrm{s}}\}$ $\mathsf{F}_{4}$ $\varnothing$ 11, 11 11
+$^{2}\!{\mathsf{E}_{6,4}^{2}}$} $\mathsf{E}_{6}$ $\{1_{\mathrm{s}},2_{\mathrm{l}},3_{\mathrm{s}},4_{\mathrm{l}},5_{\mathrm{s}},6_{\mathrm{s}}\}$ $\mathsf{F}_{4}$ $\varnothing$ 22, 11 22
+$\mathsf{E}_{7,4}^{9}$ $\mathsf{E}_{7}$ $\{1_{\mathrm{l}},3_{\mathrm{l}},4_{\mathrm{s}},6_{\mathrm{s}}\}$ $\mathsf{F}_{4}$ $3\mathsf{A}_{1}$ 44, 11 11
+$\mathsf{E}_{8,4}^{28}$ $\mathsf{E}_{8}$ $\{1_{\mathrm{s}},6_{\mathrm{s}},7_{\mathrm{l}},8_{\mathrm{l}}\}$ $\mathsf{F}_{4}$ $\mathsf{D}_{4}$ 88, 11 11
+$^{1}\!{\mathsf{E}_{6,6}^{0}}$} $\mathsf{E}_{6}$ $\{1,2,3,4,5,6\}$ $\mathsf{E}_{6}$ $\varnothing$ 11 11
+$\mathsf{E}_{7,7}^{0}$ $\mathsf{E}_{7}$ $\{1,2,3,4,5,6,7\}$ $\mathsf{E}_{7}$ $\varnothing$ 11 11
+$\mathsf{E}_{8,8}^{0}$ $\mathsf{E}_{8}$ $\{1,2,3,4,5,6,7,8\}$ $\mathsf{E}_{8}$ $\varnothing$ 11 11
+Finally, we give the complete list of isomorphisms and equivalences between isotropic irreducible Tits indices.
+Two such indices are called isomorphic (denoted by $\cong$) if they differ by an automorphism of the Dynkin diagram of $\widetilde{\Phi}$ and, possibly, by its renaming via an exceptional isomorphism.
+We say that two indices are equivalent (denoted by $\sim$) if the induced maps $u$ differ by isomorphisms between the corresponding root systems $\widetilde{\Phi}$ and $\Phi$.
+Equivalent Tits indices are indistinguishable for our purposes.
+${{}^{1}\!{\mathsf{A}_{1,1}^{(1)}}}$ $\cong\mathsf{B}_{1,1}\cong\mathsf{C}_{1,1}^{(1)},$ ${{}^{1}\!{\mathsf{D}_{4,1}^{(1)}}}$ $\cong{{}^{1}\!{\mathsf{D}_{4,1}^{(4)}}}\sim{{}^{2}\!{\mathsf{D}_{4,1}^{(1)}}},$
+$\mathsf{B}_{2,1}$ $\cong\mathsf{C}_{2,1}^{(2)},$ ${{}^{1}\!{\mathsf{D}_{4,1}^{(2)}}}$ $\sim{{}^{2}\!{\mathsf{D}_{4,1}^{(2)}}}\sim{{}^{3}\!{\mathsf{D}_{4,1}^{9}}}\sim{{}^{6}\!{\mathsf{D}_{4,1}^{9}}},$
+$\mathsf{B}_{2,2}$ $\cong\mathsf{C}_{2,2}^{(1)},$ ${{}^{1}\!{\mathsf{D}_{4,2}^{(1)}}}$ $\cong{{}^{1}\!{\mathsf{D}_{4,2}^{(2)}}}\sim{{}^{2}\!{\mathsf{D}_{4,2}^{(1)}}},$
+${{}^{2}\!{\mathsf{A}_{3,1}^{(1)}}}$ $\cong{{}^{2}\!{\mathsf{D}_{3,1}^{(2)}}},$ ${{}^{3}\!{\mathsf{D}_{4,2}^{2}}}$ $\sim{{}^{6}\!{\mathsf{D}_{4,2}^{2}}},$
+${{}^{2}\!{\mathsf{A}_{3,1}^{(2)}}}$ $\cong{{}^{2}\!{\mathsf{D}_{3,1}^{(1)}}}\sim{{}^{1}\!{\mathsf{A}_{3,1}^{(2)}}}\cong{{}^{1}\!{\mathsf{D}_{3,1}^{(1)}}},$ ${{}^{1}\!{\mathsf{E}_{6,2}^{16}}}$ $\sim{{}^{2}\!{\mathsf{E}_{6,2}^{16^{\prime\prime}}}},$
+${{}^{2}\!{\mathsf{A}_{3,2}^{(1)}}}$ $\cong{{}^{2}\!{\mathsf{D}_{3,2}^{(1)}}},$ ${{}^{1}\!{\mathsf{A}_{n,1}^{(d)}}}$ $\sim{{}^{2}\!{\mathsf{A}_{n,1}^{(d)}}}\text{ for }d\geq 3,\,n\geq 5,\,2d=n+1,$
+${{}^{1}\!{\mathsf{A}_{3,3}^{(1)}}}$ $\cong{{}^{1}\!{\mathsf{D}_{3,3}^{(1)}}},$ ${{}^{1}\!{\mathsf{D}_{n,r}^{(d)}}}$ $\sim{{}^{2}\!{\mathsf{D}_{n,r}^{(d)}}}\text{ for }d=2^{k}\mid 2n,\,d\geq 1,\,r\geq 1,\,n\geq 5,\,rd\leq n-2.$
 3 Technical lemmas
@@
   • saturated if $\Sigma=\Phi\cap\mathbb{R}_{\geq 0}\Sigma$.
-Note that any parabolic set is saturated (and every saturated set is an intersection of parabolic ones), but $\mathsf{A}_{2}\subseteq\mathsf{G}_{2}$ is not saturated.
+Note that any parabolic set is saturated (and every saturated set is an intersection of parabolic ones), but $\mathsf{A}_{2}\subseteq\mathsf{G}_{2}$ and $\mathsf{C}_{2}\subseteq\mathsf{BC}_{2}$ are not saturated.
 Any closed subset $\Sigma$ admits a unique decomposition $\Sigma=\Sigma_{\mathrm{r}}\sqcup\Sigma_{\mathrm{u}}$, where $\Sigma_{\mathrm{r}}=\Sigma\cap(-\Sigma)$ is a closed root subsystem and $\Sigma_{\mathrm{u}}=\Sigma\setminus(-\Sigma)$ is a unipotent set.
 Conversely, if $\Sigma_{\mathrm{r}}$ is a closed root subsystem, $\Sigma_{\mathrm{u}}$ is a unipotent set disjoint with $\Sigma_{\mathrm{r}}$, and $(\Sigma_{\mathrm{r}}+\Sigma_{\mathrm{u}})\cap\Phi\subseteq\Sigma_{\mathrm{u}}$, then $\Sigma_{\mathrm{r}}\cup\Sigma_{\mathrm{u}}$ is a closed set.
-A closed set $\Sigma$ is parabolic if and only if there are $w\in\mathrm{W}(\Phi)$ and $J\subseteq\Delta$ such that $w\Sigma=(\Pi+\mathbb{Z}J)\cap\Phi$.
-Here $\Pi\subseteq\Phi$ is the set of positive roots and $\Delta\subseteq\Pi$ is the set of basic roots.
+A closed set $\Sigma$ is parabolic if and only if there are $w\in\mathrm{W}(\Phi)$ and $J\subseteq\Delta$ such that $w\Sigma=(\Phi^{+}+\mathbb{Z}J)\cap\Phi$.
+Here $\Phi^{+}\subseteq\Phi$ is the set of positive roots and $\Delta\subseteq\Phi^{+}$ is the set of basic roots.
 The classes of closed sets, unipotent sets, closed root subsystems, and saturated sets are closed under intersection.
@@
   • if $\Sigma$ is saturated, then $u^{-1}(\Sigma\cup\{0\})$ is saturated.
-If $\Sigma\subseteq\Phi$ is a closed root subsystem, then the smallest closed set containing $u^{-1}(\Sigma)$ consists of $u^{-1}(\Sigma)$ and sums of pairs of roots of $u^{-1}(\Sigma)$ (with opposite images in $\Sigma$), it is also a closed root subsystem.
-For any closed subset $\Sigma\subseteq\Phi$ let $G_{\Sigma}\leq G$ and $G^{0}_{\Sigma}\leq G$ be the group subsheaves generated by $\bigcup_{\alpha\in\Sigma}U_{\alpha}$ and $L\cup\bigcup_{\alpha\in\Sigma}U_{\alpha}$, they are actually smooth closed group subschemes with connected geometric fibers over $K$ and $G^{0}_{\Sigma}$ is of type (RC) [3, §XXII.5].
+If $\Sigma\subseteq\Phi$ is a closed root subsystem, then the smallest closed set containing $u^{-1}(\Sigma)$ is also a closed root subsystem.
+For any closed subset $\Sigma\subseteq\Phi$ let $G_{\Sigma}\leq G$ and $G^{0}_{\Sigma}\leq G$ be the group subsheaves generated by $\bigcup_{\alpha\in\Sigma}U_{\alpha}$ and $L\cup\bigcup_{\alpha\in\Sigma}U_{\alpha}$, they are actually smooth closed group subschemes with connected geometric fibers over $K$ and $G^{0}_{\Sigma}$ is of type (RC) [13, §XXII.5].
 More precisely, if $\Sigma$ is a root subsystem, then $G_{\Sigma}$ and $G^{0}_{\Sigma}$ are reductive group subschemes (with induced isotropic structure).
@@
 it is an étale twisted form of some $\mathbb{A}^{N}$ as a scheme.
-Finally, in general $G_{\Sigma}=G_{\Sigma_{\mathrm{r}}}\rtimes G_{\Sigma_{\mathrm{u}}}$, and $G^{0}_{\Sigma}=G^{0}_{\Sigma_{\mathrm{r}}}\rtimes G_{\Sigma_{\mathrm{u}}}$.
+Finally, in general $G_{\Sigma}=G_{\Sigma_{\mathrm{r}}}\rtimes G_{\Sigma_{\mathrm{u}}}$ and $G^{0}_{\Sigma}=G^{0}_{\Sigma_{\mathrm{r}}}\rtimes G_{\Sigma_{\mathrm{u}}}$.
 If $\Sigma$ is parabolic, then $G^{0}_{\Sigma}$ is a parabolic group subscheme.
@@
 Suppose that $K$ is semi-local. Then every isotropic reductive group $G$ has a Gauss decomposition
-$G(K)=G_{\Pi}(K)\,G_{-\Pi}(K)\,G_{\Pi}(K)\,L(K).$
+$G(K)=G_{\Phi^{+}}(K)\,G_{\Phi^{-}}(K)\,G_{\Phi^{+}}(K)\,L(K).$
 Moreover, for any closed subset $\Sigma\subseteq\Phi$ there is a decomposition
-$G^{0}_{\Sigma}(K)=G_{\Sigma_{\mathrm{r}}\cap\Pi}(K)\,G_{\Sigma\cap(-\Pi)}(K)\,G_{\Sigma\cap\Pi}(K)\,L(K).$
+$G^{0}_{\Sigma}(K)=G_{\Sigma_{\mathrm{r}}\cap\Phi^{+}}(K)\,G_{\Sigma\cap\Phi^{-}}(K)\,G_{\Sigma\cap\Phi^{+}}(K)\,L(K).$
 Proof.
 We may assume that $\operatorname{Spec}(K)$ is connected.
-In this case there is a maximal isotropic structure $(L^{\prime},G^{\prime}_{\alpha})_{\alpha\in\Phi^{\prime}}$ and a map $v\colon\Phi^{\prime}\to\Phi\cup\{0\}$ such that $\Phi\subseteq v(\Phi^{\prime})$, $G^{\prime}_{\alpha}=G_{v^{-1}(\alpha)}$, $L^{\prime}=G_{\Phi\cap v^{-1}(0)}$ [3, XXVI.7.4.2].
-Gauss decomposition holds for $G$ with respect to a maximal isotropic structure by [3, corollary XXVI.5.2], i.e.
-$G(K)=G^{\prime}_{\Pi^{\prime}}(K)\,G^{\prime}_{-\Pi^{\prime}}(K)\,G^{\prime}_{\Pi^{\prime}}(K)\,L^{\prime}(K).$
+In this case there is a maximal isotropic structure $(L^{\prime},G^{\prime}_{\alpha})_{\alpha\in\Phi^{\prime}}$ and a map $v\colon\Phi^{\prime}\to\Phi\cup\{0\}$ such that $\Phi\subseteq v(\Phi^{\prime})$, $G^{\prime}_{\alpha}=G_{v^{-1}(\alpha)}$, $L^{\prime}=G_{\Phi\cap v^{-1}(0)}$ [13, XXVI.7.4.2].
+Gauss decomposition holds for $G$ with respect to a maximal isotropic structure by [13, corollary XXVI.5.2], i.e.
+$G(K)=G^{\prime}_{\Phi^{\prime+}}(K)\,G^{\prime}_{\Phi^{\prime-}}(K)\,G^{\prime}_{\Phi^{\prime+}}(K)\,L^{\prime}(K).$
 It follows that
-$G(K)=G_{\Pi}(K)\,G_{-\Pi}(K)\,G_{\Pi}(K)\,L(K)$
+$G(K)=G_{\Phi^{+}}(K)\,G_{\Phi^{-}}(K)\,G_{\Phi^{+}}(K)\,L(K)$
 and
-$G^{0}_{\Sigma}(K)=G_{\Sigma\cap\Pi}(K)\,G_{\Sigma\cap(-\Pi)}(K)\,G_{\Sigma\cap\Pi}(K)\,L(K)$
+$G^{0}_{\Sigma}(K)=G_{\Sigma\cap\Phi^{+}}(K)\,G_{\Sigma\cap\Phi^{-}}(K)\,G_{\Sigma\cap\Phi^{+}}(K)\,L(K)$
 for any root subsystem $\Sigma$. If $\Sigma\subseteq\Phi$ is a closed subset, then
-$G^{0}_{\Sigma}(K)$ $=G_{\Sigma_{\mathrm{r}}\cap\Pi}(K)\,G_{\Sigma_{\mathrm{r}}\cap(-\Pi)}(K)\,G_{\Sigma_{\mathrm{u}}}(K)\,G_{\Sigma_{\mathrm{r}}\cap\Pi}(K)\,L(K)$
-$=G_{\Sigma_{\mathrm{r}}\cap\Pi}(K)\,G_{\Sigma\cap(-\Pi)}(K)\,G_{\Sigma\cap\Pi}(K)\,L(K).\qed$
-For any closed subsets $\Sigma,\Sigma^{\prime}\subseteq\Phi$ the intersection $G^{0}_{\Sigma}\cap G^{0}_{\Sigma^{\prime}}$ is a smooth closed subscheme and its fiberwise connected component is $G^{0}_{\Sigma\cap\Sigma^{\prime}}$ by [3, proposition XXII.5.4.5].
+$G^{0}_{\Sigma}(K)$ $=G_{\Sigma_{\mathrm{r}}\cap\Phi^{+}}(K)\,G_{\Sigma_{\mathrm{r}}\cap\Phi^{-}}(K)\,G_{\Sigma_{\mathrm{u}}}(K)\,G_{\Sigma_{\mathrm{r}}\cap\Phi^{+}}(K)\,L(K)$
+$=G_{\Sigma_{\mathrm{r}}\cap\Phi^{+}}(K)\,G_{\Sigma\cap\Phi^{-}}(K)\,G_{\Sigma\cap\Phi^{+}}(K)\,L(K).\qed$
+For any closed subsets $\Sigma,\Sigma^{\prime}\subseteq\Phi$ the intersection $G^{0}_{\Sigma}\cap G^{0}_{\Sigma^{\prime}}$ is a smooth closed subscheme and its fiberwise connected component is $G^{0}_{\Sigma\cap\Sigma^{\prime}}$ by [13, proposition XXII.5.4.5].
 Lemma 2.
@@
 Proof.
-For the first claim it suffices to prove that $G^{0}_{\Sigma}\cap G^{0}_{\Sigma^{\prime}}=G^{0}_{\Sigma\cap\Sigma^{\prime}}$ if $\Sigma^{\prime}\supseteq\Pi$ is parabolic and $K$ is local.
+For the first claim it suffices to prove that $G^{0}_{\Sigma}\cap G^{0}_{\Sigma^{\prime}}=G^{0}_{\Sigma\cap\Sigma^{\prime}}$ if $\Sigma^{\prime}\supseteq\Phi^{+}$ is parabolic and $K$ is local.
 Take $g\in G^{0}_{\Sigma}(K)\cap G^{0}_{\Sigma^{\prime}}(K)$.
-By lemma 1 we may multiply $g$ from both sides by elements of $G^{0}_{\Sigma\cap\Sigma^{\prime}}$ to get a new element $g^{\prime}\in G_{\Sigma\cap(-\Sigma^{\prime}_{\mathrm{u}})}(K)\cap G^{0}_{\Sigma^{\prime}}(K)$.
-But such $g^{\prime}$ is necessarily trivial.
-Now let us prove the second claim. Without loss of generality $\Sigma^{\prime}\subseteq\Pi$, so
-$G^{0}_{\Sigma}\cap G_{\Sigma^{\prime}}=G^{0}_{\Sigma}\cap G^{0}_{\Pi}\cap G_{\Sigma^{\prime}}=G^{0}_{\Sigma\cap\Pi}\cap G_{\Sigma^{\prime}}=G_{\Sigma\cap\Pi\cap\Sigma^{\prime}}=G_{\Sigma\cap\Sigma^{\prime}}.\qed$
-Recall [8, §4] that a $2$-step $K$-module $(M,M_{0})$ consists of
+By lemma 1 we can multiply $g$ from both sides by elements of $G^{0}_{\Sigma\cap\Sigma^{\prime}}$ to get a new element $g^{\prime}\in G_{\Sigma\cap\Phi^{-}}(K)\cap G^{0}_{\Sigma^{\prime}}(K)$.
+But such $g^{\prime}$ necessarily lies in $G_{\Sigma\cap\Sigma^{\prime}\cap\Phi^{-}}$.
+Now let us prove the second claim. Without loss of generality $\Sigma^{\prime}\subseteq\Phi^{+}$, so
+$G^{0}_{\Sigma}\cap G_{\Sigma^{\prime}}=G^{0}_{\Sigma}\cap G^{0}_{\Phi^{+}}\cap G_{\Sigma^{\prime}}=G^{0}_{\Sigma\cap\Phi^{+}}\cap G_{\Sigma^{\prime}}=G_{\Sigma\cap\Phi^{+}\cap\Sigma^{\prime}}=G_{\Sigma\cap\Sigma^{\prime}}.\qed$
+Recall [25, §4] that a $2$-step $K$-module $(M,M_{0})$ consists of
 […]
-  • a central subgroup $M_{0}\leq M$ such that $[M,M]^{\cdot}\leq M_{0}$;
+  • a central subgroup $M_{0}\leq M$ such that $M,M]^{\cdot}\leq M_{0}$;
   • a left $K$-module structure on $M_{0}$;
-  • a right action $({-})\cdot({=})\colon M\times K\to M$ of the multiplicative monoid by group automorphisms such that $[m\cdot k,m^{\prime}\cdot k^{\prime}]^{\cdot}=kk^{\prime}[m,m^{\prime}]^{\cdot}$, $m\cdot(k+k^{\prime})=m\cdot k\dotplus kk^{\prime}\tau(m)\dotplus m\cdot k^{\prime}$ for some (uniquely determined) $\tau(m)\in M_{0}$, $m_{0}\cdot k=k^{2}m_{0}$ for $m_{0}\in M_{0}$.
+  • a right action $({-})\cdot({=})\colon M\times K\to M$ of the multiplicative monoid by group automorphisms such that $m\cdot k,m^{\prime}\cdot k^{\prime}]^{\cdot}=kk^{\prime}[m,m^{\prime}]^{\cdot}$, $m\cdot(k+k^{\prime})=m\cdot k\dotplus kk^{\prime}\tau(m)\dotplus m\cdot k^{\prime}$ for some (uniquely determined) $\tau(m)\in M_{0}$, $m_{0}\cdot k=k^{2}m_{0}$ for $m_{0}\in M_{0}$.
 Then
@@
 $\tau(\dot{0})$ $=\dot{0},$ $\tau(m\dotplus m^{\prime})$ $=\tau(m)+[m,m^{\prime}]^{\cdot}+\tau(m^{\prime}),$
-$\tau(\mathbin{\dot{-}}m)$ =˙$=\mathbin{\dot{-}}\tau(m),$ $\tau(m\cdot k)$ $=k^{2}\tau(m),$
+$\tau(\mathbin{\dot{-}}m)$ $=\mathbin{\dot{-}}\tau(m),$ $\tau(m\cdot k)$ $=k^{2}\tau(m),$
 $\tau(m_{0})$ $=2m_{0}\text{ for }m_{0}\in M_{0}.$
-A pair of subsets $(X,X_{0})\subseteq(M,M_{0})$ generates $(M,M_{0})$ if $X_{0}$ generates the $K$-module $M_{0}$ and $X$ generates the $K$-module $M/M_{0}$.
+A pair of subsets $(X,X_{0})\subseteq(M,M_{0})$ (i.e.
+$X\subseteq M$ and $X_{0}\subseteq M_{0}$) generates $(M,M_{0})$ if $X_{0}$ generates the $K$-module $M_{0}$ and $X$ generates the $K$-module $M/M_{0}$.
 We say that a $2$-step nilpotent $K$-module $(M,M_{0})$ is locally free if $M_{0}$ and $M/M_{0}$ are finitely generated projective $K$-modules.
@@
 $\tau(m_{0}\mathbin{\dot{\oplus}}m_{1})$ $=(2m_{0}-c(m_{1},m_{1}))\mathbin{\dot{\oplus}}0.$
-Every locally free $2$-step nilpotent $K$-module $(M,M_{0})$ determines a representable fpqc sheaf $\mathbb{M}_{0}(E)=E\otimes_{K}M_{0}$, $\mathbb{M}(E)=M\boxtimes E=E\otimes_{K}M_{0}\mathbin{\dot{\oplus}}E\otimes_{K}M_{1}$ of locally free $2$-step nilpotent modules (where $K\to E$ is a ring homomorphism), the latter one is independent on the choice of the splitting [8, §4].
+Every locally free $2$-step nilpotent $K$-module $(M,M_{0})$ determines a representable fpqc sheaf $\mathbb{M}_{0}(E)=E\otimes_{K}M_{0}$, $\mathbb{M}(E)=M\boxtimes E=E\otimes_{K}M_{0}\mathbin{\dot{\oplus}}E\otimes_{K}M_{1}$ of locally free $2$-step nilpotent modules (where $K\to E$ is a ring homomorphism), the latter one is independent on the choice of the splitting [25, §4].
 Also, locally free $2$-step nilpotent $K$-modules satisfy the fpqc descent.
 If $G$ is a group scheme acting on such a sheaf $(\mathbb{M},\mathbb{M}_{0})$ by automorphisms of $2$-step nilpotent modules and $G$ stabilizes some generating set $(X,X_{0})\subseteq(M,M_{0})$, then the action is trivial.
-By [9, §2] there are canonical homomorphisms $t_{\alpha}\colon\mathfrak{g}_{\alpha}\to G_{\alpha}(K)$ for non-ultrashort $\alpha\in\Phi$ and $t_{\alpha}\colon\mathfrak{g}_{\alpha}\mathbin{\dot{\oplus}}\mathfrak{g}_{2\alpha}\to G_{\alpha}(K)$ for ultrashort $\alpha$.
-In the second case $\mathfrak{g}_{\alpha}\dotplus\mathfrak{g}_{2\alpha}$ is a locally free $2$-step nilpotent $K$-module.
-Since $G$ is root graded, the Lie bracket $\mathfrak{g}_{\alpha}\times\mathfrak{g}_{\beta}\to\mathfrak{g}_{\alpha+\beta}$ is non-degenerate on the second argument if $\alpha$ is long and $\frac{\pi}{2}<\angle(\alpha,\beta)<\pi$, i.e.
+By [26, §2] there are canonical homomorphisms $t_{\alpha}\colon\mathfrak{g}_{\alpha}\to G_{\alpha}(K)$ for non-ultrashort $\alpha\in\Phi$ and $t_{\alpha}\colon\mathfrak{g}_{\alpha}\mathbin{\dot{\oplus}}\mathfrak{g}_{2\alpha}\to G_{\alpha}(K)$ for ultrashort $\alpha$.
+In the second case $\mathfrak{g}_{\alpha}\mathbin{\dot{\oplus}}\mathfrak{g}_{2\alpha}$ means a locally free $2$-step nilpotent $K$-module with the canonical short exact sequence $0\to\mathfrak{g}_{\alpha}\to\mathfrak{g}_{\alpha}\mathbin{\dot{\oplus}}\mathfrak{g}_{2\alpha}\to\mathfrak{g}_{2\alpha}\to 0$, though its splitting is non-canonical.
+Since $G$ is root graded locally in Zariski topology (by our additional assumption on isotropic data), the Lie bracket $\mathfrak{g}_{\alpha}\times\mathfrak{g}_{\beta}\to\mathfrak{g}_{\alpha+\beta}$ is non-degenerate on the second argument if $\alpha$ is long and $\frac{\pi}{2}<\angle(\alpha,\beta)<\pi$, i.e.
 $\mathfrak{g}_{\beta}\to\mathrm{Hom}(\mathfrak{g}_{\alpha},\mathfrak{g}_{\alpha+\beta})$ is injective.
@@
 Proof.
-In the orthogonal-like case $\mathfrak{g}_{\alpha}\cong R$ for a ring $R$ (an Azumaya algebra over $K$ or its quadratic étale extension) and $\mathfrak{g}_{\beta}\cong M_{R}$ for some right faithfully projective $A$-module, the pairing is just $M\times R\to M,\,(m,r)\mapsto mr$ up to an isomorphism.
-In the symplectic-like case $\mathfrak{g}_{\beta}\cong R$ for a ring $R$ with a $\lambda$-involution and $\mathfrak{g}_{\alpha}\cong\Lambda$ for a form parameter $\Lambda\leq R$, the pairing is $R\times\Lambda\to R,\,(r,u)\mapsto ru$ up to an isomorphism.
-This pairing is also non-degenerate, the intersection $\Lambda\cap R^{*}$ is non-empty because $G$ is root graded.
-∎
+Without loss of generality, $G$ splits.
+For the Tits indices $\mathsf{B}_{n,2}$ and $^{k}\!{\mathsf{D}_{n,r}^{(1)}}$} we have $\mathfrak{g}_{\alpha}\cong K$ and $\mathfrak{g}_{\beta}\cong K^{n}$, the pairing is isomorphic to $K^{n}\times K\to K^{n},\,(m,x)\mapsto mx$.
+For the Tits indices $\mathsf{C}_{2d,2}^{(d)}$ or $^{1}\!{\mathsf{D}_{2d,2}^{(d)}}$} we have $\mathfrak{g}_{\beta}\cong\operatorname{M}(d,K)$ and
+$\mathfrak{g}_{\alpha}\cong\Lambda=\{x\in\operatorname{M}(d,K)\mid x^{\mathrm{t}}=\pm x\},$
+the pairing is isomorphic to $\operatorname{M}(d,K)\times\Lambda\to\operatorname{M}(d,K),\,(x,u)\mapsto xu$.
+This pairing is also non-degenerate because the sign - appears in $\Lambda$ for the Tits index $^{1}\!{\mathsf{D}_{2d,2}^{(d)}}$} and in this case $d\geq 2$.
+Finally, if the Tits index is $^{2}\!{\mathsf{A}_{4d-1,2}^{(d)}}$}, then $\mathfrak{g}_{\beta}\cong\operatorname{M}(d,K)\times\operatorname{M}(d,K)$, $\mathfrak{g}_{\alpha}\cong\operatorname{M}(d,K)$, and the pairing is isomorphic to
+$(\operatorname{M}(d,K)\times\operatorname{M}(d,K))\times\operatorname{M}(d,K)\to\operatorname{M}(d,K)\times\operatorname{M}(d,K),\,((x,y),u)\mapsto(xu,-uy).\qed$
 Lemma 4.
@@
 Proof.
-If $G$ is classical, then this follows from the non-degeneracy of the hermitian form on the defining representation.
-Otherwise there are the cases ${{}^{2}\!{\mathrm{E}_{6,2}^{16^{\prime}}}}$, $\mathrm{E}_{7,2}^{31}$, and $\mathsf{E}_{8,2}^{66}$.
-It is possible to check using the root diagrams of $\mathsf{E}_{\ell}$ that the required Lie bracket is non-degenerate.
-∎
+In the classical cases fppf locally the pairing is isomorphic to the product map
+$\operatorname{M}(d,n,K)\times\operatorname{M}(n,d,K)\to\operatorname{M}(d,K)$
+or to a pair of such maps. The exceptional cases $^{2}\!{\mathrm{E}_{6,2}^{16^{\prime}}}$}, $\mathrm{E}_{7,2}^{31}$, and $\mathsf{E}_{8,2}^{66}$ can be checked using [22]. ∎
 4 Some normalizers and centralizers
@@
 Let also $\alpha\in\Phi$ be a long root and $X\subseteq\mathfrak{g}_{\alpha}$ be a generating set of a $K$-module.
-Then the group scheme
+Then the group subscheme
 $N=\{g\in G\mid\forall x\in X\enskip{{}^{g}\!{t_{\alpha}(x)}}\in G_{\alpha}\}$
@@
 Clearly, $P\leq N$ and we may assume that $K$ is local.
-By lemma 1 it suffices to check that $G_{\Sigma}\cap N=1$, where $\Sigma=\{\beta\mid\angle(\alpha,\beta)>\frac{\pi}{2}\}$.
+By lemma 1 it suffices to check that $G_{\Sigma}(K)\cap N(K)=1$, where $\Sigma=\{\beta\mid\angle(\alpha,\beta)>\frac{\pi}{2}\}$.
 Choose a linear map $f\colon\mathbb{R}\Phi\to\mathbb{R}$ such that $\operatorname{Ker}(f)\cap\Phi=\mathbb{R}\alpha\cap\Phi$ and let $\Phi_{s}=\{\beta\in\Phi\mid\mathrm{sign}(f(\beta))=s\}$, $\Sigma_{s}=\Phi_{s}\cap\Sigma$ for $s\in\{{-},0,{+}\}$, so $\Phi_{\pm}$ are unipotent subsets.
-Take any element $g=g_{-}g_{0}g_{+}\in G_{\Sigma}\cap N$ with $g_{s}\in G_{\Sigma_{s}}$, then $g_{\pm}=\prod_{\beta\in\Sigma_{\pm}\setminus 2\Sigma_{\pm}}t_{\beta}(y_{\beta})$ for any linear orders on $\Sigma_{\pm}\setminus 2\Sigma_{\pm}$ and $g_{0}\in G_{-\alpha}$ (or $g_{0}\in G_{-\alpha/2}$ if $\Phi$ is of type $\mathsf{BC}_{\ell}$).
+Take any element $g=g_{-}g_{0}g_{+}\in G_{\Sigma}(K)\cap N(K)$ with $g_{s}\in G_{\Sigma_{s}}$, then $g_{\pm}=\prod_{\beta\in\Sigma_{\pm}\setminus 2\Sigma_{\pm}}t_{\beta}(y_{\beta})$ for any linear orders on $\Sigma_{\pm}\setminus 2\Sigma_{\pm}$ and $g_{0}\in U_{-\alpha}(K)$ (or $g_{0}\in U_{-\alpha/2}(K)$ if $\Phi$ is of type $\mathsf{BC}_{\ell}$).
 For any $x\in X$ there is $x^{\prime}\in\mathfrak{g}_{\alpha}$ such that $g\,t_{\alpha}(x)=t_{\alpha}(x^{\prime})\,g$.
-Comparing both sides of $g_{-}\,g_{0}\,t_{\alpha}(x)\,g_{+}^{t_{\alpha}(x)}={{}^{t_{\alpha}(x^{\prime})}\!{g_{-}}}\,t_{\alpha}(x^{\prime})\,g_{0}\,g_{+}$ we get $[g_{+},t_{\alpha}(x)]=1$ and $[g_{-},t_{\alpha}(x^{\prime})]=1$.
-Since $x$ and $x^{\prime}$ runs generating sets of $\mathfrak{g}_{\alpha}$ and $G$ is root graded (so the commutator map $\mathfrak{g}_{\alpha}\times\mathfrak{g}_{\beta}\to\mathfrak{g}_{\alpha+\beta}$ is non-degenerate on the second argument for $\beta\in\Sigma_{\pm}$) both $g_{+}$ and $g_{-}$ are trivial.
-It follows that $g=g_{0}\in G_{-\alpha^{\prime}}$ for $\alpha^{\prime}\in\{\alpha,\frac{\alpha}{2}\}$.
+Comparing both sides of $g_{-}\,g_{0}\,t_{\alpha}(x)\,g_{+}^{t_{\alpha}(x)}={{}^{t_{\alpha}(x^{\prime})}\!{g_{-}}}\,t_{\alpha}(x^{\prime})\,g_{0}\,g_{+}$ we get $g_{+},t_{\alpha}(x)]=1$ and $g_{-},t_{\alpha}(x^{\prime})]=1$.
+Since $x$ runs over a generating set of $\mathfrak{g}_{\alpha}$ and the commutator map $\mathfrak{g}_{\alpha}\times\mathfrak{g}_{\beta}\to\mathfrak{g}_{\alpha+\beta}$ is non-degenerate on the second argument for $\beta\in\Sigma_{\pm}$, $g_{+}$ is trivial, so $g_{-}$ is also trivial by changing the sign of $f$.
+It follows that $g=g_{0}\in U_{-\alpha^{\prime}}(K)$ for $\alpha^{\prime}\in\{\alpha,\frac{\alpha}{2}\}$.
 Now suppose that $\Phi$ is not of the type $\mathsf{BC}_{\ell}$ and take a root $\beta\in\Phi$ such that $\frac{\pi}{2}<\angle(\alpha,\beta)<\pi$ and $\beta$ is long if $\Phi$ is of the type $\mathsf{G}_{2}$.
 Conjugate $t_{\beta}(z)$ by both sides of $g\,t_{\alpha}(x)=t_{\alpha}(x^{\prime})\,g$.
-We get $\bigl{[}\log g,[x,z]\bigr{]}=1$, where $[{-},{=}]\colon\mathfrak{g}_{\alpha}\times\mathfrak{g}_{\beta}\to\mathfrak{g}_{\alpha+\beta}$ is the commutator map and $\log g\in\mathfrak{g}_{-\alpha}$ is the element such that $g=t_{-\alpha}(\log g)$.
-Since $G$ is root graded, $\mathfrak{g}_{\alpha+\beta}=[\mathfrak{g}_{\alpha},\mathfrak{g}_{\beta}]$, so $[\log g,\mathfrak{g}_{\alpha+\beta}]=0$ and $\log g=0$ (by lemma 3 if $\beta$ is short), i.e.
+We get $\bigl[\log g,[x,z]\bigr]=0$, where $\log g\in\mathfrak{g}_{-\alpha}$ is the element such that $g=t_{-\alpha}(\log g)$.
+Since $\mathfrak{g}_{\alpha+\beta}=[\mathfrak{g}_{\alpha},\mathfrak{g}_{\beta}]$, $\log g,\mathfrak{g}_{\alpha+\beta}]=0$ and $\log g=0$ (by lemma 3 if $\beta$ is short), i.e.
 $g=1$.
 Finally, suppose that $\Phi$ is of type $\mathsf{BC}_{\ell}$.
-Take an ultrashort root $\beta$ orthogonal to $\alpha$ and conjugate $t_{\beta}(u)$ by both sides of $g\,t_{\alpha}(x)=t_{\alpha}(x^{\prime})\,g$.
-We get $[t_{\alpha}(x^{\prime}),[g,t_{\beta}(u)]]=1$, so $[g,t_{\beta}(u)]=1$ and $g\in G_{-\alpha}$ by lemma 4.
-In other words, we reduce to the previous case of $\mathsf{C}_{\ell}\subseteq\mathsf{BC}_{\ell}$.
+Take an ultrashort root $\beta$ orthogonal to $\alpha$ and conjugate $t_{\beta}(u)$ by both sides of $t_{\alpha}(x)\,g^{-1}=g^{-1}\,t_{\alpha}(x^{\prime})$.
+We get $t_{\alpha}(x^{\prime}),[g^{-1},t_{\beta}(u)]]=1$, so $g^{-1},t_{\beta}(u)]=1$ and $g\in U_{-\alpha}(K)$ by lemma 4.
+In other words, we reduce to the already considered case of $\mathsf{C}_{\ell}\subseteq\mathsf{BC}_{\ell}$.
 ∎
@@
 is a sheaf of locally free $2$-step nilpotent modules.
-The action of the group scheme $G^{0}_{\Phi\cap\operatorname{Ker}(f)}$ on it commutes with the operations of $2$-step nilpotent modules, this may be easily checked by passing to a split form of $G$.
+The action of the group scheme $G^{0}_{\Phi\cap\operatorname{Ker}(f)}$ on it commutes with the operations of $2$-step nilpotent modules, this can be easily checked by passing to a split form of $G$.
 Lemma 5.
 Let $G$ be an isotropic reductive group scheme and $f\colon\mathbb{R}\Phi\to\mathbb{R}$ a linear map such that $f(\Phi)\subseteq\{-2,-1,0,1,2\}$ and $\Phi\cap f^{-1}(2)\neq\varnothing$. Then
-$C=\{g\in G^{0}_{\Phi_{0}}\mid[g,t_{\alpha}(x_{\alpha,i})]=1\text{ for }f(\alpha)>0\}$
-coincides with the scheme center $\mathrm{C}(G)\leq\mathrm{C}(L)$, where $\Phi_{0}=\Phi\cap\operatorname{Ker}(f)$ and $x_{\alpha,i}\in\mathfrak{g}_{\alpha}$ are $K$-module generators if $\alpha$ is not ultrashort and $x_{\alpha,i}\in\mathfrak{g}_{\alpha}\mathbin{\dot{\oplus}}\mathfrak{g}_{2\alpha}$ generate the $K$-module $\mathfrak{g}_{\alpha}$ otherwise.
+$C=\bigl\{g\in G^{0}_{\Phi\cap\operatorname{Ker}(f)}\mid[g,t_{\alpha}(x_{\alpha,i})]=1\text{ for all }\alpha\text{ with }f(\alpha)>0\bigr\}$
+coincides with the scheme center $\mathrm{C}(G)\leq\mathrm{C}(L)$, $x_{\alpha,i}\in\mathfrak{g}_{\alpha}$ are $K$-module generators if $\alpha$ is not ultrashort and $x_{\alpha,i}\in\mathfrak{g}_{\alpha}\mathbin{\dot{\oplus}}\mathfrak{g}_{2\alpha}$ generate the $K$-module $\mathfrak{g}_{\alpha}$ otherwise.
 Proof.
-Clearly, $C$ is the scheme centralizer of $G_{\Phi_{>0}}$ in $G^{0}_{\Phi_{0}}$, where $\Phi_{>0}=\{\alpha\in\Phi\mid f(\alpha)>0\}$.
-We may assume that $G$ splits and $\Phi=\widetilde{\Phi}$.
+Clearly, $C$ is the scheme centralizer of $G_{\Phi_{>0}}$ in $G^{0}_{\Phi\cap\operatorname{Ker}(f)}$, where $\Phi_{>0}=\{\alpha\in\Phi\mid f(\alpha)>0\}$.
+We can assume that $G$ splits and $\Phi=\widetilde{\Phi}$.
 Firstly, let us show that $C$ is contained in the maximal torus $T=L$.
 By theorem 1 and lemma 2 the group scheme $C$ is contained in $G^{0}_{\Sigma}$, where $\Sigma$ consists of the roots $\alpha\in\Phi_{0}$ such that $\angle(\alpha,\beta)\leq\frac{\pi}{2}$ for all long roots $\beta\in\Phi_{>0}$.
-By considering rank $2$ root subsystems it may be shown that roots of $\Sigma$ are actually orthogonal to long roots from $\Phi_{>0}$.
-Now if the set of long roots of $\Phi$ is indecomposable root system (of full rank), then the set of long roots from $\Phi\setminus\Phi_{0}$ generates this root subsystem e.g.
-by [9, lemma 1].
+By considering rank $2$ root subsystems it follows that roots of $\Sigma$ are actually orthogonal to long roots from $\Phi_{>0}$.
+Now if the set of long roots of $\Phi$ is an indecomposable root system of full rank, then the set of long roots from $\Phi\setminus\operatorname{Ker}(f)$ generates this root subsystem e.g.
+by [26, lemma 1].
 In this case necessarily $\Sigma=\varnothing$.
 Otherwise $\Phi=\mathsf{C}_{\ell}$, possibly $\mathsf{B}_{2}=\mathsf{C}_{2}$.
-Up to the choice of the base $\Delta\subseteq\Phi$ any $5$-grading of $\mathsf{C}_{\ell}$ maps basic roots to 00 except one, and this distinguished root is mapped to $1$ (if it is short) or $2$ (if it is long).
+Up to the choice of the base $\Delta\subseteq\Phi$ any $5$-grading of $\mathsf{C}_{\ell}$ maps basic roots to 0 except one, and this distinguished root is mapped to $1$ (if it is short) or $2$ (if it is long).
 It easily follows that $\Sigma=\mathsf{C}_{m}\subseteq\mathsf{C}_{\ell}$ is a proper root subsystem generated by a connected Dynkin subdiagram containing the long basic root or it is empty.
-But the representation of $\mathrm{Sp}(2m,K)$ on $G_{\{\alpha\in\mathsf{C}_{m+1}\mid\alpha\cdot\mu>0\}}/G_{\mu}$ is the defining representation of symplectic group scheme, where $\mu\in\mathsf{C}_{m+1}$ is the highest root.
+But the representation of $\mathrm{Sp}(2m,K)$ on $G_{\{\alpha\in\mathsf{C}_{m+1}\mid\alpha\cdot\mu>0\}}/U_{\mu}$ is the defining representation of symplectic group scheme, where $\mu\in\mathsf{C}_{m+1}$ is the highest root.
 It follows that $C\leq T$ in all cases.
@@
 ∎
-If $\Phi$ is of type $\mathsf{BC}_{\ell}$ let $\mathrm{C}^{\mathrm{us}}(G)=\bigcap_{\alpha\in\mathsf{C}_{\ell}}\operatorname{C}_{L}(G_{\alpha})$ be the scheme centralizer of non-ultrashort root subgroups.
-This group scheme clearly contains the center.
+If $\Phi$ is of type $\mathsf{BC}_{\ell}$ with $\ell\geq 2$, let $\mathrm{C}^{\mathrm{us}}(G)=\bigcap_{\alpha\in\mathsf{C}_{\ell}}\operatorname{C}_{L}(U_{\alpha})$ be the scheme centralizer of non-ultrashort root subgroups.
+This group scheme clearly contains the center and it is contained in $L$ by theorem 1.
 It turns out that it is reductive or finite of multiplicative type, the second case holds only for $\mathsf{E}_{8,2}^{66}$.
-The derived subgroup of $\mathrm{C}^{\mathrm{us}}(G)$ corresponds to the odd component of $u^{-1}(0)$ in the classical cases and to the component $\mathsf{A}_{1}$ in the case $\mathsf{E}_{7,2}^{31}$.
-More precisely, assuming that $G$ splits and $\operatorname{C}(G)=1$ we have
+The derived subgroup of $\mathrm{C}^{\mathrm{us}}(G)$ corresponds to the “middle” component of $u^{-1}(0)$ in the classical cases and to the component $\mathsf{A}_{1}$ in the case $\mathsf{E}_{7,2}^{31}$.
+More precisely, assuming that $G$ is split and adjoint connected, we have the following.
 Tits index $\operatorname{C}^{\mathrm{us}}(G)$ conditions
-${{}^{2}\!{\mathsf{A}_{n,r}^{(d)}}}$ $\mathbb{GL}_{n+1-2rd}$ $d\mid n+1$, $n\geq 1$, $2rd\leq n$, $r\geq 2$
-$\mathsf{C}_{n,r}^{(d)}$ $\mathbb{S}\mathrm{p}_{2n-2rd}$ $2\leq d=2^{k}\mid 2n$, $rd<n$, $r\geq 2$
-${{}^{1}\!{\mathsf{D}_{n,r}^{(d)}}}$, ${{}^{2}\!{\mathsf{D}_{n,r}^{(d)}}}$ $\mathbb{SO}_{2n-2rd}$ $r\geq 2$, $2\leq d=2^{k}\mid 2n$, $rd<n$
-${{}^{2}\!{\mathsf{E}_{6,2}^{16^{\prime}}}}$ $\mathbb{G}_{\mathrm{m}}$ none
-$\mathsf{E}_{7,2}^{31}$ $\mathbb{SL}_{2}$ none
-$\mathsf{E}_{8,2}^{66}$ $\mu_{2}$ none
-In the classical cases the group scheme $\operatorname{C}^{\mathrm{us}}(G)$ can be computed using the block structure on $G$ as a matrix group.
-In the exceptional cases we apply lemma 1 to the splitting isotropic structure (the pinning) on $G$ and get $\operatorname{C}^{\mathrm{us}}(G)\leq T$ (for ${{}^{2}\!{\mathsf{E}_{6,2}^{16^{\prime}}}}$ and $\mathsf{E}_{8,2}^{66}$) or $\mathbb{SL}_{2}\leq\operatorname{C}^{\mathrm{us}}(G)\leq T\,\mathbb{SL}_{2}$.
-The intersection $T\cap\operatorname{C}^{\mathrm{us}}(G)$ is easy to compute using the root diagrams.
-For any root $\alpha\in\Phi$ let $\Gamma_{\alpha}=\{\beta\in\Phi\mid\alpha+\beta\notin\Phi\cup\{0\}\}$ and $Z_{\alpha}=\bigcup_{\beta\in\Gamma_{\alpha}}t_{\beta}(X_{\beta})\subseteq G(K)$ for some generating sets $X_{\beta}\subseteq\mathfrak{g}_{\beta}$ if $\beta$ is not ultrashort and generating sets $(X_{\beta},X_{2\beta})\subseteq(\mathfrak{g}_{\beta}\mathbin{\dot{\oplus}}\mathfrak{g}_{2\beta},\mathfrak{g}_{2\beta})$ for ultrashort $\beta$.
-It is easy to see by considering all rank $2$ and rank $3$ irreducible root systems that $\Gamma_{\alpha}$ is closed.
+$^{2}\!{\mathsf{A}_{n,r}^{(d)}}$} $\mathbb{GL}_{n+1-2rd}$ $d\mid n+1$, $2rd\leq n$, $d\geq 1$, $r\geq 2$
+$\mathsf{C}_{n,r}^{(d)}$ $\mathbb{S}\mathrm{p}_{2n-2rd}$ $2\leq d=2^{k}\mid 2n$, $rd\leq n-1$, $r\geq 2$
+$^{1}\!{\mathsf{D}_{n,r}^{(d)}}$}, $^{2}\!{\mathsf{D}_{n,r}^{(d)}}$} $\mathbb{SO}_{2n-2rd}$ $r\geq 2$, $2\leq d=2^{k}\mid 2n$, $rd\leq n-1$
+$^{2}\!{\mathsf{E}_{6,2}^{16^{\prime}}}$} $\mathbb{G}_{\mathrm{m}}$
+$\mathsf{E}_{7,2}^{31}$ $\mathbb{SL}_{2}$
+$\mathsf{E}_{8,2}^{66}$ $\mu_{2}$
+In the classical cases the group scheme $\operatorname{C}^{\mathrm{us}}(G)$ can be computed using the block structure on a finite central extension of $G$ as a matrix group.
+In the exceptional cases we apply theorem 1 to the splitting isotropic structure on $G$ and get $\operatorname{C}^{\mathrm{us}}(G)\leq T$ (for $^{2}\!{\mathsf{E}_{6,2}^{16^{\prime}}}$} and $\mathsf{E}_{8,2}^{66}$) or $\mathbb{SL}_{2}\leq\operatorname{C}^{\mathrm{us}}(G)\leq T\,\mathbb{SL}_{2}$, where $T$ is the maximal torus from the splitting isotropic structure.
+The intersection $T\cap\operatorname{C}^{\mathrm{us}}(G)$ is easy to compute using the root diagrams [22].
+For any root $\alpha\in\Phi$ let $\Gamma_{\alpha}=\{\beta\in\Phi\mid\alpha+\beta\notin\Phi\cup\{0\}\}$ and $Z_{\alpha}=\bigcup_{\beta\in\Gamma_{\alpha}}t_{\beta}(X_{\beta})\subseteq G(K)$ for some generating sets $X_{\beta}\subseteq\mathfrak{g}_{\beta}$ for non-ultrashort $\beta$ and generating sets $(X_{\beta},X_{2\beta})\subseteq(\mathfrak{g}_{\beta}\mathbin{\dot{\oplus}}\mathfrak{g}_{2\beta},\mathfrak{g}_{2\beta})$ for ultrashort $\beta$.
+It is easy to see by considering all rank $2$ and rank $3$ irreducible root systems that $\Gamma_{\alpha}$ is closed and $\angle(\alpha,\beta)\leq\frac{\pi}{2}$ for all $\beta\in\Gamma_{\alpha}$.
 Theorem 2.
 Let $K$ be a commutative ring, $G$ an isotropic reductive group scheme over $K$ with root system $\Phi$ of rank at least $2$, and $\alpha\in\Phi$ a root.
-  • If the type of $\Phi$ is $\mathsf{BC}_{\ell}$ and $\alpha$ is ultrashort, then $\operatorname{C}_{G}(Z_{\alpha})=\operatorname{C}^{\mathrm{us}}(G)\,G_{\alpha}$.
-  • If $\alpha=\mathrm{e}_{i}+\mathrm{e}_{j}$ (using the convention $\mathrm{e}_{-i}=-\mathrm{e}_{i}$) is short and the type of $\Phi$ is $\mathsf{C}_{\ell}$ or $\mathsf{BC}_{\ell}$, including $\mathsf{B}_{2}=\mathsf{C}_{2}$, then $\operatorname{C}_{G}(Z_{\alpha})=\operatorname{C}(G)\,G_{2\mathrm{e}_{i}}\,G_{\mathrm{e}_{i}+\mathrm{e}_{j}}\,G_{2\mathrm{e}_{j}}$.
-  • Otherwise $\operatorname{C}_{G}(Z_{\alpha})=\operatorname{C}(G)\,G_{\alpha}$.
+  • If the type of $\Phi$ is $\mathsf{BC}_{\ell}$ and $\alpha$ is ultrashort, then $\operatorname{C}_{G}(Z_{\alpha})=\operatorname{C}^{\mathrm{us}}(G)\,U_{\alpha}$.
+  • If $\alpha=\mathrm{e}_{i}+\mathrm{e}_{j}$ (using the convention $\mathrm{e}_{-i}=-\mathrm{e}_{i}$) is short and the type of $\Phi$ is $\mathsf{C}_{\ell}$ or $\mathsf{BC}_{\ell}$, including $\mathsf{B}_{2}=\mathsf{C}_{2}$, then $\operatorname{C}_{G}(Z_{\alpha})=\operatorname{C}(G)\,U_{2\mathrm{e}_{i}}\,U_{\mathrm{e}_{i}+\mathrm{e}_{j}}\,U_{2\mathrm{e}_{j}}$.
+  • Otherwise $\operatorname{C}_{G}(Z_{\alpha})=\operatorname{C}(G)\,U_{\alpha}$.
 Proof.
@@
 $\Sigma=\{\gamma\in\Phi\mid\angle(\beta,\gamma)\leq\frac{\pi}{2}\text{ for all long }\beta\in\Gamma_{\alpha}\}.$
-Considering the saturated root subsystem of rank $2$ (or rank $3$ in the doubly laced cases) containing $\alpha$ and arbitrary $\beta\in\Sigma\setminus\mathbb{R}\alpha$ we see that
+Considering saturated root subsystems of rank $2$ and $3$ containing $\alpha$ and arbitrary $\beta\in\Sigma$ (or just explicitly calculating $\Gamma_{\alpha}$ in all cases) we see that
   • If $\Phi$ is of type $\mathsf{C}_{\ell}$ with $\ell\geq 2$ and $\alpha=\mathrm{e}_{i}+\mathrm{e}_{j}$ is short (recall that $\mathrm{e}_{-i}=-\mathrm{e}_{i}$), then $\Sigma=\{2\mathrm{e}_{i},\mathrm{e}_{i}+\mathrm{e}_{j},2\mathrm{e}_{j}\}$.
@@
 Now apply lemma 4 to get
-$\operatorname{C}_{G}(Z_{\alpha})\leq\begin{cases}G^{0}_{\{\alpha,2\alpha\}}&\text{ if $\Phi$ is of type $\mathsf{BC}_{\ell}$, $\ell\geq 2$, $\alpha$ is ultrashort};\\ G^{0}_{\{2\mathrm{e}_{i},\mathrm{e}_{i}+\mathrm{e}_{j},2\mathrm{e}_{j}\}}&\text{ if $\Phi$ is of type $\mathsf{C}_{\ell}$ or $\mathsf{BC}_{\ell}$, $\ell\geq 2$, $\alpha=\mathrm{e}_{i}+\mathrm{e}_{j}$ is short};\\ G^{0}_{\alpha}&\text{ otherwise}.\end{cases}$
-It remains to show that $\operatorname{C}_{G}(Z_{\alpha})\cap L=\mathrm{C}(G)$ or $\mathrm{C}^{\mathrm{us}}(G)$.
-If $\alpha$ is not ultrashort we are done by lemma 5 applied to $f(\beta)=2\frac{\alpha\cdot\beta}{\alpha\cdot\alpha}$.
-Otherwise note that $\Gamma_{\alpha}$ is a parabolic subset of $\mathsf{C}_{\ell}\subseteq\mathsf{BC}_{\ell}$.
-Applying lemma 5 to $G^{0}_{\{-\beta,\beta\}}$ for various $\beta\in\Gamma_{\alpha}$ with $f(\gamma)=2\frac{\gamma\cdot\beta}{\beta\cdot\beta}$ (more precisely, to its simple subgroup containing $G_{\beta}$ after a splitting étale extension) we see that $\operatorname{C}_{L}(t_{\beta}(X_{\beta}))$ centralizes both $G_{\beta}$ and $G_{-\beta}$, so $\operatorname{C}_{G}(Z_{\alpha})\cap L=\operatorname{C}^{\mathrm{us}}(G)$.
+[…]
+It remains to show that $\operatorname{C}_{L}(Z_{\alpha})=\mathrm{C}(G)$ or $\mathrm{C}^{\mathrm{us}}(G)$.
+If $\alpha$ is neither ultrashort nor short in $\Phi=\mathsf{G}_{2}$, we are done by lemma 5 applied to $f(\beta)=2\frac{\alpha\cdot\beta}{\alpha\cdot\alpha}$.
+In the case of $\mathsf{G}_{2}$ first apply the same argument to $G_{\{-\alpha,\alpha\}}$ to get $\operatorname{C}_{L}(Z_{\alpha})\leq\operatorname{C}(L)$ (using $L,L]\leq G_{\{-\alpha,\alpha\}}$), and then note that $\operatorname{C}_{\operatorname{C}(L)}(Z_{\alpha})=\operatorname{C}(G)$ because the preimage of $\Gamma_{\alpha}$ in $\widetilde{\Phi}$ spans the whole root lattice e.g.
+by considering root diagrams [22].
+It remains to consider $\Phi=\mathsf{BC}_{\ell}$ and ultrashort $\alpha$.
+Note that $\Gamma_{\alpha}$ is a parabolic subset of $\mathsf{C}_{\ell}\subseteq\mathsf{BC}_{\ell}$.
+Applying lemma 5 to $G_{\mathsf{C}_{\ell}}$ with $f(\beta)=2\frac{\beta\cdot\alpha}{\alpha\cdot\alpha}$ we see that $\operatorname{C}_{L}(Z_{\alpha})$ centralizes all $U_{\beta}$ with $\beta\in\mathsf{C}_{\ell}$ (note that $G_{\mathsf{C}_{\ell}}\leq G^{0}_{\mathsf{C}_{\ell}}$ is a simple factor e.g.
+by [26, lemma 1] and all other factors centralize $G_{\mathsf{C}_{\ell}}\supset Z_{\alpha}$), so $\operatorname{C}_{L}(Z_{\alpha})=\operatorname{C}^{\mathrm{us}}(G)$.
 ∎
@@
 Then
-$\operatorname{C}_{G}\bigl{(}\bigcup_{\alpha\in\Phi}t_{\alpha}(X_{\alpha})\bigr{)}=\operatorname{C}(G),\quad\{g\in G\mid\forall\alpha\in\Phi\enskip{{}^{g}\!{t_{\alpha}(X_{\alpha})}}\subseteq G_{\alpha}\}=L.$
+$\operatorname{C}_{G}\bigl(\bigcup_{\alpha\in\Phi}t_{\alpha}(X_{\alpha})\bigr)=\operatorname{C}(G),\quad\{g\in G\mid\forall\alpha\in\Phi\enskip{{}^{g}\!{t_{\alpha}(X_{\alpha})}}\subseteq U_{\alpha}\}=L.$
 Proof.
-The first claim follows from lemma 2 applied to opposite long roots and lemma 2. The second one is a corollary of theorem 1 applied to all long roots and lemma 2. ∎
+The first claim follows from theorem 2 applied to opposite long roots and lemma 2. The second one is a corollary of theorem 1 applied to all long roots and lemma 2. ∎
 5 Main results
+Recall that $\mathrm{E}_{G}(K)=\bigl\langle U_{\alpha}(K)\mid\alpha\in\Phi\bigr\rangle$ is the elementary subgroup [20, 26].
+If $\mathrm{E}_{G}(K)\leq H\leq G(K)$ is an intermediate subgroup, then $\operatorname{C}(G(K))\cap H=\operatorname{C}(H)=\operatorname{C}_{H}(\mathrm{E}_{G}(K))$ by theorem 3.
 Theorem 4.
 Let $G$ be an isotropic reductive group scheme over $K$ such that the rank of $\Phi$ is at least $2$.
-Then the sets $L(K)$, $\operatorname{C}(G)(K)$, and all $\operatorname{C}(G)(K)\,G_{\alpha}(K)$ are Diophantine in the group $G(K)$.
-Moreover, the root subgroups $G_{\alpha}(K)$ are Diophantine in the group $G(K)$ unless the type is $\mathsf{B}_{2}=\mathsf{C}_{2}$, $\mathsf{BC}_{2}$, or $\alpha$ is short and the type is $\mathsf{G}_{2}$.
+Then the sets $L(K)$, $\operatorname{C}(G)(K)$, and all $\operatorname{C}(G)(K)\,U_{\alpha}(K)$ are Diophantine in the group $G(K)$.
+Moreover, the root subgroups $U_{\alpha}(K)$ are Diophantine in the group $G(K)$ unless the type is $\mathsf{B}_{2}=\mathsf{C}_{2}$, $\mathsf{BC}_{2}$, or $\alpha$ is short and the type is $\mathsf{G}_{2}$.
+If $\mathrm{E}_{G}(K)\leq H\leq G(K)$ is an intermediate subgroup, then the intersections of all above subgroups with $H$ are Diophantine in $H$.
 Proof.
-By theorem 2 the following groups are Diophantine.
-  • $\operatorname{C}^{\mathrm{us}}(G)(K)\,G_{\alpha}(K)$ for ultrashort $\alpha$ and $\Phi$ of type $\mathsf{BC}_{\ell}$;
-  • $\operatorname{C}(G)(K)\,G_{2\mathrm{e}_{i}}(K)\,G_{\mathrm{e}_{i}+\mathrm{e}_{j}}(K)\,G_{2\mathrm{e}_{j}}(K)$ for $\Phi$ of type $\mathsf{C}_{\ell}$ or $\mathsf{BC}_{\ell}$, $\ell\geq 2$;
-  • $\mathrm{C}(G)(K)\,G_{\alpha}(K)$ for $\alpha$ long or $\alpha$ short and $\Phi$ of type $\mathsf{B}_{\ell}$, $\ell\geq 3$, $\mathsf{F}_{4}$, or $\mathsf{G}_{2}$.
-Intersections of Diophantine groups, their products and their commutators with single elements are also Diophantine.
+By theorem 2 the following subgroups are Diophantine.
+  • $\operatorname{C}^{\mathrm{us}}(G)(K)\,U_{\alpha}(K)$ for ultrashort $\alpha$ and $\Phi$ of type $\mathsf{BC}_{\ell}$;
+  • $\operatorname{C}(G)(K)\,U_{2\mathrm{e}_{i}}(K)\,U_{\mathrm{e}_{i}+\mathrm{e}_{j}}(K)\,U_{2\mathrm{e}_{j}}(K)$ for $\alpha=\mathrm{e}_{i}+\mathrm{e}_{j}$ and $\Phi$ of type $\mathsf{C}_{\ell}$ or $\mathsf{BC}_{\ell}$, $\ell\geq 2$;
+  • $\mathrm{C}(G)(K)\,U_{\alpha}(K)$ for $\alpha$ long or $\alpha$ short and $\Phi$ of type $\mathsf{B}_{\ell}$, $\ell\geq 3$, $\mathsf{F}_{4}$, or $\mathsf{G}_{2}$.
+Intersections of Diophantine subgroups, their products and their commutators with single elements are also Diophantine.
 Clearly,
-$\operatorname{C}(G)(K)=\operatorname{C}(G)(K)\,G_{\alpha}(K)\cap\operatorname{C}(G)(K)\,G_{-\alpha}(K)$
+$\operatorname{C}(G)(K)=\operatorname{C}(G)(K)\,U_{\alpha}(K)\cap\operatorname{C}(G)(K)\,U_{-\alpha}(K)$
 is Diophantine, where $\alpha$ is any long root. By theorem 1 applied to $G/\operatorname{C}(G)$ we have
-$\{g\in G(K)\mid{{}^{g}\!{t_{\alpha}(X)}}\subseteq\operatorname{C}(G)(K)\,G_{\alpha}(K)\}=G^{0}_{\{\beta\mid\angle(\alpha,\beta)\leq\pi/2\}}(K)$
+$\{g\in G(K)\mid{{}^{g}\!{t_{\alpha}(X)}}\subseteq\operatorname{C}(G)(K)\,U_{\alpha}(K)\}=G^{0}_{\{\beta\mid\angle(\alpha,\beta)\leq\pi/2\}}(K)$
 for every long root $\alpha$, where $X\subseteq\mathfrak{g}_{\alpha}$ is a finite generating set. The intersection of all these parabolic subgroups is $L(K)$ by lemma 2.
 If $\alpha$ is long and the type of $\Phi$ is neither $\mathsf{C}_{\ell}$ nor $\mathsf{BC}_{\ell}$, then
-$G_{\alpha}(K)=\bigl{[}\operatorname{C}(G)(K)\,G_{\beta}(K),t_{\alpha-\beta}(e)\bigr{]}$
-is Diophantine, where $\beta$ is long, $\angle(\alpha,\beta)=\frac{\pi}{3}$, and $t_{\alpha-\beta}(e)\in G_{\alpha-\beta}$ is a factor of a Weyl element.
+$U_{\alpha}(K)=\prod_{x\in X}\bigl[\operatorname{C}(G)(K)\,U_{\beta}(K),t_{\alpha-\beta}(x)\bigr]$
+is Diophantine, where $\beta$ is long, $\angle(\alpha,\beta)=\frac{\pi}{3}$, and $X\subseteq\mathfrak{g}_{\alpha-\beta}$ is a finite generating set.
 If $\alpha=\mathrm{e}_{i}$ is short and the type of $\Phi$ is $\mathsf{B}_{\ell}$, $\ell\geq 3$, then
-$G_{\alpha}(K)=\operatorname{C}(G)(K)\,G_{\alpha}(K)\cap\bigl{[}\operatorname{C}(G)(K)\,G_{\mathrm{e}_{j}}(K),t_{\mathrm{e}_{i}-\mathrm{e}_{j}}(e)\bigr{]}\,G_{\mathrm{e}_{i}+\mathrm{e}_{j}}(K)$
-is Diophantine, where $t_{\mathrm{e}_{i}-\mathrm{e}_{j}}(e)$ is a factor of a Weyl element. It follows that the same holds for the type $\mathsf{F}_{4}$.
+$U_{\alpha}(K)=\operatorname{C}(G)(K)\,U_{\alpha}(K)\cap G_{\mathrm{e}_{i}+\mathrm{e}_{j}}(K)\,\prod_{x\in X}\bigl[\operatorname{C}(G)(K)\,G_{\mathrm{e}_{j}}(K),t_{\mathrm{e}_{i}-\mathrm{e}_{j}}(x)\bigr]$
+is Diophantine, where $j\neq\pm i$ is another index and $X\subseteq\mathfrak{g}_{\mathrm{e}_{i}-\mathrm{e}_{j}}$ is a finite generating set.
+It follows that all root subgroups are Diophantine for the type $\mathsf{F}_{4}$.
 Now suppose that $\Phi$ is of type $\mathsf{C}_{\ell}$ or $\mathsf{BC}_{\ell}$ and $\ell\geq 3$. We have
-$G_{\mathrm{e}_{i}+\mathrm{e}_{j}}(K)=\bigl{[}\bigl{[}\operatorname{C}(G)(K)\,G_{2\mathrm{e}_{j}}(K)\,G_{\mathrm{e}_{j}+\mathrm{e}_{k}}(K)\,G_{2\mathrm{e}_{k}}(K),t_{\mathrm{e}_{i}-\mathrm{e}_{j}}(e)\bigr{]},t_{\mathrm{e}_{j}-\mathrm{e}_{k}}(e^{\prime})\bigr{]},$
-where constants are factors of Weyl elements. Next, the group
-$G_{2\mathrm{e}_{i}}(K)=\operatorname{C}(G)\,G_{2\mathrm{e}_{i}}(K)\cap\bigl{[}\operatorname{C}(G)\,G_{2\mathrm{e}_{j}}(K),t_{\mathrm{e}_{i}-\mathrm{e}_{j}}(e)]\,G_{\mathrm{e}_{i}+\mathrm{e}_{j}}(K)$
-is Diophantine. If $\Phi$ is of type $\mathsf{BC}_{\ell}$, then
-$G_{\mathrm{e}_{i}}(K)=\operatorname{C}^{\mathrm{us}}(G)(K)\,G_{\mathrm{e}_{i}}(K)\cap\bigl{[}\operatorname{C}^{\mathrm{us}}(G)(K)\,G_{\mathrm{e}_{j}}(K),t_{\mathrm{e}_{i}-\mathrm{e}_{j}}(e)\bigr{]}\,G_{\mathrm{e}_{i}+\mathrm{e}_{j}}(K)\,G_{\mathrm{e}_{i}-\mathrm{e}_{j}}(K)$
-is also Diophantine.
+$U_{\mathrm{e}_{i}+\mathrm{e}_{j}}(K)=\prod_{x\in X,\,x^{\prime}\in X^{\prime}}\bigl[\bigl[\operatorname{C}(G)(K)\,U_{2\mathrm{e}_{j}}(K)\,U_{\mathrm{e}_{j}+\mathrm{e}_{k}}(K)\,U_{2\mathrm{e}_{k}}(K),t_{\mathrm{e}_{i}-\mathrm{e}_{j}}(x)\bigr],t_{\mathrm{e}_{j}-\mathrm{e}_{k}}(x^{\prime})\bigr],$
+where $k\notin\{-i,i,-j,j\}$ is a new index and $X\subseteq\mathfrak{g}_{\mathrm{e}_{i}-\mathrm{e}_{j}}$, $X^{\prime}\subseteq\mathfrak{g}_{\mathrm{e}_{j}-\mathrm{e}_{k}}$ are finite generating sets.
+Next, the group
+$U_{2\mathrm{e}_{i}}(K)=\operatorname{C}(G)\,U_{2\mathrm{e}_{i}}(K)\cap U_{\mathrm{e}_{i}+\mathrm{e}_{j}}(K)\,\prod_{x\in X}\bigl[\operatorname{C}(G)\,U_{2\mathrm{e}_{j}}(K),t_{\mathrm{e}_{i}-\mathrm{e}_{j}}(x)]$
+is Diophantine, where $X\subseteq\mathfrak{g}_{\mathrm{e}_{i}-\mathrm{e}_{j}}$ is a finite generating set. If $\Phi$ is of type $\mathsf{BC}_{\ell}$, then
+$U_{\mathrm{e}_{i}}(K)=\operatorname{C}^{\mathrm{us}}(G)(K)\,U_{\mathrm{e}_{i}}(K)\cap U_{\mathrm{e}_{i}-\mathrm{e}_{j}}(K)\,U_{\mathrm{e}_{i}+\mathrm{e}_{j}}(K)\,\prod_{x\in X}\bigl[\operatorname{C}^{\mathrm{us}}(G)(K)\,U_{\mathrm{e}_{j}}(K),t_{\mathrm{e}_{i}-\mathrm{e}_{j}}(x)\bigr]$
+is also Diophantine, where $X\subseteq\mathfrak{g}_{\mathrm{e}_{i}-\mathrm{e}_{j}}$ is a finite generating set.
 Finally, consider the exceptional cases $\mathsf{C}_{2}$ and $\mathsf{BC}_{2}$.
-The group $\operatorname{C}(G)(K)\,G_{\mathrm{e}_{i}}(K)$ is Diophantine (in the case of $\mathsf{BC}_{2}$) since it contains $g\in\operatorname{C}^{\mathrm{us}}(K)\,G_{\mathrm{e}_{i}}(K)$ if and only if
-$\bigl{[}g,t_{\mathrm{e}_{i}}(X_{0})\bigr{]}\subseteq\operatorname{C}(G)(K)\,G_{2\mathrm{e}_{i}}(K)$
-for some generating set $X_{0}\subseteq\mathfrak{g}_{\mathrm{e}_{i}}\mathbin{\dot{\oplus}}\mathfrak{g}_{2\mathrm{e}_{i}}$. We have
-$\operatorname{C}(G)\,G_{2\mathrm{e}_{i}}(K)\,G_{\mathrm{e}_{i}+\mathrm{e}_{j}}(K)=\operatorname{C}(G)\,G_{2\mathrm{e}_{i}}(K)\,\prod_{x\in X}\bigl{[}\operatorname{C}(G)\,G_{2\mathrm{e}_{i}}(K)\,G_{\mathrm{e}_{i}+\mathrm{e}_{j}}(K)\,G_{2\mathrm{e}_{j}}(K),t_{\mathrm{e}_{i}-\mathrm{e}_{j}}(x)\bigr{]}$
-for a finite generating set $X\subseteq\mathfrak{g}_{\mathrm{e}_{i}-\mathrm{e}_{j}}$ by lemma 3, so $\operatorname{C}(G)\,G_{\mathrm{e}_{i}+\mathrm{e}_{j}}(K)$ is also Diophantine. ∎
+The group $\operatorname{C}(G)(K)\,U_{\mathrm{e}_{i}}(K)$ is Diophantine (in the case of $\mathsf{BC}_{2}$) because it contains $g\in\operatorname{C}^{\mathrm{us}}(K)\,U_{\mathrm{e}_{i}}(K)$ if and only if
+$\bigl[g,t_{\mathrm{e}_{i}}(x)\bigr]\in\operatorname{C}(G)(K)\,U_{2\mathrm{e}_{i}}(K)$
+for all $x\in X$, where $X\subseteq\mathfrak{g}_{\mathrm{e}_{i}}\mathbin{\dot{\oplus}}\mathfrak{g}_{2\mathrm{e}_{i}}$ is a finite generating set. We have
+$\operatorname{C}(G)\,U_{2\mathrm{e}_{i}}(K)\,U_{\mathrm{e}_{i}+\mathrm{e}_{j}}(K)=\operatorname{C}(G)\,U_{2\mathrm{e}_{i}}(K)\,\prod_{x\in X}\bigl[\operatorname{C}(G)\,U_{2\mathrm{e}_{i}}(K)\,U_{\mathrm{e}_{i}+\mathrm{e}_{j}}(K)\,U_{2\mathrm{e}_{j}}(K),t_{\mathrm{e}_{i}-\mathrm{e}_{j}}(x)\bigr]$
+for a finite generating set $X\subseteq\mathfrak{g}_{\mathrm{e}_{i}-\mathrm{e}_{j}}$ by lemma 3, so $\operatorname{C}(G)\,U_{\mathrm{e}_{i}+\mathrm{e}_{j}}(K)$ is also Diophantine.
+The last claim about intersections with $\mathrm{E}_{G}(K)\leq H\leq G(K)$ easily follows from explicit Diophantine formulas for $L(K)$, $\operatorname{C}(G)(K)$, $\operatorname{C}(G)(K)\,U_{\alpha}(K)$, and $U_{\alpha}(K)$.
+∎
 Since $G$ is a finitely presented affine scheme, the point group $G(K)$ is e-interpretable in $K$. The following theorem shows the converse.
 Theorem 5.
-Let $G$ be an isotropic reductive group scheme over $K$ with the root system $\Phi$ of rank at least $2$. Then the ring $K$ is e-interpretable in $G(K)$.
+Let $G$ be an isotropic reductive group scheme over $K$ with the root system $\Phi$ of rank at least $2$.
+Then the ring $K$ is e-interpretable in any intermediate subgroup $\mathrm{E}_{G}(K)\leq H\leq G(K)$.
 Proof.
-We assume that the type of $\Phi$ is neither $\mathsf{G}_{2}$ nor $\mathsf{BC}_{\ell}$ since these cases easily reduce to $\mathsf{A}_{2}$ and $\mathsf{C}_{\ell}$ respectively.
+We assume that the type of $\Phi$ is neither $\mathsf{G}_{2}$ nor $\mathsf{BC}_{\ell}$ because these cases easily reduce to $\mathsf{A}_{2}$ and $\mathsf{C}_{\ell}$ respectively.
 Fix some finite $K$-module generating sets $X_{\alpha}\subseteq\mathfrak{g}_{\alpha}$ for all roots $\alpha$ and let $Y_{\alpha}=X_{\alpha}\cup\bigcup_{\beta,\alpha-\beta\in\Phi}[X_{\beta},X_{\alpha-\beta}]$.
+Below we use that up to canonical isomorphisms the groups $\mathfrak{g}_{\alpha}$ are e-interpretable in $H$ by theorem 4 as Diophantine subgroups $U_{\alpha}(K)\leq H$ or as factor-groups $\operatorname{C}(H)\,U_{\alpha}(K)/\operatorname{C}(H)$.
+Under these e-interpretations all commutator maps $\mathfrak{g}_{\alpha}\times\mathfrak{g}_{\beta}\to\mathfrak{g}_{\alpha+\beta}$ with $\alpha,\beta,\alpha+\beta\in\Phi$ have Diophantine graphs.
 Let $\widetilde{K}$ be the set of families
-$(k_{\alpha}\colon Y_{\alpha}\to\operatorname{C}(G)(K)\,G_{\alpha}(K)/\operatorname{C}(G)(K))_{\alpha\in\Phi}$
+$(k_{\alpha}\colon Y_{\alpha}\to\mathfrak{g}_{\alpha}(K))_{\alpha\in\Phi}$
 such that
-$[k_{\alpha}(y_{\alpha}),t_{\beta}(y_{\beta})]\equiv[t_{\alpha}(y_{\alpha}),k_{\beta}(y_{\beta})]\pmod{\operatorname{C}(G)(K)\,\prod_{\begin{subarray}{c}i\alpha+j\beta\in\Phi\\ i,j\geq 1\\ (i,j)\neq(1,1)\end{subarray}}G_{i\alpha+j\beta}}$
+$k_{\alpha}(y_{\alpha}),y_{\beta}]=[y_{\alpha},k_{\beta}(y_{\beta})]\in\mathfrak{g}_{\alpha+\beta}$
 for $y_{\alpha}\in Y_{\alpha}$, $y_{\beta}\in Y_{\beta}$ and
-$[k_{\alpha}(x_{\alpha}),t_{\beta}(x_{\beta})]\equiv k_{\alpha+\beta}([x_{\alpha},x_{\beta}])\pmod{\operatorname{C}(G)(K)\,\prod_{\begin{subarray}{c}i\alpha+j\beta\in\Phi\\ i,j\geq 1\\ (i,j)\neq(1,1)\end{subarray}}G_{i\alpha+j\beta}}$
-for $x_{\alpha}\in X_{\alpha}$, $x_{\beta}\in X_{\beta}$.
-Here $\alpha$ and $\beta$ are all roots such that $\alpha+\beta$ is also a root.
-Clearly, $\widetilde{K}$ is the factor-set of a Diophantine set by a Diophantine equivalence relation.
-Every $k\in K$ determines a corresponding element $\widetilde{k}\in\widetilde{K}$, $\widetilde{k}(y_{\alpha})=t_{\alpha}(ky_{\alpha})$.
-Conversely, consider elements $u_{y_{\alpha}}\in\mathfrak{g}_{\alpha}$ for $y_{\alpha}\in Y_{\alpha}$ such that $k_{\alpha}(y_{\alpha})\in t_{\alpha}(y_{x_{\alpha}})\,\operatorname{C}(G)$.
-For any root $\alpha$ there is long $\beta\in\Phi$ such that $\alpha+\beta\in\Phi$, so $[u_{y_{\alpha}},y_{\beta}]=[y_{\alpha},u_{y_{\beta}}]$.
-Recall that the Lie bracket $[{-},{=}]\colon\mathfrak{g}_{\alpha}\times\mathfrak{g}_{\beta}\to\mathfrak{g}_{\alpha+\beta}$ is non-degenerate (by lemma 3 if $\alpha$ is short).
-For any linear relation $\sum_{y_{\alpha}\in Y_{\alpha}}y_{\alpha}a_{\alpha}=0$ the equation implies that $\sum_{y_{\alpha}\in Y_{\alpha}}u_{y_{\alpha}}a_{\alpha}=0$, so there is a unique linear map $U_{\alpha}\colon\mathfrak{g}_{\alpha}\to\mathfrak{g}_{\alpha}$ such that $u_{y_{\alpha}}=U_{\alpha}(y_{\alpha})$.
-The defining congruences imply that $[U_{\alpha}(x),y]=U_{\alpha+\beta}([x,y])=[x,U_{\beta}(y)]$.
-We claim that such maps $U_{\alpha}$ are necessarily scalar.
-It suffices to consider $\Phi$ of rank $2$ and to check the claim for only one root $\alpha$.
-This is clear if one of $\mathfrak{g}_{\alpha}$ is one-dimensional.
-  • In the linear-like case (${{}^{1}\!{\mathsf{E}_{6,2}^{28}}}$ and ${{}^{1}\!{\mathsf{A}_{3d-1,2}^{(d)}}}$) we have
-    $U_{\mathrm{e}_{1}-\mathrm{e}_{2}}(x)\,y=U_{\mathrm{e}_{1}-\mathrm{e}_{3}}(xy)=x\,U_{\mathrm{e}_{2}-\mathrm{e}_{3}}(y)$
-    for $x,y\in A$, where $A$ is a composition algebra of rank $8$ or an Azumaya algebra over $K$.
-Hence $U_{\mathrm{e}_{1}-\mathrm{e}_{2}}(x)=U_{\mathrm{e}_{1}-\mathrm{e}_{3}}(x)$ (i.e.
-all maps $U_{\alpha}$ coincide) and $U_{\alpha}(x)=x\,U_{\alpha}(1)=U_{\alpha}(1)\,x$, so $U_{\alpha}(1)$ lies in the center of $A$, and this center is precisely $K$.
-  • In the orthogonal-like case ($\mathsf{B}_{n,2}$, ${{}^{1}\!{\mathsf{D}_{n,2}^{(d)}}}$, ${{}^{2}\!{\mathsf{D}_{n,2}^{(d)}}}$) we have
-    $U_{\mathrm{e}_{2}}(m)\,r=m\,U_{\mathrm{e}_{1}-\mathrm{e}_{2}}(r)=U_{\mathrm{e}_{1}}(mr)$
-    for $r\in R$ and $m\in M$, where $R$ is an Azumaya algebra over $K$ and $M_{R}$ is a faithfully projective module.
-We have $U_{\mathrm{e}_{1}}(m)=U_{\mathrm{e}_{2}}(m)=m\,U_{\mathrm{e}_{1}-\mathrm{e}_{2}}(1)$ and $U_{\mathrm{e}_{1}-\mathrm{e}_{2}}(r)=r\,U_{\mathrm{e}_{1}-\mathrm{e}_{2}}(1)=U_{\mathrm{e}_{1}-\mathrm{e}_{2}}(1)\,r$, so $U_{\mathrm{e}_{1}-\mathrm{e}_{2}}(1)\in K=\operatorname{C}(R)$.
-  • In the remaining symplectic-like case we have
-    $U_{\mathrm{e}_{1}-\mathrm{e}_{2}}(r)\,u=r\,U_{2\mathrm{e}_{2}}(u)=U_{\mathrm{e}_{1}+\mathrm{e}_{2}}(ru)$
-    for $r\in R$ and $u\in\Lambda$, where $R$ is an Azumaya algebra over $K$ or its quadratic étale extension and $\Lambda\leq R$ is a form parameter containing invertible element $\iota$.
-It follows that $U_{2\mathrm{e}_{2}}(u)=U_{\mathrm{e}_{1}-\mathrm{e}_{2}}(1)\,u$, $U_{\mathrm{e}_{1}-\mathrm{e}_{2}}(r)=r\,U_{\mathrm{e}_{1}-\mathrm{e}_{2}}(1)$, and $U_{\mathrm{e}_{1}+\mathrm{e}_{2}}(r)=r\iota^{-1}\,U_{\mathrm{e}_{1}-\mathrm{e}_{2}}(1)\,\iota$.
-Then $U_{\mathrm{e}_{1}-\mathrm{e}_{2}}(1)\,u\iota^{-1}=u\iota^{-1}\,U_{\mathrm{e}_{1}-\mathrm{e}_{2}}(1)$ and $U_{\mathrm{e}_{1}-\mathrm{e}_{2}}(1)\in\Lambda\iota^{-1}$.
-In the subcase ${{}^{2}\!{\mathsf{A}_{n,2}^{(d)}}}$ the group $\Lambda=R_{0}$ is an Azumaya algebra over $K$ (so $R=R_{0}\otimes\operatorname{C}(R)$) and $U_{\mathrm{e}_{1}-\mathrm{e}_{2}}(1)\in\Lambda\iota^{-1}\cap\operatorname{C}(R)=K$.
-In the subcase $\mathsf{C}_{n,2}^{(d)}$ the form parameter is
-    $\Lambda=\{x\in\mathrm{M}(d,K)\mid x_{ij}=x_{d+1-j,d+1-i}\}$
-    and $\iota=1$tale locally, so $U_{\mathrm{e}_{1}-\mathrm{e}_{2}}(1)\in\operatorname{C}_{R}(\Lambda)=K$. In the last subcase ${{}^{1}\!{\mathsf{D}_{n,2}^{(d)}}}$ we have
-    $\Lambda=\{x\in\mathrm{M}(d,K)\mid x_{ij}=-x_{d+1-j,d+1-i},x_{i,d+1-i}=0\}$
-    and $\iota=e_{11}+\ldots+e_{\frac{d}{2},\frac{d}{2}}-e_{\frac{d}{2}+1,\frac{d}{2}+1}-\ldots-e_{dd}$tale locally, so again $U_{\mathrm{e}_{1}-\mathrm{e}_{2}}(1)\in\operatorname{C}_{R}(\Lambda\iota^{-1})\cap\Lambda\iota^{-1}=K$ (if $d\geq 2$, then already $\operatorname{C}_{R}(\Lambda\iota^{-1})=K$).
+$k_{\alpha}(x_{\alpha}),x_{\beta}]=k_{\alpha+\beta}([x_{\alpha},x_{\beta}])\in\mathfrak{g}_{\alpha+\beta}$
+for $x_{\alpha}\in X_{\alpha}$, $x_{\beta}\in X_{\beta}$. Here $\alpha$ and $\beta$ are all roots such that $\alpha+\beta$ is also a root. Clearly, $\widetilde{K}$ is e-interpretable in $H$ as a set.
+Every $k\in K$ determines a corresponding element of $\widetilde{K}$ given by $y_{\alpha}\mapsto ky_{\alpha}$.
+Conversely, take $(k_{\alpha})_{\alpha\in\Phi}\in\widetilde{K}$.
+For every root $\alpha$ there is long $\beta\in\Phi$ such that $\alpha+\beta\in\Phi$, so $k_{\alpha}(y_{\alpha}),y_{\beta}]=[y_{\alpha},k_{\beta}(y_{\beta})]$.
+Recall that the Lie bracket $\colon\mathfrak{g}_{\alpha}\times\mathfrak{g}_{\beta}\to\mathfrak{g}_{\alpha+\beta}$ is non-degenerate (by lemma 3 if $\alpha$ is short).
+For any linear relation $\sum_{y_{\alpha}\in Y_{\alpha}}y_{\alpha}a_{\alpha}=0$ the equation implies that $\sum_{y_{\alpha}\in Y_{\alpha}}k_{\alpha}(y_{\alpha})a_{\alpha}=0$, so $k_{\alpha}$ continues to a unique linear map $k_{\alpha}\colon\mathfrak{g}_{\alpha}\to\mathfrak{g}_{\alpha}$.
+The equations on $(k_{\alpha})_{\alpha}$ imply that $k_{\alpha}(x),y]=k_{\alpha+\beta}([x,y])=[x,k_{\beta}(y)]$.
+We claim that such maps $k_{\alpha}$ are necessarily scalar.
+It suffices to consider $\Phi$ of rank $2$ and to check the claim for only one root $\alpha$ (by lemma 3).
+This is clear if one of root subspaces has rank $1$.
+To prove the claim we can also assume that $G$ splits and choose convenient parametrizations of the root subspaces $\mathfrak{g}_{\alpha}$.
+  • In the case $^{1}\!{\mathsf{A}_{3d-1,2}^{(d)}}$} we have
+    $k_{\mathrm{e}_{1}-\mathrm{e}_{2}}(x)\,y=k_{\mathrm{e}_{1}-\mathrm{e}_{3}}(xy)=x\,k_{\mathrm{e}_{2}-\mathrm{e}_{3}}(y)$
+    for $x,y\in\operatorname{M}(d,K)$.
+Hence $k_{\mathrm{e}_{1}-\mathrm{e}_{2}}(x)=k_{\mathrm{e}_{1}-\mathrm{e}_{3}}(x)=k_{\mathrm{e}_{2}-\mathrm{e}_{3}}(x)$ (i.e.
+all maps $k_{\alpha}$ coincide) and $k_{\alpha}(x)=x\,k_{\alpha}(1)=k_{\alpha}(1)\,x$, so $k_{\alpha}(1)$ lies in the center of $\operatorname{M}(d,K)$, and this center is precisely $K$.
+  • The case $^{1}\!{\mathsf{E}_{6,2}^{28}}$} is similar to the previous one. We have
+    $k_{\mathrm{e}_{1}-\mathrm{e}_{2}}(x)\,y=k_{\mathrm{e}_{1}-\mathrm{e}_{3}}(xy)=x\,k_{\mathrm{e}_{2}-\mathrm{e}_{3}}(y)$
+    for $x,y\in\mathrm{Z}(K)$, where $\mathrm{Z}(K)$ is the split octonion algebra over $K$.
+Hence all maps $k_{\alpha}$ coincide and $k_{\alpha}(x)=x\,k_{\alpha}(1)=k_{\alpha}(1)\,x$, so $k_{\alpha}(1)\in K$.
+  • In the case $^{2}\!{\mathsf{A}_{4d-1,2}^{(d)}}$} we have
+    $k_{\mathrm{e}_{1}-\mathrm{e}_{2}}(x,y)\circ z=(x,y)\circ k_{2\mathrm{e}_{2}}(z)=k_{\mathrm{e}_{1}+\mathrm{e}_{2}}((x,y)\circ z)$
+    for $x,y,z\in\operatorname{M}(d,K)$, where $(x,y)\circ z=(zx,yz)$.
+It follows that $k_{\mathrm{e}_{1}+\mathrm{e}_{2}}(x,y)=(x,y)\circ k_{\mathrm{e}_{2}}(z)$ and $k_{\mathrm{e}_{2}}(z)=zk_{\mathrm{e}_{2}}(1)=k_{\mathrm{e}_{2}}(1)z$, so $k_{\mathrm{e}_{2}}(1)\in K$.
+  • In the case $\mathsf{C}_{2d,2}^{(d)}$ we have
+    $k_{\mathrm{e}_{1}-\mathrm{e}_{2}}(x)\,y=x\,k_{2\mathrm{e}_{2}}(y)=k_{\mathrm{e}_{1}+\mathrm{e}_{2}}(xy)$
+    for $x\in\operatorname{M}(d,K)$ and symmetric matrix $y=y^{\mathrm{t}}\in\operatorname{M}(d,K)$.
+Then $k_{2\mathrm{e}_{2}}(y)=k_{\mathrm{e}_{1}+\mathrm{e}_{2}}(y)$ and $k_{\mathrm{e}_{1}+\mathrm{e}_{2}}(x)=k_{\mathrm{e}_{1}-\mathrm{e}_{2}}(x)=x\,k_{2\mathrm{e}_{2}}(1)$.
+So $k_{2\mathrm{e}_{2}}(y)=k_{2\mathrm{e}_{2}}(1)\,y=y\,k_{2\mathrm{e}_{2}}(y)$, and it easily follows that $k_{2\mathrm{e}_{2}}(1)\in K$.
+  • In the last case $^{1}\!{\mathsf{D}_{2d,2}^{(d)}}$} we have
+    $k_{\mathrm{e}_{1}-\mathrm{e}_{2}}(x)\,y=x\,k_{2\mathrm{e}_{2}}(y)=k_{\mathrm{e}_{1}+\mathrm{e}_{2}}(xy)$
+    for $x\in\operatorname{M}(d,K)$ and alternating $y\in\operatorname{M}(d,K)$ (i.e.
+$y_{ii}=0$ and $y_{ij}=-y_{ji}$).
+Since $d\geq 2$ is even (actually, a power of $2$), there is an invertible alternating matrix $J\in\operatorname{M}(d,K)$, say, the matrix of the split symplectic form.
+As in the previous case, $k_{2\mathrm{e}_{2}}(y)=k_{\mathrm{e}_{1}+\mathrm{e}_{2}}(y)$, $k_{\mathrm{e}_{1}-\mathrm{e}_{2}}(x)=x\,k_{2\mathrm{e}_{2}}(J)\,J^{-1}$, $k_{\mathrm{e}_{1}+\mathrm{e}_{2}}(x)=xJ^{-1}\,k_{\mathrm{e}_{1}+\mathrm{e}_{2}}(J)$.
+Then $k_{2\mathrm{e}_{2}}(y)=k_{2\mathrm{e}_{2}}(J)\,J^{-1}y=yJ^{-1}\,k_{2\mathrm{e}_{2}}(J)$ and $k_{2\mathrm{e}_{2}}(J)$ is alternating, so it easily follows that $k_{2\mathrm{e}_{2}}(J)\,J^{-1}\in K$.
 It remains to define ring operations on $\widetilde{K}$.
-The group operation is directly induced from $G(K)$.
-Finally, the product of $(k_{\alpha})_{\alpha}$ and $(l_{\alpha})_{\alpha}$ is such a family $(m_{\alpha})_{\alpha}$ that
-$[k_{\alpha}(x_{\alpha}),l_{\beta}(x_{\beta})]\equiv m_{\alpha+\beta}([x_{\alpha},x_{\beta}])\pmod{\operatorname{C}(G)(K)\,\prod_{\begin{subarray}{c}i\alpha+j\beta\in\Phi\\ i,j\geq 1\\ (i,j)\neq(1,1)\end{subarray}}G_{i\alpha+j\beta}}$
+The group operation is directly induced from $\mathfrak{g}_{\alpha}$.
+The product of $(k_{\alpha})_{\alpha}$ and $(l_{\alpha})_{\alpha}$ is such a family $(m_{\alpha})_{\alpha}$ that
+$k_{\alpha}(x_{\alpha}),l_{\beta}(x_{\beta})]=m_{\alpha+\beta}([x_{\alpha},x_{\beta}])$
 for $x_{\alpha}\in X_{\alpha}$, $x_{\beta}\in X_{\beta}$, and roots $\alpha$, $\beta$ such that $\alpha+\beta$ is also a root.
-It suffices to impose this condition only for one pair $(\alpha,\beta)$.
+It suffices to impose this condition only for one pair $(\alpha,\beta)$ with long $\alpha$.
 ∎
@@
 Let $K$ be a commutative ring with an enumeration by $\mathbb{N}$ or its finite subset such that the ring operations are computable.
-Let also $G$ be an isotropic reductive group over $K$ such that the rank of $\Phi$ is at least $2$.
+Let also $G$ be an isotropic reductive group scheme over $K$ such that the rank of $\Phi$ is at least $2$.
 Then the Diophantine problems for $K$ and $G(K)$ are equivalent, i.e.
 they reduce to each other by algorithms.
-Here the enumeration on $G(K)$ is obtained from some closed embedding $G\subseteq\mathbb{A}^{n}_{K}$, it is independent of the embedding up to a computable permutation.
+Here the enumeration on $G(K)$ is obtained from some closed embedding $G\subseteq\mathbb{A}^{n}_{K}$, it is independent of embedding up to a computable permutation.
 Proof.
 Clearly, the group operations on $G(K)$ are also computable.
-Every Diophantine subset of $G(K)^{k}$ may be considered as a Diophantine subset of $K^{kn}$ using the e-interpretation of $G(K)$ in $K$ and this transformation is clearly computable in terms of defining formulae.
+Every Diophantine subset of $G(K)^{k}$ can be considered as a Diophantine subset of $K^{kn}$ using the e-interpretation of $G(K)$ in $K$ and this transformation is clearly computable in terms of defining formulae.
 In the other direction apply theorem 5.